A Complete Bibliography of Publications in The Journal of Supercomputing: 2025–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

30 September 2025 Version 1.08

Title word cross-reference

(K_{11} − C_{11})^{n} [218]. (n, k) [550]. ^{2} [507, 647]. C [439]. d [1120]. γ [250]. H [216]. K [23, 65, 79, 274, 456, 520, 633, 707, 744, 1051]. n [744]. p [53, 197]. Q [103, 611]. r, s, t [173]. U [987]. Z [230].

-ary [744]. -Block [230]. -cube [744]. -dimensional [1120]. -dominant [216]. -factor-critical [456]. -graphyne [250]. -means [65, 274, 439, 520, 707]. -nearest [23, 79, 633, 1051]. -network-based [611]. -rung [103]. -sequence [197]. -spherical [173]. -star [550]. -type [53].

1 [201]. 1-Extra [164]. 1-memristor [1101]. 1-transistor [1101]. 16 [709]. 19 [113, 693].

1M1T [1101].

2 [690]. 2-connected [501]. 2-disjoint [273]. 2-transformer [690]. 2-tuple [103]. 2.0 [398]. 23 [398]. 2D [203, 625].

3-component [164]. 3-path [900].

3-path-connectivity [245]. 3000 [38]. 3D [19, 39, 276, 449, 455, 466, 487, 684, 758, 805, 864, 889, 1026, 1108, 1156].

4 [454]. 4-D [454]. 4.0 [1078]. 4D [472].

512 [355]. 5G [72, 328, 593, 685]. 5G/ [593].

6 [1014]. 6-DoF [1014]. 6G [555, 576, 593, 932, 958].

AAMB [695]. ABAC [911]. ABC [777, 878]. ABC-SVM [878]. ABIDS [708]. ABIDS-VEM [708]. accelerate [734, 1065]. accelerated [453, 1165]. Accelerating [315, 349, 685, 915]. Acceleration [264, 364, 589]. accelerator [190, 586, 842, 1148]. accelerators [23, 536, 573]. access [44, 91, 213, 233, 396, 436, 537, 911, 1064]. access-control [91]. accident [546, 698]. AccPRET [536]. accuracy [399, 714, 865, 967, 988, 1038, 1089]. accurate [347, 373, 546, 977, 1003, 1092, 1112]. achieve [541]. ACIMNet [855]. ACO [257]. acoustic [560, 595]. acquiring [617]. across [249]. action [10, 12, 37, 311, 408]. Activation [945]. activity [342, 887, 956, 957]. actor [350]. actor-critic [350]. adaptation [376, 500, 575, 1098, 1113]. adapting [137, 868]. Adaptive [66, 80, 92, 98, 112, 139, 212, 302, 336, 392, 458, 481, 501, 514, 518, 572, 588, 601, 626, 639, 649, 678, 683, 690, 730, 735, 743, 750, 754, 775, 835, 860, 866, 882, 894, 935, 937, 941, 943, 948, 954, 971, 1032, 1043, 1049, 1072, 1092, 1110, 1126, 1135, 1137, 1142, 1159, 1164]. AdApTS [1164]. AdaRNN [1055]. adder [482]. additive [551]. Addressing [615, 1143]. ADH [505]. ADH-YOLO [505]. adhoc [1069]. adjusted [503]. ADOX [212]. ADOX-SSA [212]. Adpl [847]. ADV [34]. ADV-YOLO [34]. advanced [547, 604]. Advancements [362]. Advancing [19, 409, 1128, 1156]. Adversarial [168, 239, 289, 342, 609, 726, 756, 815, 888, 943, 959, 1140]. adversarial-majority [959]. AE [331, 657]. AE-MCDM [657]. AE-UNet [331]. Aerial [95, 105, 302, 414, 492, 605, 662, 830, 898, 1034, 1157]. AF [727]. AF-YOLOv8 [727]. AFF [754]. AFF-DSTnet [754]. affine [500]. AFMD [920]. against [518, 806, 959, 962, 1145]. agent [150, 279, 381, 772, 784, 844, 872, 998]. agent-based [772]. aggregate [607]. aggregation [43, 63, 269, 311, 698, 781, 823, 837, 952, 967, 1133]. aggregations [294]. agnostic [1048]. agreement [9, 149]. agriculture [280]. AGV [786]. AI [272, 280, 340, 349, 383, 441, 539, 553, 602, 603, 818, 933, 1014, 1020, 1151]. AI-based [280]. AI-driven [272, 383, 818]. AI-powered [553, 603, 933, 1014]. AIGC [417]. AIGC-generated [417]. AIoT [462]. air [73, 394]. airborne [481]. aircraft [337, 551, 577]. airport [505, 1162]. AISBench [441]. Aitken [264]. ALCARO [330]. algebra [327]. algebra-enhanced [327]. Algebraic [53]. algorithm [7, 8, 13, 18, 21, 22, 47, 52, 54, 55, 66, 79, 94, 98, 99, 110, 113, 116, 124, 135, 137, 144, 150, 170, 182, 209, 227, 236, 244, 251, 252, 266, 285, 290, 292, 295, 302, 313, 318, 322, 323, 348, 353, 377, 392, 403, 432, 440, 450, 483, 494, 517, 518, 521, 532, 565, 571, 579, 582, 596, 597, 599, 619, 622, 623, 643, 662, 692, 694, 703, 707, 709, 711, 721, 727, 734, 750, 758, 766, 767, 802, 813, 822, 830, 846, 860, 877, 894, 897, 929, 935, 944, 950, 955, 998, 1011, 1016, 1047, 1052, 1060, 1076, 1118, 1119, 1129, 1138, 1142]. algorithm [32, 67, 91, 188, 202, 261, 264, 365, 374, 380, 384, 433, 434, 471, 557, 606, 612, 624, 665, 680, 719, 769, 942, 948, 951, 969, 974, 987, 996, 1018, 1043, 1059, 1116]. algorithm-based [137]. algorithm-optimized [1016]. algorithmic [797]. algorithms [20, 88, 132, 172, 247, 257, 468, 475, 660, 777, 891, 903, 910, 1039]. Alignment [358, 376, 599, 741, 914, 940, 1041, 1122]. All-MLP-based [795]. all-weather [1162]. allergic [53]. allocation [11, 118, 192, 237, 333, 360, 459, 504, 557, 561, 787, 820, 825, 923, 989, 1066, 1076, 1105, 1166]. ALMASH [301]. along [379]. alternative [703]. altitude [414, 854]. Alzheimer [423]. AlzONet [423]. Ameliorating [877]. among [736]. analysis [1, 9, 46, 51, 53, 55, 58, 64, 86, 93, 131, 136, 179, 195, 329, 341, 342, 351, 447, 477, 484, 497, 498, 531, 550, 600, 610, 616, 635, 644, 668, 715, 746, 748, 778, 795, 801, 804, 876, 881, 895, 906, 918, 939, 943, 980, 994, 1039, 1149]. analytical [268, 1015]. analytics [547]. Analyzing [598, 625, 651]. Android [491, 834]. ANFIS [201, 1039]. angle [98]. angle-based [98]. annealing [147]. annotated [1112]. annotation [587]. Anomalies [243]. anomaly [195, 282, 297, 320, 375, 404, 448, 495, 614, 708, 754, 763, 875, 933, 1010, 1087]. anomaly-based [708]. anonymity [301]. anonymity-based [301]. anonymization [134]. anonymous [1150]. answering [1071]. ant [47, 972]. antenna [318]. antenna-array [318]. Antivirus [151]. any [465]. anything [868]. AODV [379]. AoSoA [1064]. AoSoA-based [1064]. AP [72, 333]. Apache [214, 506]. API [581]. APPBoost [518]. appearance [1077]. Application [7, 53, 118, 197, 205, 280, 294, 616, 646, 681, 758, 767, 779, 835, 883, 950, 975, 1027, 1054, 1064, 1133]. applications [94, 108, 109, 136, 167, 173, 253, 360, 428, 445, 588, 666, 686, 740, 764, 916, 923, 949, 1056, 1061]. applied [159, 422]. applying [770]. approach [1, 6, 24, 29, 58, 88, 92, 123, 128, 134, 137, 142, 165, 166, 184, 198, 206, 207, 254, 260, 263, 274, 283, 285, 314, 316, 339, 362, 403, 419, 424, 457, 478, 479, 501, 506, 526, 547, 560, 564, 637, 647, 650, 657, 693, 712, 731, 772, 786, 820, 891, 906, 929, 936, 964, 965, 991, 992, 1000, 1095, 1110, 1144, 1166]. approaches [20, 76, 347, 747, 763, 851, 921, 1039]. approximate [109, 121, 399, 1164]. Approximating [930]. approximation [458, 1133]. Arabic [187]. arbitrary [305]. arbitration [473]. architecture [172, 252, 271, 330, 344, 359, 536, 555, 661, 667, 739, 753, 773, 818, 845, 912, 1030, 1083, 1094, 1101, 1112]. architectures [315, 766]. archive [751]. area [6, 291, 515, 965, 1133]. areas [845]. ARGCN [86]. ARIMA [366]. arithmetic [23, 302]. arm [214, 1129]. Arnold [39, 1135]. arrangement [218]. arrangements [39]. array [318, 573]. artery [267, 687]. articles [381]. Artificial [62, 110, 150, 270, 779, 882, 950]. ary [744]. ASOD [605]. ASOD-YOLOX [605]. Aspect [46, 64, 125, 131, 189, 200, 341, 531, 620, 644, 668]. Aspect-aware [64]. Aspect-based [64, 131, 341, 531, 644]. aspect-focused [341]. aspect-level [46, 125, 189, 200, 668]. Assessing [180, 648]. assessment [73, 205, 285, 436, 461, 967]. Assignment [95, 1054]. assist [410]. assisted [8, 149, 165, 231, 308, 576, 646, 1011, 1089, 1129]. associated [243]. association [328, 708]. association-based [328]. associations [36, 72]. assurance [247]. AST [587]. asymmetric [656]. asynchronous [7, 486, 621, 749]. atomic [484]. atomistic [542]. Attack [62, 444, 541, 645, 848]. attacks [69, 252, 313, 474, 506, 815, 908, 959, 962, 1089, 1145]. Attention [18, 26, 28, 31, 40, 70, 74, 82, 115, 125, 126, 157, 191, 199, 240, 242, 267, 293, 317, 331, 345, 405, 429, 470, 474, 527, 529, 551, 560, 591, 595, 618, 620, 638, 656, 688, 706, 732, 778, 794, 804, 807, 814, 826, 840, 841, 861, 876, 892, 907, 908, 920, 934, 956, 980, 993, 997, 999, 1019, 1025, 1033, 1050, 1057, 1074, 1083, 1111, 1126, 1158, 1159]. attention-based [317, 980, 1033]. attention-enhanced [125]. attention-guided [840, 934]. attentional [86, 943, 968]. attentionrub [1094]. attentive [847]. attribute [44, 167, 181, 421, 528, 816]. attribute-based [44, 421, 816]. attributes [707]. attribution [25]. attributive [1133]. auction [816]. audio [678, 1002]. audio-visual [678]. Audiovisual [31]. auditing [539]. augmentation [114, 208, 375, 530, 828, 1089]. augmented [185, 897, 1114]. authenticated [9, 149, 1150]. authentication [5, 119, 233, 301, 695, 831, 850, 905, 1130]. authentication-based [119]. authentications [831]. authorial [151]. authorization [641]. auto [228, 346, 470, 583, 1139, 1152]. auto-correlation [1139]. auto-encoder [346, 470]. auto-encoder-based [583]. auto-negotiation [228]. auto-tuning [1152]. autoencoder [5, 16, 62, 118, 231, 234, 657, 942]. autoencoder-based [234, 657]. autoencoder-enhanced [942]. automata [686, 772]. Automated [1, 600]. Automatic [597, 893, 995, 1064]. automatically [150]. automation [685]. automaton [99]. automotive [1031]. autonomic [360]. autonomous [379, 380, 718, 764, 866, 951, 1046, 1073]. autoregression [752]. autoregressive [724]. autoscaling [923]. autotuning [283]. Availability [436]. average [99, 1086]. average-roulette [99]. avoidance [604, 1129]. AVX [355]. AVX-512 [355]. aware [2, 3, 15, 27, 45, 52, 64, 121, 122, 226, 299, 306, 376, 377, 426, 485, 682, 687, 688, 700, 725, 762, 776, 784, 905, 924, 965, 1022, 1025, 1026, 1029, 1032, 1069]. aware-based [1022]. awareness [84, 375, 637, 749]. axiomatic [1000].

B [775]. B-spline [775]. B5G [538]. backoff [213]. backpropagation [398]. backscatter [362]. backup [435]. backward [65]. bag [285]. BAHA [950]. balance [458, 607]. Balanced [60, 273, 678, 950]. Balancing [52, 109, 328, 443, 469, 643, 865, 874, 893, 929, 953, 963, 972, 1149]. bald [251]. band [976]. Bandwidth [175]. bank [278]. barbed [222]. Barcelona [1006]. bare [1165]. bare-bone [1165]. base [27]. Based [4, 6, 10, 16, 18, 21, 24, 33, 34, 57, 65, 66, 69, 98, 100, 113, 121, 124, 129, 160, 162, 171, 210, 215, 240, 241, 250, 302, 333, 337, 348, 353, 363, 379, 382, 385, 448, 463, 483, 487, 499, 528, 543, 544, 561, 564, 579, 590, 596, 611, 614, 621, 643, 644, 657, 660, 673, 678, 701, 707, 710, 733, 736, 752, 754, 768, 772, 773, 802, 813, 821, 829, 833, 848, 863, 868, 890, 895, 909, 912, 926, 970, 975, 978, 995, 997, 999, 1003, 1009, 1012, 1021, 1028, 1040, 1068, 1076, 1081–1083, 1103, 1104, 1110, 1123, 1127, 1133, 1138, 1160, 1165]. based [9, 13, 22, 25, 28, 31, 37, 39, 42, 58, 62, 68, 85, 89, 128, 140, 144, 153, 156, 173, 181, 184, 203, 212, 255, 270, 280, 292, 304, 317, 318, 322, 323, 343, 349, 351, 369, 372, 378, 381, 402, 406, 413, 415, 424, 430, 431, 440, 447, 450, 458, 462, 468, 474, 485, 491, 494, 519, 540, 582, 585, 588, 627, 637, 639, 640, 652, 667, 671, 672, 685, 711, 743, 750, 753, 762, 788, 793, 795, 827, 828, 834, 836, 855, 857, 867, 872, 891, 945, 955, 960, 973, 990, 1008, 1037, 1041, 1051, 1062, 1064, 1073, 1079, 1090, 1094, 1100, 1112, 1121, 1137, 1145, 1164, 1166]. based [5, 8, 12, 35, 44, 47, 54, 76, 79, 96, 97, 110, 123, 126, 130, 132, 134, 137, 142, 188, 189, 196, 209, 219, 227–229, 232, 260, 262, 266, 267, 276, 277, 290, 300, 301, 328, 330, 335, 352, 365, 373, 389, 405, 421, 433, 451, 453, 475, 476, 490, 495, 520, 529, 532, 549, 551, 558, 583, 600, 613, 624, 626, 645, 650, 708, 721, 728, 729, 740, 741, 748, 759, 796, 811, 830, 832, 835, 854, 881, 892, 899, 908, 919, 920, 922, 929, 935, 944, 954, 963, 968, 980, 1001, 1019, 1022, 1025, 1033, 1042, 1044, 1056, 1087, 1111, 1115, 1118, 1119, 1140, 1141, 1149]. based [14, 40, 49, 64, 67, 111, 119, 131, 143, 148, 169, 172, 213, 221, 234, 248, 264, 269, 324, 341, 344, 350, 435, 438, 454, 461, 471, 488, 492, 505, 508, 515, 523, 525, 531, 534, 554, 557, 562, 587, 605, 633, 665, 695, 702, 706, 769, 815, 816, 838, 844, 858, 878, 889, 898, 927, 941, 951, 974, 987, 988, 1038, 1043, 1045, 1055, 1059, 1066, 1108, 1120, 1134, 1148]. bases [243]. basis [152, 801]. basket [826]. bat [137]. Batched [1045]. batteries [973]. Bayesian [55]. BDLT [271]. BDLT-IoMT [271]. beam [275, 642]. bearing [262]. Beetle [90, 403]. behavior [457, 540, 616, 635, 702, 1116]. behavioral [768, 793]. behaviour [284]. beluga [1138]. benchmark [441, 475, 635, 755]. benefits [648]. Bernoulli [201]. BERT [163, 187, 317, 381]. best [837]. best-worst [837]. better [785]. betweenness [1125]. Beyond [72, 347, 576, 593, 685, 823]. BFL [378]. BFT [486]. bi [31, 177, 342, 356, 529, 532]. bi-clique [356]. bi-directional [342]. bi-layer [31]. bi-level [529]. bias [540, 1093]. bias-corrected [1093]. Biclustering [949]. bidirectional [117, 778, 866, 963, 983, 1159]. bifurcation [939]. big [165, 181, 615, 949]. BiGRU [1095]. Bilateral [136]. Bilinear [437]. BiLSTM [162, 502]. binary [190, 209, 225, 356, 871, 1054, 1165]. binary/ternary [190]. bioinformatics [949]. biological [939]. biomarkers [927]. biomedical [171]. biometric [528, 570]. bionic [236]. bipartite [48, 356]. Bipolar [294]. bird [512]. bit [183]. bit-level [183]. Bitcoin [77]. bits [210]. BKA [461]. BL [502]. black [612]. blades [1097]. blindness [514]. block [77, 160, 224, 230, 422, 1138, 1159]. block-level [1138]. blockage [511]. Blockchain [61, 97, 111, 128, 166, 219, 232, 271, 317, 321, 332, 378, 411, 463, 469, 498, 522, 585, 661, 797, 800, 816, 833, 850, 905, 1075, 1078, 1081]. blockchain-based [463, 585, 816, 1081]. Blockchain-enabled [469, 797, 905, 1078]. blockchain-federated [378]. blockchains [695]. blocks [888]. blood [410, 669, 1104]. Bluelight [562]. BN [545]. boards [974]. body [106, 342, 965]. body-worn [342]. bone [1165]. boosted [368]. boosting [229, 518, 1099]. border [1133]. borderline [756]. bot [598]. botnet [506]. bottleneck [1096]. bound [102, 1149]. boundary [913]. bounds [535]. boxes [806]. BPTCN [733]. BR [55]. brain [19, 235, 345, 423, 502, 684, 1132]. brake [723]. braking [723]. branch [127, 507, 591, 718, 733, 927, 1134, 1140]. brands [1076]. breakpoints [222]. breast [674, 836, 1155]. broadcast [247]. broadcasting [621]. brooms [53]. bubble [164]. bubble-sort [164]. budget [286, 994]. budget-feasible [286]. buffer [1086]. bug [1146]. building [706, 859, 1025]. built [109]. built-in [109]. bullying [284]. burden [607]. bushing [365]. BXIv3 [525]. BXIv3-based [525]. bypass [753]. Byzantine [711, 959]. Byzantine-robust [959].

C [836]. C&M [502]. C2 [330]. C3 [558]. CA [433]. cache [45, 400]. caching [637, 1037, 1058]. calculation [218]. calculations [465]. calculus [90]. calls [659]. camera [25, 1041]. camouflage [157]. camouflaged [559]. cancellation [595]. cancer [674, 836, 1155]. Canny [230]. capabilities [196]. capability [757]. capacity [822, 1149]. Capsule [86, 424, 489, 849]. caption [644]. capuchin [236]. Caputo [714]. carbon [488]. cardiac [893]. cardiovascular [402, 1039]. Carlo [374, 562]. carotid [267]. carrier [576, 1053]. cars [1164]. Cartesian [517]. cascade [720]. case [308, 351, 445, 447, 542, 728, 1151]. cases [681]. cat [113, 318]. categories [213]. categorization [335]. categorization-based [335]. category [375]. category-guided [375]. causal [247, 358, 386, 575]. causal-contrastive [575]. cause [195]. cavity [385]. CBM [155]. CBR [77]. CCLNet [658]. CD [190, 854]. CD-MAC [190]. CD-YOLOv8s [854]. CDA [726]. CDMANet [242]. CDNRocks [57]. CE [579]. CELAT [1111]. CELAT-DiffNet [1111]. cell [410, 459, 649, 686, 721, 1104]. Cellular [99, 322, 333, 686, 772, 976]. censoring [760]. center [4, 211, 1006]. centers [184]. central [242]. centric [166, 537, 736]. CEPDNet [100]. Cerebral [106]. certificateless [653]. CGWRIME [368]. chain [89, 372, 396, 561, 922]. chain-based [89, 922]. chains [324]. challenges [291, 648, 747, 780, 932, 949, 958, 1006].

Change [54, 706, 710]. change-point [710]. changes [71]. Channel [126, 169, 252, 270, 405, 455, 461, 469, 499, 591, 762, 855, 993, 1026, 1050, 1087, 1111]. channel-based [762]. channel-enhanced [1111]. channels [878]. chaotic [108, 152, 521, 665, 829, 916]. character [617]. characterization [887]. chargers [298]. charging [298, 475, 837]. ChatGPT [629]. chatHPC [194]. Chebyshev [179, 252, 519]. ChebyshevNet [179]. check [296, 703]. chemical [352, 828]. chest [809, 1096]. chicken [770]. China [351, 447]. Chinese [50, 200]. Chinese-oriented [200]. chip [753, 855, 1086]. chlorophyll [265]. chlorophyll-a [265]. churn [178, 1151]. CILOS [103]. CIMS [229]. CIMS-based [229]. cipher [224]. circuit [117, 354, 391, 450, 974]. circular [1133]. cities [97, 403, 506]. city [637, 818]. civil [776]. CL [1161]. clairvoyant [1085]. class [16, 74, 167, 248, 630, 659, 701]. class-consistent [167]. class-imbalanced [248, 659]. class-wise [16]. classical [347, 455, 1005, 1080]. Classification [23, 26, 54, 66, 101, 114, 125, 126, 189, 200, 217, 248, 285, 344, 345, 370, 376, 387, 395, 397, 405, 409, 500, 502, 507, 541, 608, 630, 636, 654, 667, 670, 679, 683, 730, 746, 771, 782, 834– 836, 841, 842, 846, 871, 979, 1007, 1044, 1096, 1104, 1118, 1134, 1159]. classifier [154, 896, 1129]. classifiers [1165]. classify [526]. classifying [1098]. CLDE [556]. Click [984]. Click-through [984]. client [69, 196, 552]. client-side [69]. CLIP [174, 804]. CLIP-driven [174, 804]. CLIP3D [805]. clique [356]. clockwork [412]. closeness [1079]. closeness-based [1079]. closure [67]. Cloud [4, 12, 65, 89, 127, 128, 135, 141, 165, 176, 181, 237, 295, 361, 387, 421, 424, 462, 469, 474, 539, 544, 564, 588, 590, 611, 615, 619, 643, 666, 755, 770, 774, 825, 845, 859, 874, 880, 921, 923, 929, 947, 964, 969, 971, 972, 994, 1008, 1023, 1108, 1128]. cloud-assisted [165]. Cloud-based [12, 462]. cloud-edge [539]. cloud-edge-terminal [611]. cloud-fog [755]. cloud-native [923]. Cloud-WAVECAP [424]. clouds [1026]. Cluster [2, 470, 561, 616, 1012]. cluster-based [561]. Cluster-guided [470]. clustered [863, 877]. Clustering [2, 65, 98, 135, 143, 150, 171, 193, 209, 274, 379, 390, 439, 594, 640, 649, 707, 851, 874, 899, 1018, 1062, 1092, 1143, 1153]. Clustering-based [143, 171, 1062]. clustering-efficient [1143]. clusters [104, 427, 460, 701]. cnn [5, 24, 100, 144, 162, 403, 453, 462, 502, 628, 658, 684, 777, 824, 966, 1094]. CNN-attentionrub [1094]. CNN-based [5, 24, 100]. CNN-LSTM [658]. CNN-swin [684]. CNN-transformer [824]. CNNs [60, 399, 1096]. CNSD [768]. CNSD-Par [768]. co [452, 571, 842]. co-design [842]. co-execution [452]. co-location [571]. coal [968]. coal-gangue [968]. coalition [132]. coalitions [557]. CoAP [906]. Coarse [480]. Coarse-to-fine [480]. coati [996]. code [486, 587, 741, 785, 1146]. coding [85, 210]. codistillation [977]. coefficient [394, 779]. coefficients [675]. coevolutionary [59, 287]. cognitive [133, 547]. coherence [887, 1040]. cold [234, 1061]. collaborated [825]. collaboration [84, 158, 368, 497]. Collaborative [54, 59, 79, 114, 124, 287, 369, 396, 487, 557, 791, 1063]. Collatz [663, 1163]. collection [123]. collective [766]. collusion [839]. collusion-resistant [839]. colony [972]. color [108, 858, 883]. colored [247]. coloring [309, 1043]. combat [73, 279]. Combating [69]. combinatorial [728, 1147]. combined [161, 326]. combining [94, 258, 819, 845, 990, 993]. Comment [149]. commerce [229]. commercialization [163, 843]. common [571, 658, 710]. communication [11, 283, 333, 362, 529, 538, 576, 642, 784, 796, 818, 865, 1012, 1088]. communication-constrained [11]. communications [277]. community [182, 216, 450, 729]. CoMP [784]. CoMP-LG [784]. compact [77]. compactions [156]. comparative [922, 1039]. comparison [665, 1120, 1133]. compensation [109, 668]. competition [368]. competition-boosted [368]. Competitive [556, 577]. compiler [442]. complementarity [559]. complementary [765, 868, 991]. complete [180, 334]. completion [572, 938]. Complex [36, 159, 182, 402, 503, 569, 612, 713, 767, 778, 985, 995, 1002]. complex-valued [778, 985]. component [120, 164, 550, 918]. components [918]. composite [331, 1056]. composites [13]. Composition [940, 1060]. compound [352]. comprehension [50]. Comprehensive [291, 359, 475, 676, 711, 763, 932, 958, 1031, 1128, 1149]. compressed [144, 1140]. compression [96, 392, 782, 810, 865, 973]. computable [57]. Computation [175, 459, 563, 764, 825, 1005]. Computation-efficient [764]. computational [152, 477, 921]. computationally [1014]. computer [172, 568]. computers [1080]. computility [1066]. Computing [15, 24, 29, 35, 41, 52, 76, 84, 100, 116, 121, 127, 186, 190, 192, 196, 254, 257, 288, 289, 295, 310, 360, 377, 379, 399, 418, 419, 421, 426, 436, 469, 506, 539, 564, 578, 602, 612, 619, 639, 643, 648, 661, 700, 705, 755, 763, 770, 773, 787, 796, 827, 839, 845, 859, 874, 882, 921, 924, 929, 949, 969, 972, 992, 1000, 1008, 1012, 1013, 1020, 1044, 1075, 1128, 1144, 1147, 1163, 1164]. computing-based [1164]. computing-inspired [1013]. concave [320]. concentrations [265]. conception [170]. concise [197, 571]. condition [563]. Conditional [14, 238, 509, 712, 1051]. conditioned [30]. conditions [713]. confidence [656]. confidence-driven [656]. configuration [1152]. conflict [909]. conflict-based [909]. confluence [332]. congestion [965, 1003]. conjecture [663]. conjugate [55]. connected [501]. connection [250, 976]. connections [207]. connectivity [72, 164, 180, 245, 524, 900]. consecutive [1163]. Consensus [61, 91, 97, 323, 711]. considering [779, 793, 1086]. consistency [299, 1093]. consistent [167]. consolidation [417]. consortium [411, 816]. constellation [989]. constrained [11, 190, 286, 558, 636, 819, 913, 1011, 1110, 1129]. constraint [116, 132, 849]. constraints [4, 463, 994, 999, 1086]. Constructing [259, 594]. construction [334, 649, 654]. constructions [552]. consumer [789]. consumption [314, 329, 403, 460, 650, 918]. container [15, 260, 451, 768, 1062]. container-based [451]. content [168, 637]. content-type [637]. contention [213, 473]. context [71, 311, 413, 594, 668, 676, 776, 856, 1026, 1091]. context-aware [776, 1026]. contextual [176]. contextualization [823]. continual [870, 896]. Continuous [5, 389]. Continuous-time [389]. contour [230, 762]. contract [372, 508, 587]. contraction [354]. contracts [44, 372, 1081]. contrast [509, 1032]. contrast-aware [1032]. contrastive [37, 101, 139, 153, 239, 256, 575, 583, 613, 814, 857, 875, 914, 971]. contribution [255]. contributions [503]. control [44, 91, 185, 204, 396, 422, 626, 696, 742, 911, 965, 1002, 1053, 1073]. Controllable [108, 528]. controlled [1028]. controller [435, 515, 963, 1125]. Conv2D [545]. Conv2D-ReLU-BN [545]. ConvBiFuseNet [345]. convective [513]. convergence [148, 663]. Convergent [465]. conversation [954]. conversational [153, 799]. convex [592, 1008]. ConvFormer [527]. convolution [26, 235, 554, 613, 761, 1050]. convolutional [18, 25, 27, 193, 231, 277, 319, 341, 390, 395, 416, 437, 455, 466, 531, 688, 733, 764, 835, 841, 912, 980, 997, 1016, 1101, 1132, 1159]. convolutions [1045, 1126]. cooperation [430]. Cooperative [59, 287, 699, 1037]. coordinate [191, 784]. coordinate-aware [784]. coordination [584]. core [59, 287, 496]. cores [248]. coronary [687]. corpora [187]. corrected [1093]. Correction [287, 383, 384, 404, 405, 419, 447, 484, 590, 603, 664, 697, 934, 958, 990, 1030]. correction-optimized [697]. correlated [495]. correlation [348, 1139]. correlation-guided [348]. correntropy [1131]. cost [116, 132, 288, 588, 615, 810]. cost-effective [288]. counterfactual [938]. counterpart [1020]. coverage [83, 944]. coverage-based [944]. covers [48, 273]. COVID [113, 693]. COVID-19 [113, 693]. CovMedCare [332]. CPCS [430]. CPU [59, 287, 428, 452, 473]. CPU-GPU [452]. CPUs [1045]. crack [721, 746]. criteria [89, 657, 837]. criterion [1]. critic [350]. critical [456]. criticality [387, 1085]. CRKG [258]. Crns [174]. Cross [127, 139, 299, 316, 431, 461, 507, 591, 683, 690, 695, 776, 849, 870, 908, 984, 1087, 1098, 1122]. cross-attention [908]. cross-aware [299]. cross-channel [1087]. cross-domain [139, 683, 695, 1122]. cross-enhancement [507]. cross-fusion [591]. cross-modal [127]. cross-modality [849]. cross-network [984]. cross-project [316]. cross-stage [431, 776]. Cross-subject [1098]. cross-task [870]. crossover [998]. crossroad [67]. crow [972]. crowd [702]. crowdfunding [263]. crowding [622]. crowdsensing [286, 457]. CRS [153]. cryptographic [585, 845]. cryptography [933]. cryptosystem [903]. Cs [711]. Cs-pbft [711]. CSCNet [1026]. CSI [990]. CT [331]. CU [795]. CU-SEMLP [795]. cube [180, 393, 524, 744, 883]. cubes [96, 334]. cubic [103]. cuckoo [755]. cue [1077]. cues [1077]. cultural [948]. curiosity [227]. currencies [433]. curvature [1108]. custom [634]. customer [178, 1151]. customized [398]. CV [201]. CVSR [985]. cyber [51, 142, 284, 547]. cyber-bullying [284]. cyber-physical [51, 547]. cyberattacks [709, 731]. cyberphysical [446]. cybersecurity [271, 277, 1005, 1088]. cybertwin [430]. cycle [1060]. cyclic [722]. cylinder [185].

D [242, 454, 559]. D2D [976]. DAG [813]. DAG-structured [813]. damage [380, 551, 776]. damped [704]. dandelion [458]. dangerous [1116]. DAP [77]. DAP-CBR [77]. Dara [376]. DARVFL [807]. DARVFL-LSTM [807]. Data [4, 41, 57, 63, 91, 123, 134, 143, 154, 165– 167, 181, 184, 210, 211, 225, 238, 271, 272, 300, 316, 342, 359, 375, 378, 383, 392, 396, 428, 435, 443, 465, 472, 495, 497, 507, 509, 530, 535, 547, 565, 571, 600, 615, 617, 629, 633, 637, 647, 673, 696, 707, 708, 712, 730, 766, 785, 792, 797, 800, 828, 833, 860, 863, 881, 902, 944, 949, 1045, 1051, 1053, 1063, 1078, 1089, 1092, 1102, 1116, 1134, 1146, 1153]. Data-dependent [535]. Data-driven [465, 696, 1146]. data-free [509]. database [81, 844]. dataset [31, 635, 1112]. datasets [118, 308, 417, 443, 628, 1063, 1160]. DBSCAN [150]. DCLMD [594]. Dcsca [591]. Dcsca-Net [591]. DDoS [474, 598, 848]. DDQN [276]. deadline [4, 994]. debiasing [530]. deblur [431]. debris [1162]. Decentralized [121, 469, 717, 835]. decimal [482]. decision [29, 229, 353, 354, 515, 544, 590, 657, 837, 866, 1073, 1118]. decision-making [29, 657, 837, 866, 1073]. decoder [157, 732]. decoding [658]. decomposition [160, 812, 832, 920, 1028, 1056]. decompositions [806]. Deep [6, 35, 72, 92, 130, 161, 213, 223, 241, 280, 284, 288, 295, 315, 336, 337, 361, 373, 390, 395, 410, 423, 438, 448, 462, 545, 555, 560, 564, 589, 596, 598, 611, 614, 631, 634, 685, 763, 801, 825, 844, 871, 874, 881, 926, 942, 952, 953, 986, 1008, 1017, 1028, 1067, 1073, 1123, 1152, 1155]. deep-reinforcement-learning-based [564]. Deepat [337]. deepfake [1140]. defect [32, 40, 155, 239, 261, 415, 489, 494, 761, 781, 788, 974]. defects [1097, 1103]. defend [962]. defense [766, 869, 1031]. defined [291, 426, 485, 846, 963]. defining [845]. Defocus [431]. deformable [480, 1126]. degradation [1085]. degree [145, 324]. degree-based [324]. degree-wise [145]. dehazing [169, 750, 824, 1019, 1032]. delay [52, 564]. delay-sensitive [564]. deletion [670, 729]. deliver [263]. delivery [637]. demand [55, 475, 1069]. democratic [771]. demonstrations [594]. denial [444]. DenLsNet [836]. DenLsNet-C [836]. denoising [3, 100, 159, 1010, 1040, 1072]. Dense [856, 1072, 1132]. DenseNet [836]. Density [143, 814]. dependencies [189]. Dependency [45, 82, 125, 688]. Dependency-aware [45]. dependency-oriented [82]. dependent [203, 535, 648, 1000]. deploy [445]. deployment [84, 523, 538]. depression [93]. DepressionFeature [93]. depth [431, 761, 910, 1041, 1091, 1123]. depth-of-field [431]. deraining [937]. derivation [225]. derivatives [714]. descent [321, 514]. Design [9, 148, 149, 251, 428, 482, 533, 577, 842, 885, 945, 963, 1099, 1146]. designing [254, 419]. Detecting [284, 709, 794, 1067]. Detection [16, 22, 28, 34, 42, 62, 105, 106, 110, 151, 177, 191, 222, 231, 249, 254, 282, 288, 297, 303, 308, 313, 320, 331, 381, 404, 410, 413, 415, 419, 424, 448–450, 455, 474, 491, 494, 514, 532, 541, 553, 558, 569, 579, 580, 583, 603, 609, 614, 618, 628, 629, 631, 632, 636, 645, 659, 662, 664, 670, 692, 703, 710, 713, 721, 727, 754, 763, 771, 776, 781, 783, 791, 792, 805, 830, 848, 854, 855, 864, 867, 885, 907, 908, 913, 933, 936, 946, 955, 960, 1014, 1016, 1021, 1035, 1068, 1089, 1097, 1099, 1103, 1112, 1114, 1115, 1136, 1154, 1155, 1161]. detection [32, 40, 67, 75, 142, 155, 202, 230, 239, 261, 275, 277, 290, 299, 307, 312, 327, 365, 375, 380, 384, 417, 438, 444, 466, 467, 476, 492, 495, 505, 508, 515, 548, 551, 554, 559, 572, 605, 638, 676, 699, 702, 706, 708, 729, 761, 790, 838, 856, 875, 878, 898, 941, 951, 974, 982, 1010, 1034, 1036, 1087, 1093, 1116, 1119, 1122, 1126, 1140, 1141, 1157]. detection-based [450]. detector [414, 783, 1162]. determination [767]. deterministic [336, 731]. DETR [290, 783, 790, 898, 907]. DETRmg [307]. Deutsch [1047]. Development [573, 871, 1015, 1156]. device [9, 119, 593, 902]. devices [232, 392, 580, 664, 831, 871, 877, 921]. DevOps [1015]. dew [149]. dew-assisted [149]. DFGIC [765]. DFGIC-Net [765]. DFPD [1154]. DFPD-YOLO [1154]. DGN [130]. DGSL [1010]. Dhcache [400]. Dhcm [584]. diabetic [166]. DiabeticChain [166]. Diagnosis [111, 262, 410, 423, 462, 527, 634]. diagnosis-based [462]. diagnostic [327]. diagonal [576]. diagram [262]. diagrams [354]. Diamond [803, 998]. difference [242, 675]. differences [648]. differencing [704]. different [710, 806]. differential [49, 116, 326, 556, 577, 589, 596, 751, 770, 930, 935, 1076, 1142]. differentiating [213]. DiffNet [1111]. DiffREE [30]. Diffusion [30, 203, 208, 224, 338, 437, 714, 726, 765, 815, 828, 883, 1111]. diffusion-based [815, 828]. DiFuseR [21]. digit [23]. Digital [5, 229, 468, 519, 648, 748, 789, 915, 1017, 1063]. dilated [499, 855]. dimension [123, 230, 600, 645, 822]. dimensional [39, 136, 154, 212, 378, 495, 514, 600, 660, 680, 757, 767, 791, 883, 948, 1009, 1120]. dimensioning [893]. directed [751, 1001]. Directing [1057]. Directional [342, 834]. directions [773]. Dirichlet [118]. disassembly [7]. disaster [294, 538]. discard [408]. discharges [1]. discontinuous [857]. discover [597]. discovering [571, 693]. discovery [182, 927]. discrete [7, 159, 517, 606, 782]. discrete-time [159]. Discriminant [136, 397]. discrimination [712]. discriminative [925, 1057]. disease [402, 423, 783, 878, 1039, 1159]. disentangled [386]. disentanglement [528]. disjoint [48, 211, 273, 393, 744]. dispatch [897]. Distance [173, 458, 501, 572, 622, 638, 710, 767, 899, 1069]. distance-based [710]. distance-induced [572]. distance-vector [1069]. DistilBERT [438]. DistilBERT-based [438]. distillation [56, 509, 691, 788, 896, 973, 978, 1004, 1007, 1109, 1140, 1141, 1143]. distillation-based [973]. distillation-guided [1007]. distinguishing [309]. distracted [1052]. Distributed [11, 21, 57, 132, 228, 254, 315, 419, 434, 444, 578, 700, 766, 818, 844, 917, 919, 963, 977, 978, 1015, 1047]. distribution [195, 205, 257, 376, 594, 884, 1013, 1143]. distribution-aware [376]. Divergence [740]. diverse [392, 1160]. divide [393, 524]. divide-and-swap [393, 524]. divided [998]. DLC [771]. DLC-IDS [771]. DLTPPM [1046]. DML [706]. DNA [858]. DNN [745]. DNS [1067]. document [725]. documents [163, 171]. DoF [1014]. DoH [1067]. Domain [114, 139, 306, 376, 510, 572, 575, 595, 683, 695, 706, 778, 840, 867, 934, 943, 975, 991, 993, 1019, 1098, 1113, 1122]. Domain-adaptive [683, 943]. domains [249]. dominance [1079]. dominant [216]. dominating [501, 1070]. Double [76, 522, 987]. double-DQN [76]. down [876]. DPC [274]. DPC-empowered [274]. DPF [684]. DPF-Unet [684]. DPSMUNet [267]. DPU [859]. DPUSwap [859]. DQN [76, 523]. DQN-MSRA [523]. drawings [855]. drawn [828]. DRCD [473]. driven [37, 174, 268, 272, 348, 383, 465, 473, 556, 620, 641, 656, 696, 747, 804, 818, 852, 874, 933, 992, 1024, 1071, 1146]. Driving [275, 288, 718, 764, 889, 951, 1046, 1052, 1164]. DRL [213, 992]. DRL-driven [992]. drone [414, 607, 830, 1034]. drowning [1068]. drowsiness [453]. DRSS [58]. drug [693]. drugs [53, 479]. DRX [329]. DSC [433]. DSC-RepVGG [433]. DSDN [1166]. DSS [14]. DSSE [421]. DSTnet [754]. DT [1010]. DT-DGSL [1010]. DTCN [221]. Dual [8, 17, 19, 58, 68, 82, 127, 131, 157, 197, 267, 390, 400, 461, 464, 531, 591, 595, 622, 662, 706, 751, 754, 782, 847, 849, 868, 879, 896, 980, 1042, 1046, 1070, 1072, 1124, 1134, 1140, 1144]. dual-archive [751]. Dual-branch [127, 591, 1134, 1140]. dual-channel [461]. Dual-constraint [849]. dual-decoder [157]. dual-domain [595, 706]. dual-energy [8]. Dual-enhanced [531, 1042]. dual-evaluation-mode [1070]. dual-hash [400]. dual-layer [1046]. dual-modal [879]. dual-modality [847]. dual-path [17, 662, 1072]. dual-pooling [267]. dual-space [622]. dual-stage [782]. dual-strategy [1144]. Dual-stream [19, 68, 879]. dual-student-teacher [754]. Dual-YOLO [662]. dung [90]. dust [168]. DV [170]. DVFS [924, 1000]. DVFS-enabled [924]. DVTXAI [280]. dwarf [483]. dyed [239]. Dynamic [11, 18, 41, 44, 54, 63, 71, 77, 89, 159, 198, 341, 348, 371, 387, 389, 411, 440, 469, 481, 544, 574, 582, 584, 590, 594, 626, 682, 722, 730, 738, 749, 753, 759, 770, 807, 811, 853, 875, 896, 897, 907, 921, 979, 1010, 1032, 1037, 1043, 1053, 1113, 1121, 1125, 1147]. Dynamical [35, 939]. dynamics [731, 1148]. DYR [574]. DYR-SLAM [574].

e-commerce [229]. E-GRACL [42]. E-harnet [957]. E-MTMYOLO [1112]. EA [415]. each [296, 686]. eagle [251]. EAR [485]. Early [394, 476, 617]. early-modern [617]. earth [989]. earthquake [191, 294, 1042]. EAViz [881]. ECDOAGM [766]. echo [30, 595]. ECKT [870]. economic [897]. ecosystems [463, 604]. EDBLSD [359]. EDBLSD-IIoT [359]. Edge [15, 29, 35, 48, 76, 84, 100, 121, 164, 180, 192, 230, 246, 257, 270, 273, 278, 288, 310, 331, 334, 349, 356, 360, 436, 539, 553, 564, 580, 603, 611, 612, 639, 661, 664, 692, 729, 796, 825, 827, 924, 946, 992, 1008, 1012, 1044, 1062, 1074, 1075, 1128, 1144, 1166]. edge-cloud [564, 825]. edge-connectivity [180]. edge-DSDN-based [1166]. edge-independent [334]. edge/cloud [1008]. editable [800]. EDSDN [1166]. education [91, 882]. EEG [126, 405, 453, 592, 658, 878, 1098]. EEG-based [453]. effect [386]. Effective [6, 229, 288, 305, 493, 509, 520, 630, 1063]. effects [779]. efficiency [77, 109, 154, 184, 292, 484, 953]. Efficient [4, 8, 15, 23, 38, 40, 50, 61, 103, 111, 113, 116, 123, 140, 149, 195, 209, 215, 237, 252, 264, 274, 298, 321, 343, 354, 355, 408, 424, 434, 453, 482, 496, 509, 537, 542, 571, 592, 636, 690, 715, 717, 721, 732, 737, 742, 764, 777, 781, 783, 790, 810, 812, 822, 825, 831, 844– 846, 862, 867, 869, 898, 912, 930, 931, 944, 957, 972, 1008, 1012, 1014, 1052, 1058, 1060, 1078, 1080, 1085, 1101, 1120, 1143, 1144, 1148, 1152]. EfficientNetV2 [746, 990]. EfficientNetV2-based [990]. EFFN [221]. EFFN-Transformer [221]. EGAL [781]. EGAL-Net [781]. elastic [417, 578]. elasticity [666]. elasticnet [526]. ELDM [1166]. ELDM-EDSDN [1166]. election [426]. electric [241, 314, 475, 837]. Electrical [49, 918]. electromagnetic [432]. electronic [396, 816, 1092]. element [398]. elementary [619]. elephant [766]. Ellipse [569]. elliptic [458]. embedded [349]. embedding [218, 543, 1053]. EMD [1149]. EMD-based [1149]. emergency [294, 351, 447]. EMGODV [140]. EMGODV-Hop [140]. emission [612, 897]. emitter [1090]. emotion [31, 71, 560, 799, 890, 954]. emphysema [308].

Empirical [628, 851]. employing [541]. empowered [274]. Empowering [194]. Emsd [898]. Emsd-detr [898]. emulators [516]. enabled [195, 446, 469, 797, 827, 905, 924, 1078, 1135, 1159]. enabling [967]. encirclement [279]. encirclement-combat [279]. encoder [346, 470, 583, 644, 732, 861, 868]. encoder-decoder [732]. Encoders [940]. encoding [282, 792, 858]. encrypted [225, 835, 905]. encryption [12, 39, 63, 96, 108, 119, 181, 210, 454, 552, 651, 816, 817, 858, 883, 902, 916, 1027]. end [380, 429, 798]. end-to-end [380, 429, 798]. Energy [2, 4, 8, 15, 20, 52, 116, 123, 184, 209, 226, 237, 314, 359, 403, 427, 434, 460, 482, 484, 485, 542, 640, 647, 650, 742, 793, 845, 862, 918, 924, 953, 965, 976, 1008, 1012, 1085, 1095]. Energy-aware [2, 485, 924]. Energy-efficient [4, 8, 15, 116, 237, 434, 482, 542, 742, 845, 862, 1008, 1012, 1085]. Energy-harvesting-aware [226]. Energy-saving [976]. engine [259]. engineering [148, 188, 251, 368, 483, 533, 629, 884]. enhance [789, 857, 1089]. Enhanced [64, 75, 90, 110, 113, 117, 125, 140, 153, 155, 170, 220, 223, 300, 303, 319, 321, 327, 359, 370, 380, 388, 438, 449, 457, 458, 504, 513, 518, 531, 555, 574, 591, 608, 632, 674, 679, 705, 713, 716, 737, 755, 776, 782, 801, 805, 809, 826, 841, 866, 869, 872, 874, 876, 888, 898, 942, 944, 970, 982, 992, 1018, 1039, 1042, 1063, 1100, 1111, 1154]. enhancement [144, 160, 169, 244, 408, 413, 467, 507, 528, 734, 821, 840, 848, 856, 888, 934, 966, 1019, 1021, 1111]. Enhancing [10, 50, 72, 74, 77, 78, 178, 282, 292, 358, 394, 417, 538, 541, 560, 636, 668, 792, 814, 815, 845, 852, 853, 870, 885, 886, 906, 907, 928, 938, 946, 971, 1016, 1017, 1081, 1109, 1125, 1126, 1131]. ensemble [110, 178, 600, 628, 633, 665, 708, 730, 794, 809, 899, 955, 964, 979, 994]. entities [845]. entity [652, 1106, 1109]. entropy [65, 85, 167, 250, 1051]. entropy-based [65, 1051]. environment [142, 181, 237, 529, 573, 579, 588, 611, 641, 770, 825, 968, 1015]. environmental [54, 612]. environments [11, 223, 377, 474, 558, 569, 615, 630, 636, 774, 924, 995, 1128]. epilepsy [881]. epileptiform [1]. epochs [265]. equation [49, 589, 704]. equations [596, 625, 675, 731, 930]. equilibrium [708]. equipment [7]. erasure [486]. ERBFT [486]. error [9, 62, 109, 477, 703]. error-resilient [109]. ERW [786]. estate [118]. Estimating [386, 650, 961]. Estimation [270, 429, 458, 471, 759, 892, 1038, 1041, 1091]. Evaluating [103, 288, 349, 629]. evaluation [247, 385, 468, 586, 628, 665, 1070, 1105, 1149, 1160]. evasive [635, 1089]. event [126, 288, 402, 405, 742]. event-related [126, 405]. evidence [205, 1063]. evidential [1114]. evolution [116, 556, 577, 602, 751, 770, 933, 1076, 1142]. Evolutionary [54, 98, 388, 401, 440, 935, 942, 1011, 1118, 1138]. Evolutionary-enhanced [388]. examples [609]. Exaptation [1006]. exchange [1009, 1150]. exchangers [975]. excitation [1072]. execution [452, 774, 1080]. exit [745]. expansion [896]. expected [324]. expensive [1011]. Experience [214, 504, 789]. Experience-guided [214]. experimentation [212]. expert [1112]. expert-annotated [1112]. experts [407]. explainable [62, 280, 1112, 1151]. explanation [1087]. explicit [704]. exploitation [452]. Exploiting [183]. exploration [350, 722, 780, 939]. Exploring [227, 428, 681, 778, 1020]. Exponential [680, 704, 884]. Exponential-trigonometric [680]. Exponentially [204]. exponentiation [296]. exportable [465]. expression [238]. expressiveness [390]. extended [658, 1123]. extension [304, 398]. extensions [355]. external [281, 527]. Extra [164]. extracted [49]. extracting [299]. extraction [82, 240, 438, 620, 688, 775, 801, 870, 872, 896, 901, 970, 999, 1074]. extractive [171]. extrapolation [30]. ExtraTrees [665]. extreme [1131]. eye [1094]. eye-rubbing [1094]. eyes [1068].

F [647]. FaaS [1128]. FaasFlows [206]. Fabdb [753]. fabric [239, 369, 702, 754, 781]. face [119, 528]. faced [236]. Facial [454, 1084]. facilitated [304]. factor [456, 626]. factoring [102]. Factorization [79, 87, 102, 292, 627, 801, 1124]. factors [120, 281, 693, 1015]. failure [379]. fair [97, 372]. fake [177, 249, 299, 313, 791, 960]. fall [422, 792, 982]. families [948]. FANET [274]. Fast [100, 182, 252, 468, 580, 664, 977, 1011]. fastener [494]. faster [522]. FastGA [495]. FastMS [726]. FastMS-CDA [726]. Fault [262, 273, 316, 365, 462, 527, 564, 626, 672, 711, 753, 891, 1029, 1060, 1144, 1148]. fault-tolerance [891]. fault-tolerant [564, 626, 672, 753, 1060]. faults [48, 183, 597]. FCAFormer [993]. FDAG [840, 934]. FDAG-GAN [840, 934]. FDL [867]. FDL-YOLO [867]. FDPNet [631]. Feasibility [746]. feasible [286]. Feature [17, 19, 28, 30, 40, 64, 106, 129, 147, 169, 193, 212, 220, 238, 254–256, 285, 299, 358, 370, 406, 413, 419, 438, 489, 491, 503, 513, 532, 551, 592, 608, 632, 633, 636, 640, 645, 654, 657, 712, 735, 754, 757, 765, 769, 775, 788, 791, 794, 838, 840, 848, 864, 872, 885, 907, 913, 934, 937, 941, 947, 951, 952, 955, 967, 984, 987, 1032, 1051, 1059, 1082, 1084, 1087, 1099, 1100, 1124, 1137, 1155]. feature-attention [794]. feature-conditioned [30]. feature-guided [256, 765]. features [36, 49, 150, 306, 346, 381, 385, 526, 560, 609, 834, 845, 946, 975, 1057, 1074, 1104, 1155]. FedBat [137]. FedCK [1143]. federal [540]. Federated [26, 33, 63, 137, 158, 196, 226, 232, 246, 269, 317, 326, 378, 630, 661, 697, 707, 717, 747, 749, 773, 817, 863, 877, 959, 962, 983, 1037, 1075, 1115, 1143, 1153]. FedShield [1115]. FedSSU [717]. feedback [89, 762, 992]. feedforward [339]. Few [139, 583, 654, 870, 908, 1007, 1090, 1093, 1107, 1122]. Few-shot [139, 583, 654, 870, 908, 1007, 1090, 1093, 1107, 1122]. Fgcfl [1153]. Fibonacci [197, 675]. FIBS [468]. FIBS-fast [468]. field [431, 432, 465, 544, 573, 590]. fields [757, 889]. file [260]. Filling [998]. film [779]. filter [520, 821, 955, 1040, 1090]. filtering [79, 510, 520, 736, 1063]. financial [531, 1071]. find [1060]. Finding [36, 322, 356, 660]. fine [44, 50, 361, 409, 480, 622, 737, 1153]. fine-grained [44, 361, 409, 622, 1153]. fine-tuning [50, 737]. fingerprinting [748]. finite [626]. finite-time [626]. firefighting [462]. firewalls [614]. first [23]. fisher [709]. FIT4HPC [915]. fitness [440, 458]. fitness-distance-similarity [458]. fitting [230]. five [1, 39]. five-criterion [1]. five-dimensional [39]. flaws [1146]. flexible [293, 478, 552, 717]. flight [478]. flight-follow [478]. floating [477]. floating-point [477]. flock [655]. flock-inspired [655]. flow [129, 198, 281, 428, 434, 749, 785]. FlowLoader [1058]. flowshops [948]. flying [538]. FM [201]. FM/FM/1 [201]. FMO [709]. FO [1034]. FO-YOLO [1034]. focal [16, 501]. focus [1123]. focused [341]. focusing [758, 1154]. fog [52, 116, 377, 379, 426, 506, 755, 839, 1075, 1128]. fog-edge [1075]. foggy [604]. folded [320, 524, 900]. folded-concave [320]. follow [478]. FootprintNet [570]. footprints [570]. forecast [431]. forecasting [20, 49, 55, 68, 115, 281, 416, 488, 499, 566, 690, 798, 886, 928, 939, 986, 991, 993, 1083, 1095, 1139]. foreground [1057]. foreign [632, 1162]. forensic [748]. forensics [510]. forest [7]. forests [541]. forgery [631]. forging [732]. formal [738]. forward [874]. foundation [542]. Fourier [193, 416, 1028]. FP [885]. FP-RTDETR [885]. FPGA [121, 344, 453, 481, 774, 842, 892, 1148]. FPGA-accelerated [453]. FPGA-based [121, 892, 1148]. FPGAs [268]. FPN [1003]. fractal [600]. Fractional [90, 152, 416, 714, 874, 930]. frame [670, 759, 864, 1098]. frame-based [759]. framework [50, 69, 89, 101, 177, 234, 280, 312, 317, 331, 373, 423, 488, 521, 527, 585, 588, 600, 627, 644, 670, 707, 741, 792, 802, 815, 824, 837, 866, 896, 905, 960, 973, 978, 982, 1008, 1012, 1061, 1063– 1065, 1100, 1107, 1149, 1151, 1153, 1156, 1161]. frameworks [1020, 1031]. Fraud [659, 1115]. free [140, 509]. frequency [262, 342, 385, 510, 595, 778, 805, 826, 840, 867, 919, 934, 975, 991, 993, 1019, 1093]. frequency-domain [840, 934]. frequent [565]. friction [779]. friendly [881]. Frobenius [204]. frontier [443]. fronts [98]. Frozen [940]. FSOC [858]. FSSO [998]. fulfillment [1]. full [309, 762]. full-resolution [762]. fully [12, 296, 466, 1101]. fully-verifiable [296]. function [33, 90, 152, 320, 486, 561, 640, 945, 1131]. functional [552]. functions [619, 767]. fundus [1158]. Fusion [22, 28, 33, 43, 49, 51, 106, 129, 138, 158, 198– 200, 251, 262, 266, 306, 312, 345, 382, 387, 390, 395, 406, 407, 431, 467, 491, 545, 551, 591, 597, 640, 642, 645, 647, 656, 680, 684, 685, 735, 754, 765, 776, 788, 791, 838, 841, 855, 856, 861, 864, 876, 879, 885, 892, 913, 920, 925, 928, 940, 951, 954, 968, 978, 999, 1026, 1032, 1036, 1040, 1056, 1072, 1074, 1087, 1091, 1116, 1123, 1126, 1130, 1137, 1141]. fusion-based [920, 968, 1123]. future [36, 291, 773]. FuzOptRoute [1012]. fuzzing [780]. Fuzzy [29, 73, 103, 167, 173, 209, 294, 353, 402, 439, 626, 647, 696, 740, 837, 903, 1012, 1018, 1056, 1092, 1133]. fuzzy-based [626].

GA [339, 790]. GA-DE [339]. GA-DETR [790]. gain [318]. gait [373, 646]. Game [178, 228, 322, 401, 403, 487, 686]. game-based [322]. gaming [323]. GAN [739, 840, 888, 934, 1030, 1089, 1161]. GAN-assisted [1089]. GAN-CL-STO [1161]. gangue [968]. gap [994]. GAPPO [786]. GAPPO-ERW [786]. GARCH [366]. gate [573]. gated [823, 984]. gates [1028]. gating [82, 125]. Gaussian [205, 269]. gazelle [140]. GBADroid [491]. GCE [937]. GCE-Net [937]. GDNet [160]. GDRNet [499]. GDT [515]. GDT-IDS [515]. gearbox [829]. gem5 [887]. GEMM [601]. GEMV [496]. gene [238, 927]. general [101, 519, 563]. generalised [968]. generalization [114, 306, 535]. Generalized [102, 159, 217, 439, 675, 698, 1011]. generated [289, 417, 893]. generating [631, 635]. generation [33, 132, 138, 151, 352, 379, 587, 597, 691, 726, 741, 876, 916, 944, 1069, 1095]. Generative [239, 258, 289, 342, 888]. generator [94]. GenerCTC [101]. generic [1061]. genetic [88, 557, 596, 883, 929, 1016]. genome [355]. genotype [557]. geographical [103]. Geohash [85]. geolocation [223]. geometric [327, 947]. geophysical [704]. Geospectra [223]. gesture [534]. GFIDF [312]. GHG [365]. Ghost [61, 160]. Ghost-Weight [61]. gland [762]. global [90, 158, 326, 377, 483, 772, 819, 894, 1091]. global-local [1091]. GNN [215, 869]. GNN-based [215]. GNN-enhanced [869]. Golden [533]. good [26]. goore [322, 686]. government [800]. GPT [352]. GPT-based [352]. GPU [59, 270, 283, 287, 349, 361, 428, 452, 473, 537, 566, 734, 1064]. GPU-based [270, 349]. GPU-centric [537]. GPUs [21, 567, 601]. graceful [1085]. GRACL [42]. gradient [3, 55, 321, 514, 678, 781]. gradient-aware [3]. gradients [336]. gradual [312]. grain [899]. grained [44, 361, 409, 622, 1153]. granularity [46, 56, 122, 220, 1109]. granulator [511]. Graph [18, 27, 41–43, 76, 86, 107, 112, 125, 129, 138, 176, 193, 198, 215, 248, 293, 319, 341, 389, 406, 437, 470, 476, 488, 500, 508, 515, 531, 542, 572, 599, 627, 659, 672, 688, 691, 720, 798, 814, 835, 841, 875, 938, 961, 980, 1007, 1010, 1024, 1025, 1043, 1087, 1114, 1121, 1160]. Graph-augmented [1114]. graph-based [76, 508, 515, 672, 1025, 1160]. Graph-induced [43]. graphics [364]. Graphormer [798]. graphs [48, 120, 145, 218, 245, 309, 322, 356, 456, 543, 784, 1070]. graphyne [250]. GraSNet [1014]. grasp [1014]. grasping [952]. Grassmann [397]. gravity [1079]. gray [766, 1042]. greedy [182]. green [1156]. grey [819, 877]. grids [709, 850]. Ground [276, 424]. Ground-based [424]. group [199, 227, 831, 927, 983]. group-wise [983]. grouped [601]. grouping [499]. groups [197]. Growth [719]. GRU [462, 671, 722]. GRU-transformer [671]. GSTT [1134]. GSW [222]. GUFORMER [3]. guidance [983, 1108]. guided [80, 168, 214, 256, 348, 358, 367, 375, 470, 705, 765, 809, 840, 895, 934, 946, 1007, 1036, 1122, 1142]. GUITARES [1025]. Gumbel [33]. GWO [184, 1125].

H.266 [1053]. H.266/VVC [1053]. Hadoop [641]. Hamacher [294]. hand [828, 1038]. hand-drawn [828]. handling [294]. handover [426, 593]. HAR [956]. hardening [731]. hardware [23, 252, 344, 794]. harmonic [439, 837, 1018]. Harnessing [901]. harnet [957]. Harris [148, 606, 643]. harvesting [8, 226]. harvesting-assisted [8]. hash [400]. hate [28]. hawks [148, 606, 643]. hazy [505]. head [2, 31]. headlines [1136]. health [1092]. Healthcare [141, 181, 285, 301, 585, 797, 877, 1013]. heart [801]. heat [975]. Hebei [351, 447]. height [1163]. helmet [727]. herd [597]. heterogeneity [196, 1153]. heterogeneous [11, 38, 41, 112, 167, 268, 279, 283, 328, 333, 427, 434, 473, 481, 508, 517, 557, 588, 615, 618, 628, 633, 694, 705, 769, 893, 911, 924, 962, 969, 1051, 1087, 1125, 1128, 1152]. heuristic [257, 339, 374, 719]. heuristics [374]. hexagonal [222]. HFSL [196]. HGFT [947]. HIDE [1053]. hiding [210]. hierarchical [195, 358, 607, 661, 801, 820, 823, 962, 967]. Hierarchy [584, 1015, 1022]. HIFNet [713]. High [81, 108, 154, 224, 260, 275, 277, 318, 344, 347, 398, 451, 482, 495, 514, 539, 563, 571, 602, 609, 648, 714, 761, 763, 767, 854, 882, 902, 904, 919, 948, 949, 967, 988, 1053, 1103, 1119]. High-accuracy [967, 988]. high-accurate [347]. high-altitude [854]. high-dimensional [154, 514, 767, 948]. high-efficiency [154]. high-frequency [919]. high-integrity [1053]. high-level [451, 609]. high-performance [260, 344, 539, 602, 648, 763, 882, 1103]. High-precision [277, 761, 1119]. high-resolution [904]. high-security [108]. High-speed [482, 563]. High-utility [81]. higher [203]. higher-order [203]. highly [61, 844]. Hilbert [592]. histopathological [395]. HMAC [522]. holes [382]. holographic [1020]. homogeneous [1160]. homomorphic [12, 63, 817]. honey [181]. honey-based [181]. Hop [140, 170]. horse [188, 597]. hot [88]. HPC [194, 427, 445, 820, 1023]. HSVDetector [508]. HTTPS [1067]. Huber [1040]. Human [12, 37, 170, 342, 471, 577, 676, 956, 957, 982, 1065]. Human-inspired [1065]. human-object [676]. human-powered [577]. hummingbird [110, 950]. HW [842]. HW-SW [842]. Hybrid [13, 19, 20, 68, 72, 83, 144, 210, 254, 317, 359, 373, 381, 419, 432, 448, 452, 453, 455, 462, 464, 482, 483, 487, 521, 624, 628, 658, 690, 705, 712, 714, 715, 722, 732, 739, 761, 816, 819, 821, 824, 829, 903, 947, 948, 957, 966, 969, 972, 975, 987, 1005, 1030, 1042, 1067, 1076, 1083, 1094, 1095, 1120, 1158]. HydraGNN [542]. hydraulic [185]. hydrological [103]. hydropower [7]. hyper [374]. hyper-heuristic [374]. hyperchaotic [39, 94, 454, 858, 1027]. hypercube [273, 803, 900]. Hypergraph [613, 954]. Hypergraph-based [954]. hyperledger [369, 702]. hyperparameter [777]. hyperspectral [193, 320, 507, 841, 1007]. hypothesis [133].

I/O [616]. IaaS [994]. iAFTA [891].

iAgent [891]. ICAT [191]. ICAT-net [191]. ICN [436]. identification [1, 56, 220, 363, 382, 513, 547, 570, 849, 1137]. Identifying [263, 503, 871, 1001, 1015, 1079]. identity [695, 702]. IDS [254, 419, 515, 771]. Ifqa [1071]. Ifqa-llm [1071]. II [562, 760]. IID [63, 630, 739, 863, 1030]. IIoT [359, 1127]. IIoT-based [1127]. illumination [946]. illumination-guided [946]. Im2col [1045]. Im2col-based [1045]. Image [3, 17, 26, 70, 90, 96, 100, 105, 108, 113, 138, 160, 168, 169, 193, 240, 244, 275, 331, 336, 338, 346, 367, 376, 395, 397, 431, 454, 480, 490, 553, 591, 603, 654, 662, 734, 750, 762, 782, 824, 828, 840– 842, 854, 861, 883, 888, 907, 914, 916, 926, 934, 937, 966, 978, 1007, 1019, 1027, 1032, 1040, 1056, 1072, 1074, 1111, 1123, 1156]. image-set [397]. image-text [138, 367]. imagery [202, 303, 384, 658, 1034]. images [22, 113, 267, 410, 505, 605, 631, 634, 684, 706, 732, 776, 809, 836, 858, 898, 904, 1091, 1098, 1104, 1157, 1158]. imaging [423, 534, 553, 603]. imbalance [16, 518, 730]. imbalanced [248, 443, 659, 979]. imitation [289, 786]. Impact [183, 187, 265, 503, 648, 742]. imperceptibility [815]. implementation [398, 496, 567, 625, 1088]. implementations [812]. Implementing [1067]. Implicit [102, 687, 914]. Improve [57, 156, 399, 647, 736, 901, 1086]. Improved [34, 44, 66, 69, 116, 132, 177, 188, 230, 238, 251, 261, 266, 274–276, 290, 295, 323, 356, 382, 403, 434, 454, 471, 483, 486, 492, 497, 505, 558, 565, 582, 606, 624, 643, 663, 696, 770, 828, 848, 884, 889, 929, 935, 974, 981, 988, 996, 1042, 1055, 1117, 1119, 1166]. improvement [306, 906]. Improving [157, 262, 435, 609, 853, 963, 964]. imputation [1092]. in-context [594]. in-depth [910]. in-memory [190]. In-situ [225]. InceptionNet [346]. incompatible [948]. incorporating [52, 527, 1018]. increased [797]. increasing [2]. incremental [81, 1043]. independent [334, 530, 770, 918]. index [712, 961]. indexing [321]. indicator [1118]. indices [250, 324, 479]. indirect [1150]. indispensable [207]. indistinguishability [369]. individual [348, 743]. individual-based [743]. indoor [665, 863]. induced [43, 145, 572]. inductive [357]. Industrial [84, 359, 415, 681, 732, 905, 910, 975, 1002]. industry [950, 1078]. InEPS [427]. infant [106]. Inference [60, 73, 183, 205, 209, 358, 496, 545, 589, 745, 760, 810, 904]. infidelity [977]. infinite [859]. influence [21, 130, 556]. influential [503, 1001]. information [43, 56, 158, 255, 285, 357, 387, 467, 537, 559, 633, 671, 699, 722, 765, 769, 785, 856, 883, 1051, 1096, 1116, 1133]. informed [92, 394]. InformerAutoScale [923]. infotainment [748]. Infrared [22, 56, 220, 765, 790, 849, 885, 978, 1035, 1036, 1056]. infrastructure [446, 462, 561, 576]. infrastructures [15, 776]. inheritance [158, 577]. initialization [705]. injection [532]. Innovations [340]. Innovative [20, 72, 379, 526, 819]. input [739, 1030]. Inqasm [442]. InQuIR [442]. insider [246]. insights [851, 938]. insole [646]. inspection [549]. inspired [107, 655, 772, 1013, 1042, 1065, 1135]. instance [1117, 1165]. instance-based [1165]. Instruction [398, 667]. instrument [823]. Integrated [95, 198, 201, 300, 339, 532, 796, 1004, 1012, 1073, 1151]. integrating [379, 749, 769, 807, 943, 1050, 1077, 1081, 1095, 1128]. integration [351, 447, 668, 865, 880, 907, 937]. Integrative [1039]. integrity [522, 1053]. Intelligence [62, 270, 882, 992, 1062]. Intelligent [45, 76, 237, 353, 362, 402, 427, 462, 593, 637, 820, 833, 1002, 1068, 1071, 1081, 1127]. intensity [54]. intent [312, 685, 913, 1066]. intent-based [685]. intention [1071]. intention-driven [1071]. Inter [283, 598, 989]. inter-bot [598]. inter-GPU [283]. Inter-satellite [989]. Interaction [548, 676, 699, 795, 799, 855, 1148]. interactions [693]. Interactive [189, 808, 1140]. interconnected [112]. interconnection [525, 672, 803]. interdependencies [668]. Interest [122, 146, 221]. interference [104]. interictal [1]. intermediate [418]. internally [211]. Internet [84, 123, 271, 301, 332, 547, 600, 610, 653, 749, 773, 845, 905, 1087]. interoperability [797]. interpolating [660]. Interpretable [609, 890, 927]. interruption [201]. intersample [474]. intersections [29]. intra [367]. intra-modal [367]. Introducing [295, 468]. introspection [260]. introspective [691]. intrusion [16, 42, 110, 231, 254, 277, 419, 515, 541, 636, 645, 708, 771, 955, 1016, 1099, 1161]. intrusions [632]. invariant [306]. invasive [669]. Inverse [159, 697]. Investigating [187]. Investigation [704]. IoE [604]. IoMT [219, 271, 636, 695, 916]. ion [973]. IoT [6, 42, 61, 110, 133, 141, 149, 151, 195, 209, 231, 232, 246, 272, 359, 360, 362, 383, 392, 435, 446, 463, 529, 545, 598, 628, 651, 755, 766, 780, 793, 818, 871, 877, 906, 921, 933, 1010, 1016, 1020, 1029, 1099, 1102, 1125, 1128]. IoT-based [6, 209, 435]. IoT-cloud [141]. IoT-enabled [195, 446]. IPAQ [377]. IPFS [797]. irregular [98, 490]. IRS [576]. IRS-assisted [576]. isogeny [468]. isogeny-based [468]. isolation [768]. item [826]. itemsets [565]. iterated [182]. iteration [186, 930]. iterative [520]. iterative-partitioning [520]. iterator [343].

Jackal [533]. Jacobsthal [197]. Japanese [617]. Jaya [434]. jitter [460]. job [293, 361, 478, 606, 820, 948]. job-shop [293]. jobs [361]. Joint [84, 459, 576, 759, 787, 838, 999]. jointly [46]. Josephus [180, 334]. Jozsa [1047]. judgment [420]. juggling [677].

K-NN [66]. Kacformer [824]. KAN [872, 1024]. Kds [978]. Kds-radfnet [978]. KdV [704]. Kepler [66]. keratoconus [1094]. kernel [24, 70, 679]. kernel-based [24]. kernels [728, 893]. Key [6, 9, 57, 149, 225, 259, 343, 371, 382, 400, 568, 673, 817, 905, 919, 1009, 1015, 1150]. key-derivation [225]. key-value [57, 259]. keyframe [408]. keyphrase [688]. keystroke [1130]. kinds [806]. Kirchhoff [961]. knobs [844]. Knowledge [33, 131, 158, 258, 407, 509, 577, 689, 737, 747, 788, 814, 824, 870, 896, 938, 973, 978, 1007, 1025, 1109, 1138, 1140, 1143]. Knowledge-driven [747]. knowledge-enhanced [737]. Kolmogorov [1135]. Kolmogorov-Arnold [1135]. Korean [163, 843]. Kronecker [874]. KSIPF [520]. KubePipe [451]. Kumaraswamy [760]. KV [156].

label [122, 594, 608, 1051, 1122, 1143]. label-aware [122]. labeled [712, 1051]. labeling [834]. Lagrange [425]. landmarks [480]. landscape [347, 350, 738]. Language [50, 75, 163, 194, 258, 352, 496, 530, 737, 805, 843, 847, 895, 925, 960, 1049, 1058, 1071, 1137]. Large [41, 50, 70, 91, 104, 145, 163, 194, 496, 514, 561, 565, 737, 843, 895, 960, 1014, 1049, 1058, 1070, 1071]. Large-scale [41, 91, 104, 163, 514, 561, 565, 843, 1014]. latency [359, 733, 753, 1086]. latent [118, 168, 815]. latent-content [168]. lattice [219]. lattice-based [219]. layer [31, 322, 382, 506, 545, 728, 984, 1046, 1079, 1105]. layer-wise [984, 1105]. layered [846, 1005, 1082]. layerwise [811]. layout [333, 1045]. Ldstd [414]. leader [478]. leadership [616, 771]. leakage [165]. leakage-resilient [165]. learn [207]. Learnable [572, 656, 784]. learned [321]. Learning [9, 33, 37, 58, 63, 69, 72, 74, 113, 114, 139, 153, 158, 161, 178, 185, 200, 207, 220, 241, 246, 254, 263, 271, 284, 288, 289, 293, 314, 315, 317, 320, 325, 332, 337, 361, 370, 378, 394, 410, 419, 420, 422, 423, 448, 462, 485, 545, 555, 556, 560, 564, 594, 596–598, 614, 628, 630, 639, 644, 659, 669, 678, 685, 686, 717, 743, 747, 763, 773, 785, 786, 814, 825, 827, 847, 852, 857, 863, 867, 873, 877, 880, 882, 926, 953, 955, 964, 971, 983, 995, 1007, 1017, 1037, 1039, 1065, 1072, 1073, 1107, 1123, 1131, 1143, 1147, 1152, 1155]. learning [1, 20, 35, 92, 110, 130, 137, 148, 176, 184, 188, 196, 213, 223, 232, 256, 265, 269, 286, 295, 308, 326, 328, 373, 386, 388, 389, 438, 451, 471, 475, 478, 490, 514, 523, 575, 583, 589, 600, 613, 634, 661, 676, 708, 741, 749, 782, 817, 820, 844, 869, 875, 881, 901, 908, 914, 917, 935, 959, 962, 977, 986, 1010, 1022, 1024, 1044, 1067, 1075, 1087, 1115, 1145, 1153]. learning- [614]. learning-based [35, 113, 184, 241, 475, 485, 614, 773, 827, 881, 1037, 1145]. learning-driven [556]. least [210]. LEDDR [1052]. LEDDR-YOLO [1052]. Leech [719]. legacy [484]. legal [420]. Leipnik [152]. lemma [945]. length [652, 744, 1048]. length-representation-based [652]. lesion [327]. less [1145]. leukemia [410]. level [46, 125, 183, 189, 200, 256, 318, 346, 374, 409, 444, 451, 529, 599, 609, 668, 903, 1138]. leveling [185]. Levenberg [55]. Leveraging [1, 223, 272, 383, 426, 604, 631, 682, 708, 789, 826]. LEVYEFO [432]. LEVYEFO-WTMTOA [432]. LG [784]. LGASR [168]. library [283, 981]. LiDAR [507, 642]. LIF [398]. lifecycle [973]. lifetime [2, 853]. light [27, 160, 275, 333, 765, 946, 1021]. LightGBM [665]. LightUAV [105]. LightUAV-YOLO [105]. Lightweight [56, 105, 191, 199, 207, 224, 233, 261, 301, 305, 307, 365, 392, 493, 527, 546, 558, 580, 608, 664, 721, 727, 732, 781, 788, 790, 832, 850, 885, 886, 982, 1003, 1038, 1052, 1069, 1103, 1141, 1156]. lightweight-based [1038]. like [48, 352]. Likelihood [760]. limit [663]. limitations [629]. limited [822, 831]. limited-capacity [822]. line [549]. linear [136, 218, 318, 356, 425, 675]. Linguistic [103, 353, 381, 880]. link [357, 627, 989]. linked [1147]. lite [297, 404]. literature [253, 747, 1023, 1147]. LiteYOLO [365]. LiteYOLO-GHG [365]. lithium [973]. lithium-ion [973]. livestreaming [789]. llm [1071]. LLMs [886]. LM [55]. LMF [1021]. LMF-Net [1021]. load [49, 52, 55, 257, 328, 469, 643, 874, 893, 929, 963, 972, 1098]. load-balancing [929]. loader [1058]. LOADng [1069]. local [132, 158, 182, 302, 326, 668, 1058, 1070, 1091, 1111]. local-search [1070]. locality [397, 677]. localization [1, 140, 170, 195, 249, 311, 471, 863, 990]. Locally [899]. Location [85, 128, 221, 369, 571, 837, 1022]. lock [206, 689]. lock-in [206]. locking [1145]. LOFs [794]. log [375, 448]. log-based [448]. LogCTBL [448]. logic [696, 1012, 1145]. logic-integrated [1012]. logistic [679]. logistics [1029, 1054]. LogSD [375]. long [68, 115, 117, 407, 499, 716, 851, 928, 963, 1042, 1083]. long-tailed [407]. long-term [68, 115, 499, 928, 1083]. loop [67]. LoRA [50, 371, 737]. LoRaWAN [371]. loss [16, 33, 217, 350, 638, 971, 1137]. lossy [392]. Lotka [731]. Low [6, 160, 318, 346, 374, 414, 509, 579, 733, 753, 810, 946, 989, 1021]. low-altitude [414]. Low-cost [810]. low-earth-orbit [989]. low-latency [733, 753]. low-level [346, 374]. low-light [160, 946, 1021]. low-power [6]. low-visibility [579]. LPQAA [233]. LPTMono [1091]. LSA [1105]. LSA-MEP [1105]. LSM [57, 343]. LSM-tree-based [57, 343]. LSPP [165]. LSTM [31, 142, 177, 265, 366, 403, 453, 532, 587, 628, 658, 807, 821, 836, 1066]. LSTM-based [142]. LU [87]. lung [331]. LVAST [305]. Lynxsight [710].

M [502, 507]. M-C&M-BL [502]. MA [1033]. MA-MSTNet [1033]. MAC [190]. MACAE [231]. Machine [1, 20, 69, 78, 110, 184, 217, 237, 254, 260, 263, 271, 314, 332, 370, 394, 419, 422, 451, 475, 514, 597, 614, 669, 763, 880, 977, 1039, 1131, 1145]. machines [469]. MAFFNet [735]. MAGDM [740]. maintaining [91]. maintenance [7]. maize [783]. majority [959]. makeup [1084]. making [29, 353, 657, 837, 866, 1073]. Malgta [895]. malicious [438, 583, 702, 1067]. malware [151, 491, 635, 782, 895]. mamba [235, 449, 690, 778, 1021]. Mamba-based [1021]. Mamba-Enhanced [449]. management [103, 166, 232, 251, 278, 317, 371, 593, 673, 797, 839, 905, 964, 973, 976]. managing [766]. mandibular [1112]. manifold [397, 1166]. manipulation [952]. manipulators [626]. Mann [987]. mantis [709]. manufacturing [135]. Many [80, 124, 440, 822, 1118]. Many-objective [80, 124, 440, 1118]. many-to-many [822]. map [431, 566, 574]. MAPER [15]. mapping [193, 521, 594, 604]. MAPPO [457]. MapsTSF [690]. MAQT [429]. marine [819, 1092]. Maritime [22]. marketing [229]. Markov [89]. Marmara [265]. Marquardt [55]. mash [285]. mash-up [285]. mask [96, 876, 1041]. Masked [239, 583]. masking [490]. massive [571, 766]. matching [24, 138, 355, 367, 572, 822, 941]. MATD3 [279]. material [351, 447]. materials [13, 484, 542]. mathematical [201, 340, 619]. mating [440]. MATLAB [479]. matrices [159]. matrix [79, 175, 204, 214, 292, 496, 627, 801, 1051]. MAVs [481]. max [47]. max-min [47]. maximization [21, 130, 146, 357, 556]. maximum [356, 744]. Mayfly [619]. MC [153]. MC-CRS [153]. MCAN [299]. MCDM [657]. MCGDM [173]. Mdcsnet [387]. MDH [19]. MDH-Net [19]. MDS [982]. MDS-OpenPose [982]. mean [837]. means [65, 274, 439, 520, 707]. measure [204, 740]. measurement [230, 569, 1079]. measures [173, 250, 651, 1031]. MEC [459]. MECG [319]. mechanism [18, 28, 31, 40, 82, 104, 124, 126, 331, 336, 401, 405, 438, 440, 584, 622, 656, 807, 952, 1006, 1050, 1059, 1078, 1083, 1159]. mechanism-based [1083]. mechanisms [191, 282, 638, 943]. media [272, 284, 383, 1106]. Median [510]. Medical [26, 39, 70, 90, 96, 111, 271, 308, 317, 378, 396, 480, 547, 553, 591, 603, 861, 916, 926, 1074, 1156]. medium [475]. medium-term [475]. MEDU [762]. MEDU-Net [762]. mega [989]. mega-constellation [989]. MeghMesa [666]. MEKF [407]. membership [1018]. memory [115, 117, 189, 190, 231, 509, 582, 584, 635, 677, 963, 1042, 1064, 1098]. memory-based [582, 963]. memristive [204, 425]. memristor [1101]. Menger [180]. meniscus [526]. MEP [1105]. MERGE [809]. merging [667, 724, 1049]. meshes [909]. mesoscale [513]. message [357]. Meta [207, 339, 350, 358, 639, 719, 827, 852, 908, 1147]. meta-heuristic [339, 719]. meta-learning [207, 908, 1147]. Meta-path-driven [852]. Meta-path-guided [358]. Meta-reinforcement [827]. metaheuristic [201, 236, 623, 772, 777, 998]. metaheuristics [1147]. metal [40, 155, 788]. metaphor [467]. Method [2, 11, 62, 83, 103, 113, 221, 238, 249, 255, 262, 264, 330, 350, 351, 370, 382, 385, 411, 425, 431, 444, 447, 476, 487, 490, 491, 508, 520, 523, 558, 562, 563, 570, 581, 597, 617, 620, 640, 642, 678, 696, 701, 704, 714, 718, 728, 730, 740, 749, 777, 788, 790, 819, 827, 837, 844, 848, 854, 855, 889, 908, 911, 943, 953, 962, 968, 988, 1004, 1032, 1042, 1052, 1059, 1079, 1121, 1123, 1134, 1145, 1165, 1166]. methodology [465, 1002, 1055]. methods [204, 315, 399, 543, 1015]. metric [66, 903, 1105]. Mfe [513]. Mfe-net [513]. MFFCNN [416]. MFS [608]. MFTFF [864]. MGFormer [671]. MGSF [494]. MHAED [732]. MHAED-Net [732]. micro [988]. micro-milling [988]. microscope [758]. microscopic [410]. microservices [104]. mighty [886]. migration [184, 755, 758, 992]. migrations [15]. military [1056]. Millimeter [270, 642]. millimeter-wave [642]. milling [988]. mills [88]. MIMO [270]. mine [51]. miniature [720]. minimal [225]. minimization [306]. minimize [606]. minimizing [184, 1061]. Minimum [218, 501, 563, 1070]. mining [56, 81, 566, 911]. Mink [1014]. Mink-GraSNet [1014]. minority [248, 701]. MIoT [547]. mirroring [660]. missile [374]. missing [1092]. mitigate [183]. Mitigating [104, 446, 540, 1063]. mitigation [474, 547]. Mixed [87, 190, 214, 1033, 1085]. mixed-criticality [1085]. Mixed-precision [87, 214]. mixed-signal [190]. mixer [702]. MixeRF [757]. mixing [757]. mixture [13]. Mixup [375]. MLDDoS [444]. MLExchange [904]. MLP [795]. MLS [521]. MLTree [846]. Mobile [9, 35, 119, 123, 128, 257, 286, 433, 457, 639, 682, 702, 827, 992, 1008, 1012, 1044, 1144]. mobility [15, 1029, 1127]. mobility-aware [15, 1029]. mobility-support [1127]. modal [107, 127, 177, 367, 642, 670, 735, 805, 879]. modality [86, 319, 847, 849, 876, 1074]. modality-enhanced [319]. modality-squeeze [86]. mode [1070]. Model [30, 34, 39, 40, 65, 75, 105, 106, 143, 162, 163, 178, 179, 200, 201, 222, 228, 237, 254, 265, 273, 304, 307, 314, 337, 345, 363, 366, 378, 381, 388, 394, 396, 398, 411, 415, 419, 437, 448, 474, 495, 502, 512, 539, 544, 546, 548, 580, 587, 590, 618, 633, 634, 641, 652, 664, 714, 715, 726, 733, 746, 759, 766, 771, 795, 799, 800, 807, 836, 843, 857, 862, 868, 871, 880, 890, 895, 898, 923, 939, 940, 960, 975, 995, 1003, 1005, 1029, 1035, 1041, 1046, 1048, 1050, 1054, 1055, 1057, 1058, 1082, 1103, 1113, 1133]. model-agnostic [1048]. model-based [895]. Modeling [46, 122, 171, 247, 300, 340, 367, 412, 479, 484, 542, 690, 716, 738, 772, 1026]. modelling [650, 731]. models [20, 50, 187, 194, 258, 288, 338, 352, 356, 493, 496, 535, 542, 737, 761, 805, 1017, 1049, 1071, 1080, 1111, 1160]. modern [268, 617]. modification [1018]. modified [99, 164, 186, 746]. modular [296]. modulation [678, 735]. module [231, 267]. module-assisted [231]. modules [493, 558]. moduli [102]. MOEAD [705]. molar [1112]. momentum [1107]. Mondrian [541]. mongoose [483]. monitoring [260, 332, 435, 511, 612, 669, 702, 988, 1094]. monocular [449, 1091]. Monomm [449]. Monte [374, 562]. morphological [1]. most [23]. Motif [729]. Motif-based [729]. motion [297, 404, 1077]. motor [658]. mountain [140]. move [182]. movements [106]. movie [78, 347]. moving [385]. MPD [852]. MPD-RL [852]. MQTT [521]. MR [634]. MRI [19, 345, 423, 684, 1132]. MSA [956]. MSA-HAR [956]. MSADBO [1095]. MSBES [251]. MSF [1026]. MSF-CSCNet [1026]. MSRA [523]. MSS [621]. MST [86]. MST-ARGCN [86]. MSTF [928]. MSTNet [1033]. MT [38]. MT-3000 [38]. MTMYOLO [1112]. Multi [4, 7, 10, 19, 31, 46, 70, 74, 84, 88, 89, 96, 98, 107, 108, 113, 116, 122, 124, 135, 153, 156, 177, 185, 186, 216, 220, 255, 293, 314, 322, 348, 372, 377, 381, 409, 412, 416, 429, 431, 436, 458, 487, 491, 502, 559, 576, 599, 607, 621, 630– 632, 635, 638, 640, 642– 644, 657, 670, 671, 694, 720, 725, 735, 745, 751, 755, 757, 774, 783, 788, 791, 805, 808, 837, 838, 843, 861, 862, 864, 872, 885, 913, 926, 937, 946, 952, 956, 967, 981, 983, 996, 997, 1026– 1028, 1037, 1044, 1074, 1077, 1096– 1098, 1102, 1105, 1109, 1133, 1137, 1139, 1155]. multi [17, 32, 40, 49, 54, 56, 59, 60, 199, 251, 266, 277, 279, 287, 387, 432, 444, 449, 455, 459, 507, 533, 552, 554, 557, 608, 680, 705, 718, 722, 723, 728, 741, 784, 817, 844, 856, 884, 892, 897, 901, 928, 935, 936, 941, 942, 948, 951, 969, 1022, 1032, 1033, 1056, 1058, 1059, 1120, 1126, 1129, 1141, 1151]. multi- [251]. multi-access [436]. multi-agent [279, 381, 784, 844, 872]. Multi-attributive [1133]. multi-BiLSTM [502]. multi-branch [507, 718]. multi-carrier [576]. multi-cell [459]. multi-channel [455]. multi-class [74, 630]. multi-classifier-assisted [1129]. multi-client [552]. multi-CNN [502]. Multi-CNNs [1096]. multi-compactions [156]. multi-contrastive [153]. Multi-controlled [1028]. multi-core [59, 287]. multi-criteria [89, 657, 837]. Multi-cue [1077]. multi-cylinder [185]. Multi-dimensional [757, 791]. Multi-document [725]. multi-exit [745]. multi-feature [640, 838]. multi-frame [864, 1098]. multi-fuzzy [1056]. multi-genotype [557]. Multi-granularity [46, 56, 122, 220, 1109]. multi-head [31]. multi-hierarchical [607]. multi-information [722]. multi-information-based [671]. multi-key [817]. multi-label [608]. multi-layer [322, 728]. Multi-level [409, 444, 599]. multi-mechanism [124]. Multi-medical [96]. multi-metric [1105]. Multi-modal [107, 177, 642, 670, 735, 805]. Multi-modality [1074]. multi-model [314]. Multi-object [838, 941, 997, 1077]. Multi-objective [54, 59, 88, 98, 116, 255, 287, 293, 348, 377, 608, 643, 694, 705, 722, 723, 751, 755, 897, 942, 969, 1155]. multi-party [372, 1120]. Multi-path [913]. multi-person [7]. multi-piston [723]. multi-process [135]. Multi-purpose [843]. Multi-queue-based [4]. Multi-robot [487]. multi-round [1059]. Multi-scale [10, 17, 32, 40, 49, 70, 199, 277, 387, 416, 429, 431, 449, 507, 631, 632, 638, 783, 788, 808, 856, 861, 885, 892, 926, 928, 936, 937, 946, 951, 952, 956, 967, 1026, 1032, 1033, 1074, 1097, 1126, 1137, 1139, 1141]. multi-scenario [554]. multi-secret [621]. multi-sensor [1102]. multi-server [84]. multi-sources [635]. Multi-stage [19, 983]. multi-static [862]. multi-step [186]. Multi-strategy [266, 458, 533, 680, 884, 948, 981, 996]. Multi-stream [559]. Multi-task [644, 741, 901, 952, 1022, 1044, 1151]. multi-tasking [935]. multi-tenant [774]. multi-threshold [113]. multi-tier [1058]. Multi-time-scale [412]. multi-TPU [60]. multi-tracker [432]. multi-UAV [1037]. Multi-user [745]. multi-valued [216]. multi-verse [897]. Multi-view [491, 720]. multi-wing [108, 1027]. multichannel [200]. multiclass [423, 658, 871]. multicore [536, 912, 999]. multicores [887]. multidimensional [855]. multidisk [259]. multilayer [346]. multilayered [695]. multilevel [90, 666, 1004]. multilingual [960]. Multimodal [28, 46, 58, 64, 71, 86, 284, 299, 319, 467, 548, 622, 644, 678, 751, 791, 792, 795, 799, 804, 809, 876, 881, 890, 940, 954, 980, 1109, 1142]. multimodal-enhanced [809]. multinational [433]. multiobjective [622]. multiparty [497]. multiple [36, 39, 72, 83, 249, 279, 298, 407, 541, 686, 823, 831, 846, 883, 1076]. multiple-color [883]. multiplex [729, 1079]. multiplication [214, 1120]. multiplier [109]. Multiscale [513, 716, 732, 920]. multisense [647]. multisensor [647]. multistage [726]. multistep [523]. multitask [1138]. Multivariate [162, 282, 416, 488, 875, 955, 993, 1024, 1055, 1087, 1134]. multizone [517]. music [581, 823]. mutation [1118, 1135]. Mutual [242, 255, 301, 357, 691]. MVFL [817]. MVS [364].

N [689]. N-Lock [689]. named [652, 1106, 1109]. Namib [403]. nanotechnology [253]. Narwal [1100]. native [923]. natural [248, 352]. navigation [107, 409, 604, 866]. NBO [403]. NCA [1104]. nearest [23, 79, 633, 1051]. negative [79, 138, 801]. negotiation [228]. Neighbor [79, 309, 1051]. neighborhood [248, 503, 633, 769, 1142]. neighbors [23, 772]. NER [857]. nested [652]. Net [19, 70, 191, 513, 591, 732, 735, 762, 765, 776, 781, 926, 937, 1021]. NetQASM [442]. nets [247]. Network [2, 3, 16, 17, 19, 68, 82, 86, 100, 125, 129, 133, 139, 144, 152, 157, 160, 182, 191, 212, 233, 242, 250, 267, 291, 293, 298, 308, 325, 327, 333, 366, 395, 403, 409, 412, 416, 437, 449, 455, 461, 485, 499, 513, 541, 551, 555, 570, 591, 611, 635, 645, 658, 671, 674, 683, 684, 687, 699, 720, 728, 732, 752–754, 765, 768, 781, 787, 803, 804, 810, 834, 835, 841, 845, 849, 853, 862, 867, 871, 874, 888, 905, 908, 939, 952, 955, 957, 966, 973, 975, 980, 995, 1021, 1025, 1026, 1033, 1081, 1086, 1090, 1101, 1117, 1121, 1130, 1135, 1139, 1159]. network [27, 36, 56, 71, 239, 275, 299, 311, 341, 398, 408, 457, 481, 488, 504, 507, 515, 554, 559, 649, 655, 706, 808, 869, 879, 920, 925, 927, 984, 1008, 1024, 1036, 1132, 1140, 1158]. network-assisted [308]. network-based [835, 1121]. network-driven [874]. network-on-chip [753, 1086]. networked [538]. networking [426, 685]. networks [6, 8, 18, 25, 36, 42, 64, 97, 127, 131, 145, 146, 164, 170, 175, 180, 183, 189, 195, 204, 209, 211, 216, 231, 246, 254, 277, 279, 328, 339, 342, 359, 379, 382, 406, 419, 425, 450, 459, 464, 466, 480, 503, 525, 531, 550, 556, 598, 661, 672, 688, 698, 729, 733, 764, 779, 791, 793, 811, 846, 853, 891, 912, 922, 932, 945, 956, 958, 963, 965, 985, 989, 1001, 1012, 1016, 1060, 1066, 1079, 1088, 1102, 1127, 1166]. Networks-based [97]. Neural [3, 6, 25, 42, 49, 97, 129, 183, 191, 204, 212, 277, 308, 325, 339, 395, 398, 406, 412, 416, 425, 455, 464, 488, 674, 720, 757, 764, 779, 811, 835, 871, 889, 905, 912, 939, 945, 952, 975, 980, 1016, 1101, 1121, 1132, 1159]. NeuralWiGait [373]. neurodynamics [159]. neuromorphic [398]. neutron [562]. news [122, 177, 249, 299, 381, 791, 1136]. Newton [152]. Newtonian [779]. Next [151, 379, 826, 916, 1069]. next-basket [826]. Next-generation [151, 379, 916]. niche [751]. NLP [177]. NN [66]. no [606]. no-wait [606]. NoC [912]. NoC-based [912]. Node [140, 382, 393, 401, 667, 744, 905, 976]. Node-disjoint [744]. Node-to-set [393]. nodes [57, 503, 893]. noise [518, 520, 535, 736, 739, 1030, 1054, 1063]. noisy [968]. NOMA [576]. Non [63, 79, 630, 669, 722, 739, 779, 801, 863, 1030, 1108]. non-cyclic [722]. Non-IID [63, 739, 863, 1030]. non-invasive [669]. non-negative [79, 801]. non-Newtonian [779]. nonlinear [92, 152]. nonnegative [627]. nonterrestrial [555]. norm [204]. normalization [114, 937]. normalized [638, 802]. NoSQL [650]. NoSQL-based [650]. Note [228, 250]. NoTNER [1106]. Novel [6, 27, 65, 67, 69, 83, 94, 128, 166, 179, 219, 262, 271, 278, 280, 315, 316, 426, 435, 502, 521, 623, 647, 672, 712, 743, 771, 772, 836, 851, 880, 930, 936, 987, 1018, 1082, 1107, 1112, 1156, 1166]. NPGPT [352]. NR [329]. NR-unlicensed [329]. nuclei [174, 199]. number [94, 250]. number-based [250]. numbers [563, 802, 1163]. Numerical [188, 203, 714, 723, 743, 884]. numerous [720].

O [616]. obfuscation [1084]. object [105, 157, 202, 303, 384, 417, 449, 455, 466, 505, 558, 559, 572, 580, 605, 609, 632, 664, 676, 692, 699, 713, 724, 759, 805, 830, 838, 856, 864, 879, 898, 925, 941, 946, 951, 997, 1021, 1034, 1035, 1077, 1093, 1122, 1126, 1141, 1154, 1157, 1162]. objective [54, 59, 80, 88, 98, 116, 124, 255, 287, 293, 348, 377, 440, 577, 608, 643, 694, 705, 722, 723, 751, 755, 897, 942, 969, 1118, 1155]. objectives [725]. objects [385, 638]. oblivious [673]. obstacle [604, 1129]. OCB [567]. occluded [1154]. occlusion [363]. OCSAug [828]. OCT [472, 656]. OCTA [1158]. ODD [202, 384]. ODD-YOLOv8 [202]. ODE [1036]. ODE-guided [1036]. off [372, 396, 1038]. off-chain [372, 396]. offline [653]. offloading [29, 35, 45, 52, 76, 289, 360, 459, 504, 564, 611, 639, 787, 825, 827, 921, 1008, 1029, 1044]. OFSPNet [1036]. oil [716, 829]. on-chain [372, 396]. on-demand [1069]. one [822, 1078]. one-stop [1078]. oneAPI [428]. Online [111, 286, 321, 460, 523, 600, 653, 1115]. online/offline [653]. Onsager [97]. Ontological [693]. Ontologies [243]. ontology [93]. OP2 [1064]. open [417, 465, 780, 913, 1106]. open-world [417]. OpenPose [982]. operating [1031]. operation [1051]. operational [1002]. operator [755, 1135]. operators [43, 294, 578, 1133]. opportunities [291, 915, 932, 958]. Opposition [148, 746, 769, 935]. Opposition-based [148, 769, 935]. Opt [1049]. optical [568, 758, 828, 1040]. optima [1142]. Optimal [102, 212, 218, 377, 538, 744, 776, 878, 1066]. optimisation [545]. Optimization [8, 13, 41, 54, 55, 59, 66, 72, 80, 90, 95, 98, 99, 124, 132, 143, 150, 170, 188, 201, 209, 236, 237, 255, 264, 266, 287, 302, 313, 318, 330, 340, 348, 358, 367, 368, 403, 411, 425, 432, 460, 464, 469, 483, 487, 537, 556, 562, 576, 582, 584, 606, 622, 624, 637, 641, 643, 667, 680, 685, 700, 709, 722, 723, 729, 736, 737, 742, 743, 751, 755, 766, 769, 770, 772, 775, 777, 813, 818, 819, 877, 891, 894, 913, 931, 972, 987, 996, 998, 1008, 1011, 1012, 1045, 1064, 1082, 1086, 1100, 1110, 1122, 1129, 1135, 1138, 1142, 1147, 1151, 1165]. optimization [20, 45, 148, 533]. Optimization-based [736, 1012, 1100]. optimizations [188, 734, 884]. Optimized [6, 47, 191, 316, 333, 359, 391, 423, 429, 443, 636, 669, 697, 715, 854, 890, 1016, 1144]. optimizer [90, 140, 236, 264, 368, 458, 708, 746, 772, 884, 897, 981, 1100]. optimizers [1042]. Optimizing [29, 72, 76, 88, 134, 172, 257, 270, 289, 303, 339, 370, 400, 463, 688, 746, 820, 851, 904, 923, 992, 1095, 1106]. oracle [1145]. oracle-less [1145]. orbit [989]. orchestration [1102]. order [90, 152, 203, 247]. ordered [802]. ordinal [1048]. ordinary [49]. ordinates [517]. organ [1159]. oriented [82, 200]. orthogonal [125, 212, 801]. orthogonal-gating [125]. orthopair [103]. other [296]. outsourced [225]. Outsourcing [296, 497]. overhead [289, 343, 865]. overlapping [901]. oversampling [248, 520, 701].

P2PPO [504]. paced [979]. pack [767]. package [855, 1080]. packet [748, 846, 1086]. Packing [211]. Padovan [197]. pads [723]. PAF [907]. PAF-DETR [907]. PageRank [186]. pages [313]. pair [351, 447, 518, 802]. Paired [273]. pairing [427]. Pakistan [103]. Pallet [471]. palsy [106]. pancake [245]. pandemic [332]. PANs [26]. paradigm [1036]. Parallel [7, 24, 38, 47, 73, 135, 226, 334, 345, 450, 504, 514, 517, 562, 567, 599, 629, 694, 700, 739, 861, 1030, 1058, 1080, 1087, 1101, 1135]. parallelization [451]. parameter [266, 458, 518, 1082]. parameterized [535]. parameters [760, 1135]. parametric [1108]. Pareto [98]. parity [703]. Parkinson [878]. ParsBERT [715]. part [620]. part-of-speech-driven [620]. partial [589, 596, 752, 776, 907]. partially [712, 1051]. Particle [13, 255, 487, 743, 769, 1120, 1165]. particles [623]. particular [1052]. partition [439, 1018]. partitioned [273, 707, 1085]. partitioning [520, 745]. parts [106]. party [9, 372, 1120]. passing [357]. password [38, 1150]. patch [762]. patch-aware [762]. patent [566]. Path [17, 48, 83, 95, 245, 273, 276, 322, 353, 358, 374, 487, 662, 680, 696, 744, 758, 786, 852, 872, 900, 913, 944, 981, 1060, 1072, 1133]. pathfinding [784]. pathology [836]. paths [211, 393, 744]. patient [166]. patient-centric [166]. patients [332]. Pattern [81, 318, 677]. patterns [571, 658, 710, 1146]. payment [1115]. PB [673]. PB-UOKM [673]. PBFT [323, 711]. PCA [1104]. PCA-NCA [1104]. PCB [261, 909, 1103]. PDCF [213]. PDCF-DRL [213]. PDPA [620]. peak [660]. peak-finding [660]. pedestrian [363, 671, 849]. pelican [944]. penetration [995]. perception [430, 919, 1091, 1097]. Perceptual [464]. Performance [41, 57, 156, 187, 260, 265, 329, 344, 400, 411, 441, 468, 484, 498, 539, 588, 602, 625, 648, 677, 763, 882, 893, 902, 949, 953, 1023, 1103, 1144, 1160]. performance-to-cost [588]. Periodic [129, 675, 690, 796]. permutation [434]. Persian [715]. person [7, 56, 220, 914]. personalization [749]. Personalized [158, 821, 826, 882]. perspective [677, 994]. perspectives [949]. perturbed [203]. pervasive [1063]. PESA [477]. PET [656]. PET/oCT [656]. Petri [247]. pharmaceutical [549]. phase [191, 1123]. phishing [69, 313]. photography [492, 830]. photovoltaic [20, 458, 721, 991]. physical [51, 299, 547]. physics [92]. physics-informed [92]. picking [191]. PID [266]. PIFTA4Rec [826]. pinball [217]. pipeline [156, 1048]. pipeline-based [156]. pipelines [451]. piston [723]. pivotal [760]. pivoting [87]. pixel [230]. placement [192, 257, 302, 435, 853, 1125, 1144]. placer [268]. planar [309]. Planning [7, 83, 88, 95, 276, 353, 374, 487, 680, 696, 758, 786, 866, 872, 981, 1129]. plants [394, 833]. plaque [267]. platform [100, 219, 1063]. platforms [121, 538, 705, 789, 1000]. play [75, 493]. pLazyQML [1080]. plug [75, 493]. plug-and-play [75, 493]. PMCKV [156]. Pmsafe [861]. pneumonia [809]. Pocket [235]. point [176, 221, 387, 477, 652, 710, 822, 947, 971, 1026, 1108]. Pointat [971]. poisoning [962]. Polaris [398]. policies [869]. policy [107, 336, 358, 473, 673, 911]. policy-based [673]. pollution [275]. polynomial [179, 519]. polynomials [252]. polyphonic [823]. pooling [176, 240, 267]. popularity [540, 720]. population [582, 623]. portability [625, 893, 1023]. pose [429, 471, 1038]. position [341, 759]. position-scale [759]. positional [282]. positioning [230, 569, 665]. Post [9, 233, 1017, 1075, 1150]. Post-quantum [9, 233, 1017, 1075, 1150]. Potential [1142]. Potential-optima [1142]. potentials [126, 405]. Power [6, 20, 190, 329, 394, 443, 716, 787, 808, 976, 991, 1088, 1095]. power-constrained [190]. powered [553, 577, 603, 933, 1014]. PPG [892]. PPLBB [219]. PPO [504]. PPTopicPLM [75]. Practical [673, 711, 806]. Practical-time [806]. practice [26]. pre [75, 87, 111, 725]. pre-diagnosis [111]. pre-pivoting [87]. pre-trained [75]. pre-training [725]. Preaftrack [941]. Precise [230, 591, 758, 887, 1097]. Precision [87, 109, 214, 249, 277, 761, 1119]. predator [819]. Predefined [1054]. Predefined-time [1054]. predictability [536]. predicting [314, 779]. Prediction [18, 54, 92, 129, 162, 178, 198, 241, 297, 316, 348, 357, 366, 374, 394, 402–404, 420, 422, 475, 489, 627, 642, 655, 671, 674, 690, 698, 716, 720, 722, 733, 749, 752, 796, 807, 808, 829, 857, 918, 964, 975, 984, 1003, 1022, 1024, 1033, 1039, 1042, 1048, 1051, 1055, 1082, 1095, 1121, 1151]. prediction-based [796]. predictions [736]. predictive [325, 542]. predominant [823]. preference [27]. prefilling [77]. Preliminary [586]. Prescribed [626]. preservation [1084]. preserving [14, 111, 127, 128, 219, 246, 397, 421, 773, 797, 800, 817]. pressure [669]. pretraining [187]. prevention [598, 1094]. price [366, 488]. Pricing [278, 457]. principal [918]. principle [124]. principles [291]. printed [617, 974]. prior [327]. PRIORITI [335]. prioritisation [97]. prioritization [335, 728]. prioritized [504]. prioritizing [1015]. priority [52]. priority-aware [52]. Privacy [14, 85, 111, 127, 128, 134, 219, 246, 326, 369, 421, 497, 528, 555, 702, 773, 797, 800, 817, 906, 1020, 1046, 1078, 1084]. privacy-preserving [14, 111, 127, 128, 219, 246, 421, 773, 797, 800, 817]. private [810]. PRNU [25]. PRNU-based [25]. pro [47]. proactive [923]. probabilistic [369]. probability [348]. probability-driven [348]. Problem [102, 161, 234, 393, 435, 478, 501, 556, 606, 819, 1054, 1070, 1076, 1113]. problem-solving [819]. problems [80, 132, 148, 150, 203, 236, 251, 368, 483, 722, 884, 897, 981, 1011, 1110, 1131, 1133, 1142, 1147]. procedure [294, 930]. process [135, 1002, 1015]. processes [325]. processing [190, 271, 398, 402, 568, 919, 1002]. processor [47]. processors [214, 364]. product [352, 415]. product-like [352]. production [88, 549]. productivity [1023]. products [1009]. profiling [635]. profit [84]. program [737, 774]. programmable [573]. programming [356, 629, 1054]. programs [477]. progress [910]. progressive [548, 760]. project [316]. projection [397, 860]. projections [1009]. projects [263]. promote [927]. prompt [420, 620, 629, 644, 847]. Prompt4LJP [420]. proof [97, 689]. proof-of-reputation [97]. propagation [77]. properties [803, 1001]. Proposal [256, 536]. Proposal-level [256]. Proposed [617, 703]. prospects [747]. protection [85, 96, 326, 369, 612, 702, 1046, 1078]. protocol [9, 61, 149, 426, 621, 689, 850, 906, 1120]. protocols [310, 372, 922]. Prototype [422, 927, 1090, 1165]. Provably [653, 850]. provider [278]. providing [797]. Province [351, 447]. provision [525]. proximal [197]. pruning [1052, 1105, 1141]. pseudo [94, 1122]. pseudo-label [1122]. pseudo-random [94]. pseudocode [741]. pseudocode-based [741]. PSI [608]. PSI-MFS [608]. PSO [777, 1135]. public [1006, 1150]. publication [143]. Publisher [287, 603]. PUF [850]. pump [462]. purification [413, 466]. purpose [843]. PV [394, 1095]. PV/wind [1095]. PVGwfa [599]. pyelonephritis [479]. pyramid [988, 1132]. pyramids [775]. Pythagorean [1133].

Q [1008]. Q-Network [1008]. QAMO [1135]. QAMO-PSO [1135]. qARM [565]. QAT [1069]. QAT-LOADng [1069]. QbE [24]. QEKLR [679]. QENN [674]. Qiskit [1088]. QLW [224]. QML [422, 1080]. QNN [347]. QoS [1069]. QP [464]. QSPR [479]. QSVM [78]. quadratic [728, 1054]. QuadTree [775]. qualitative [385]. Quality [144, 525, 668]. Quality-of-service [525]. quantisation [545]. quantized [496]. Quantum [9, 78, 117, 142, 147, 154, 223, 233, 253, 308, 354, 388, 391, 418, 422, 450, 455, 498, 535, 602, 674, 679, 681, 764, 811, 910, 916, 917, 932, 933, 958, 977, 1005, 1013, 1017, 1020, 1028, 1075, 1120, 1135, 1147, 1150, 1159]. quantum-classical [455]. Quantum-driven [933]. quantum-enhanced [674, 679]. Quantum-inspired [1135]. Quantum-resilient [932, 958].

Quantum-resistant [498]. quantum-SAR [223]. quasi [204]. quasi-synchronization [204]. quaternion [204, 425]. quaternion-valued [204, 425]. quaternions [675]. qubit [1028]. query [343, 429]. query-optimized [429]. question [1071]. question-answering [1071]. queue [4, 201]. QuickEn [902].

R&D [1006]. rabbit [313]. rabbits [150]. race [629]. radar [30, 1041, 1090]. radfnet [978]. radial [152]. radiance [757, 889]. radiation [465]. radiomics [1155]. radius [120, 456]. rail [742]. railway [796]. railways [632]. ramification [708]. random [7, 94, 193, 324, 389, 478, 486, 543]. RandomForest [162]. range [140, 343, 563]. range-free-based [140]. rank [43]. rank-aggregation [43]. rate [265, 892, 984]. rating [401]. ratio [241, 588]. ratios [13, 443]. RAVDESS [31]. raw [13]. ray [809, 1096]. RBC [621]. RBC-MSS [621]. RBF [212]. RCCS [1062]. RDSF [1124]. RDSF-UFS [1124]. re [56, 220, 363, 849]. re-identification [56, 220, 363, 849]. REA [1157]. REA-YOLO [1157]. reaction [203]. reaction-diffusion [203]. read [57, 400]. ready [1026]. Real [3, 118, 121, 222, 226, 229, 260, 307, 337, 344, 449, 460, 472, 512, 669, 681, 750, 783, 789, 799, 802, 834, 854, 867, 892, 946, 974, 1068, 1102, 1162]. Real-time [222, 226, 229, 260, 307, 337, 344, 449, 460, 472, 512, 669, 750, 783, 789, 799, 834, 854, 867, 892, 946, 974, 1068, 1102, 1162]. real-world [681]. reasoning [50, 112, 205, 952, 1114]. rechargeable [298]. reciprocal [880]. recognition [10, 12, 31, 37, 275, 342, 346, 373, 407, 408, 433, 453, 512, 534, 546, 560, 596, 646, 652, 735, 799, 809, 823, 828, 854, 873, 890, 954, 956, 957, 968, 1052, 1090, 1106, 1109, 1164]. recognizers [617]. recoloring [1043]. recommendation [79, 122, 208, 221, 234, 304, 347, 412, 437, 540, 581, 613, 736, 821, 826]. recommendations [27, 78, 292]. recommender [153, 215, 227, 272, 383, 843, 942, 1025]. recommending [693]. reconfigurable [362, 536, 573]. reconfiguration [135, 578, 689]. Reconstruction [37, 62, 168, 472, 682, 889, 1106]. Reconstruction-driven [37]. records [317, 396, 1092]. recovery [38, 538, 810]. recurrent [86, 325, 412]. recursive [175, 1137]. redefining [347]. ReDepthNet [1041]. reduce [1086]. reduced [359]. reducing [206, 343, 526]. reduction [167, 514, 1002, 1043, 1054, 1165]. redundancy [183, 238, 1076]. reentrant [948]. Reference [80, 240, 353, 677]. Reference-based [240]. referring [174]. refinement [19, 215, 762, 1122]. Refining [1113]. reflection [338]. region [1041]. regional [473]. regional-contention-driven [473]. regions [103]. registers [586]. registration [19, 70, 480, 926, 947]. registry [15]. regression [380, 679, 802, 1048, 1131]. regularization [55, 320, 627, 807]. reinforcement [35, 113, 130, 161, 184, 185, 188, 213, 286, 293, 295, 328, 478, 485, 523, 564, 596, 639, 741, 786, 820, 825, 827, 844, 852, 869, 995, 1037, 1044, 1065, 1073]. related [126, 405]. relation [688, 814, 870, 896, 901, 914, 970]. relation-aware [688]. Relational [112, 357, 999]. relationship [27, 476, 952]. relationship-aware [27]. relay [77, 976]. released [689]. relevance [238]. reliability [180, 205, 435, 550, 1055, 1076, 1086]. reliability-redundancy [1076]. reliable [237, 256, 641]. ReLU [545, 810]. remote [17, 303, 332, 370, 673, 706]. removal [338]. renewable [1095]. replacement [810, 1000, 1037]. replay [504]. representation [58, 197, 346, 376, 386, 571, 652, 809, 860, 873, 971]. representations [319, 395, 418]. reputation [97, 323]. RepVGG [433]. requirements [557, 793]. resampling [730]. rescaled [217]. Research [28, 221, 228, 241, 306, 351, 363, 447, 540, 553, 603, 646, 738, 833, 932, 958]. reserve [351, 447]. reservoir [300]. reshaping [602]. residual [131, 499, 504, 563, 888, 984, 1072]. residuals [240, 510]. Resilience [51, 358, 538, 852]. Resilient [109, 165, 932, 958, 1144]. resistance [460]. resistant [498, 839, 858]. resisting [252]. ResNet [1040]. ResNeXt [874]. resolution [17, 240, 762, 904, 985, 997]. Resource [135, 192, 278, 292, 333, 360, 391, 459, 463, 504, 557, 558, 615, 636, 700, 787, 825, 831, 923, 931, 964, 1013, 1061, 1062, 1066, 1166]. resource-aware [700]. resource-constrained [558, 636]. resource-limited [831]. resources [1152]. respiration [892]. response [92, 206]. resting [878]. restoration [840, 934]. restricted [563]. results [941]. Rethink [655]. Retraction [228, 250]. retrieval [127, 258, 336, 346, 537, 914, 967, 1078]. reuse [17]. reversed [125]. reversible [210]. Review [253, 418, 700, 747, 910, 949, 986, 1023, 1031, 1147]. reviews [610, 960]. revisited [1146]. reward [336, 374]. Reynolds [655]. RFAConv [155]. RFAConv-CBM-ViT [155]. RFD [134]. RFD-based [134]. RGB [242, 559, 879, 1116]. RGB-D [242, 559]. RGB-T [879]. RHFL [962]. rice [1042]. rice-inspired [1042]. Riesz [714]. RIME [368]. RIMS [1127]. RIMS-RPL [1127]. ring [9, 14, 117]. RIOKV [343]. RISC [398]. RISC-V [398]. risk [693, 1002]. risks [275, 446]. RL [485, 852, 1127]. RL-based [1127]. RL-EAR [485]. RNN [184]. road [380, 430, 746, 1024, 1035, 1154]. Roaen [125]. RoboSoccer [604]. Robot [95, 487, 981]. robotic [626, 952, 1129]. robotics [1014]. robots [682]. Robust [25, 84, 185, 244, 271, 406, 866, 871, 959, 962, 1072, 1115, 1124, 1161]. robustness [518, 961, 1131]. Rock [422]. RocksDB [57]. root [195]. rotation [677]. rough [633, 769]. roulette [99]. round [1059]. route [637]. route-based [637]. routing [8, 274, 345, 379, 426, 485, 529, 750, 793, 909, 922, 965, 1012, 1069]. routing-next [1069]. RPL [793, 1127]. RPL-based [793]. RSA [102]. RSSI [990]. RT [290, 307, 898]. RT-DETR [898]. RT-DETRmg [307]. RTAB [574]. RTDETR [830, 885]. RTL [1146]. rubbing [1094]. Rubik [883]. Rule [205, 243, 402, 834]. rule-based [402, 834]. rules [655]. rumor [75, 476]. rung [103]. runway [1162]. rupture [1148]. RV [398]. RV-SNN [398]. RVV [172].

S [806]. S-boxes [806]. SA [705]. Sac [1066]. Sac-lstm [1066]. sacroiliitis [634]. Safety [205, 288, 866, 1031]. Saffe [940]. salesman [161]. salient [106, 699]. salp [1059]. SALSTM [115]. salted [522]. SAM [868]. same [976, 1036, 1163]. sample [138, 289, 726]. samples [756]. sampling [68, 145, 387, 476, 961]. sampling-based [68]. sand [113, 168, 318]. sand-dust [168]. sandbox [151]. Sanitation [583]. SAPFIS [73]. SAR [34, 223, 1141]. Sarcasm [548, 1136]. SatEC [504]. satellite [233, 989]. saving [647, 976]. SBERT [851]. SBEW [951]. SBEW-YOLOV8 [951]. SCA [366]. SCADA [1088]. scalability [797]. Scalable [259, 451, 512, 542, 872, 1161]. scale [10, 17, 32, 40, 41, 49, 70, 91, 104, 163, 199, 277, 387, 412, 416, 429, 431, 449, 484, 507, 514, 561, 565, 631, 632, 638, 759, 783, 788, 808, 843, 855, 856, 861, 885, 892, 926, 928, 936, 937, 946, 951, 952, 956, 967, 1014, 1026, 1032, 1033, 1074, 1097, 1126, 1137, 1139, 1141]. scaled [55]. scales [425]. scaling [626, 1000, 1015]. scenario [554]. scenarios [692, 838, 1117]. Scene [107, 138, 505, 612, 682, 691, 1035]. scenes [67, 889]. SCG [55]. scheduler [427]. schedules [298]. Scheduling [4, 116, 226, 293, 295, 361, 377, 434, 478, 588, 604, 615, 619, 643, 694, 705, 745, 755, 770, 813, 820, 924, 948, 969, 994, 1000, 1029, 1062, 1085, 1152]. scheme [14, 85, 96, 119, 203, 213, 233, 289, 301, 369, 371, 372, 430, 454, 519, 539, 653, 673, 695, 768, 816, 858]. schemes [552, 651]. Schwarz [945]. scientometric [610]. SCNet [554]. SCNet-YOLO [554]. scoring [104, 335, 711, 919]. scoring-based [711]. scrambling [883]. script [69]. script-based [69]. SD [291, 1125]. SD-IoT [1125]. SD-WAN [291]. SDHNet [68]. SDN [379, 436, 1161]. SDSMS [371]. SDSMS-LoRa [371]. Sea [265]. sealed [385, 816]. search [124, 212, 216, 251, 302, 321, 501, 565, 582, 755, 810, 909, 972, 1070]. search-based [212, 302]. searchable [119]. seasons [894]. secret [621]. Secure [9, 12, 94, 271, 296, 371, 522, 552, 653, 673, 774, 793, 800, 845, 850, 858, 1075, 1120]. Securing [181, 246]. security [69, 72, 108, 121, 165, 310, 359, 362, 461, 521, 555, 585, 636, 651, 731, 903, 906, 932, 933, 958, 1020, 1023, 1081, 1149]. segment [868]. segmentation [60, 74, 90, 113, 157, 174, 176, 199, 235, 242, 267, 331, 410, 591, 640, 656, 684, 687, 718, 732, 762, 776, 788, 861, 868, 904, 978, 1026, 1041, 1117, 1132, 1156, 1158]. segmented [115, 956]. Segrè [1009]. select [283]. selectable [39]. selected [490]. selecting [837]. selection [2, 89, 134, 147, 150, 212, 238, 254, 255, 316, 351, 370, 419, 440, 447, 532, 564, 608, 622, 633, 636, 657, 667, 712, 769, 820, 878, 955, 984, 987, 1051, 1059, 1082, 1099, 1100, 1124, 1155]. selections [592]. Selective [133]. selector [489]. Self [28, 56, 58, 115, 121, 137, 267, 336, 338, 390, 474, 481, 490, 560, 631, 633, 873, 979, 997, 1004, 1106, 1159, 1164]. self-adapting [137]. self-adaptive [336, 481, 1159]. self-attention [28, 115, 267, 474, 560, 997]. Self-aware [121]. self-distillation [56, 1004]. self-driving [1164]. self-expressiveness [390]. self-forgery [631]. self-information-based [633]. self-optimizing [1106]. self-paced [979]. self-supervised [58, 338, 490, 873].

Semantic [64, 74, 242, 299, 346, 375, 508, 529, 572, 654, 682, 718, 940, 970, 978, 1026, 1041]. Semantic-Alignment [940]. Semantic-aware [682]. semantic-physical [299]. semantics [967]. Semi [169, 210, 269, 376, 749, 799, 863, 1085, 1156]. semi-asynchronous [749]. semi-clairvoyant [1085]. semi-reversible [210]. Semi-supervised [169, 269, 376, 799, 863, 1156]. SEMLP [795]. sensing [17, 303, 370, 702, 706]. sensitive [564]. sensitivity [477]. sensor [2, 8, 133, 170, 209, 298, 342, 485, 853, 862, 891, 922, 956, 1094, 1102]. sensor-based [1094]. sensors [401, 534]. sensory [792]. sentiment [46, 58, 64, 86, 125, 131, 189, 200, 341, 531, 620, 644, 668, 683, 715, 795, 804, 876, 943, 980]. separable [761]. SeqMatcher [355]. sequence [7, 197, 355, 599, 677, 821, 873]. sequence-to-graph [599]. sequencing [88]. sequential [81, 412]. SER [888]. SER-GAN [888]. SERDNet [1072]. serialization [1061]. series [68, 114, 129, 179, 282, 416, 493, 499, 716, 778, 792, 807, 875, 886, 928, 986, 993, 1010, 1050, 1055, 1076, 1087, 1130, 1134]. server [84, 257, 441, 460, 1144]. serverless [206, 1061, 1062]. servers [296]. Service [84, 89, 201, 278, 436, 444, 525, 561, 611, 992, 1055]. services [685]. session [292, 304, 371, 716]. session-based [292, 304]. session-enhanced [716]. set [225, 304, 351, 353, 393, 397, 447, 501, 633, 769, 1070]. sets [173, 565, 740]. settings [297, 404]. SF-CLIP3D [805]. SFC [523]. SFOD [256]. SFVE [467]. SGB [511]. SGB-YOLOv5 [511]. SGSLNet [176]. shallow [625]. Shannon [65]. SHAP [890]. SHAP-based [890]. shaped [3, 752]. shard [689]. sharding [411]. shared [707]. sharing [361, 378, 396, 430, 586, 621, 673, 800, 833]. shark [746]. sharpness [306]. sharpness-aware [306]. SHAVER [586]. shift [1113]. shifts [490]. ship [34, 554, 885, 1119, 1141]. shop [293, 434, 478, 606]. short [55, 75, 115, 117, 343, 808, 878, 963, 991, 1042]. short-range [343]. short-term [55, 115, 117, 963, 991, 1042]. short-text [75]. short-time [878]. shorter [1060]. shortest [322, 1133]. shot [139, 583, 654, 870, 908, 1007, 1090, 1093, 1107, 1122]. siamese [481, 570, 879]. Siamsdt [481]. side [69, 252, 598]. sidelobe [318]. sign [530, 618, 1164]. signal [190, 735, 1090]. signals [660, 892, 956]. signature [14, 44, 468, 519, 522, 607, 653]. signature* [468]. signatures [1017]. signer [530]. signer-independent [530]. signers [530]. significant [23, 210]. signing [372]. signs [307]. SIKE [463]. Sim [527]. Sim-ConvFormer [527]. SimAM [527]. SIMD [398]. SIMD-style [398]. Similar [581]. similarity [124, 150, 458, 592, 768, 1092]. simple [1063]. simplified [634]. Simulation [132, 201, 247, 300, 450, 573, 723, 893]. simulator [484]. simulators [516]. Simultaneous [95, 1086]. Single [169, 338, 577, 649, 652, 703, 937, 952, 1028, 1120]. single-cell [649]. Single-image [169, 338, 937]. single-objective [577]. single-particle [1120]. single-point [652]. single-qubit [1028]. singular [812, 832]. Singularity [445]. singularly [203]. sink [853, 862]. site [351, 447]. situ [225]. situation [73, 461]. size [648, 1086]. size-dependent [648]. skeleton [10, 37, 873, 1116]. skeleton-based [10, 37]. sketch [21, 444]. sketch-based [21]. skew [1143]. skewed [919]. skin [327, 551]. skyline [216]. SLAM [574, 682]. Slerp [1049]. Slerp-Opt [1049]. slice [499]. sliding [321]. slopes [422]. Small [202, 303, 307, 308, 384, 414, 415, 505, 605, 830, 856, 886, 898, 936, 951, 1034–1036, 1154, 1157]. Smart [44, 97, 297, 372, 403, 404, 446, 506, 508, 587, 646, 709, 789, 796, 818, 850, 1081, 1128]. smartphone [272, 383]. smoke [612]. Smooth [217]. smoothed [860]. SMSERT [783]. SMSERT-DETR [783]. snake [624, 981]. SNN [398]. SoC [481]. Social [27, 145, 146, 208, 215, 227, 272, 284, 383, 437, 556, 613, 1106]. social-curiosity-based [227]. social-media [284]. soft [294, 528]. soft-biometric [528]. SoftMax [33]. software [291, 426, 485, 489, 597, 846, 917, 963, 1015]. software-defined [291, 426, 485, 963]. SOG [1035]. SOG-YOLO [1035]. solar [20, 458]. solution [151, 435, 561]. solutions [604, 796, 797, 930, 932, 958, 1006]. solve [161, 714]. solver [589]. Solving [102, 132, 236, 722, 767, 819, 897, 1011, 1054, 1110]. sort [164, 1077]. sound [288, 801]. source [25, 316, 598]. source-side [598]. sources [635]. space [622, 654, 714, 761, 810, 950, 998, 1124]. spam [1114]. span [857]. spanning [334]. Spark [506, 514]. sparse [175, 466, 690, 1070]. sparsity [1063, 1105, 1131]. Spatial [169, 281, 387, 507, 510, 571, 591, 658, 710, 783, 805, 808, 873, 997, 1093, 1139]. spatial-channel [169, 591]. Spatial-frequency [1093]. spatial-frequency-enhanced [805]. spatial-spectral [507]. spatial-temporal [281, 808, 1139]. spatially [490, 888]. Spatio [18, 129, 389, 698, 798, 986]. Spatio-temporal [18, 129, 389, 698, 798, 986]. Spatiotemporal [10, 325, 690, 699, 997, 1024]. special [622]. specific [93, 218]. specifications [314]. Spectral [120, 456, 507, 778, 801]. Spectre [1089]. spectrogram [385]. spectrum [14, 702]. Speech [28, 560, 620, 1159]. Speed [18, 241, 482, 563, 1038]. Speed-accuracy [1038]. speeding [726]. spherical [173, 582, 740, 837]. spiking [398]. spiral [39]. spirals [619]. spiro [324]. spline [775]. split [63, 196]. splitting [186, 203, 467].

Splp [1162]. Splp-yolo [1162]. Spotify [581]. spreaders [1001, 1079]. spreading [1001]. spyware [834]. Sql [532]. SQUAD [917]. squaring [391]. squeeze [86, 1072]. squeeze-and-excitation [1072]. SqueezeNet [892]. SS [303]. SSA [212]. SSCENet [507]. SSD [516]. SSIM [490]. SSMIM [799]. SSO [123]. SSO-based [123]. ST [798, 1003]. ST-FPN [1003]. St-Graphormer [798]. stability [425]. stacked [178]. Stackelberg [228]. Stacking [794, 955]. stacking-based [955]. stage [19, 62, 101, 431, 476, 512, 639, 676, 751, 776, 782, 983]. staleness [749]. Star [550, 672, 970]. Star-transformer [970]. start [234, 1061]. State [185, 469, 761, 878]. State-space [761]. states [1120]. Static [198, 534, 862]. station [7, 837]. Statistical [366]. stay [1048]. STD [24]. STDP [398]. steel [833]. steepest [148]. steganography [1149]. Steiner [211]. STEM [882]. step [186]. Stepwise [802]. Stff [830]. Stff-rtdetr [830]. STGEN [698]. STMACN [1139]. STO [1161]. stochastic [322, 514]. Stock [366, 1033, 1082]. stop [1078]. Storage [57, 321, 537, 650, 673, 833, 902]. storage-efficient [321]. store [259, 400]. stores [156, 343]. storytellers [263]. STPNet [325]. strategic [1006]. strategies [45, 204, 395, 457, 705, 972, 998, 1005, 1043, 1049, 1056, 1099, 1152]. strategy [4, 8, 54, 58, 80, 87, 125, 134, 138, 168, 170, 251, 266, 326, 458, 483, 500, 533, 564, 578, 601, 624, 680, 743, 745, 758, 811, 884, 948, 981, 996, 1018, 1037, 1044, 1073, 1144]. stratified [176]. straw [511]. streaks [1163]. stream [19, 68, 139, 311, 559, 578, 600, 730, 879, 919]. streaming [428, 588, 666]. streamlines [300]. streams [766, 1053]. strengthened [440]. strengths [629]. strip [88, 706]. strongly [180]. structural [92, 176, 382, 422, 785]. Structure [132, 199, 524, 803, 828, 1010, 1134]. structured [623, 813]. structures [659, 806]. student [754]. study [152, 265, 292, 308, 326, 351, 401, 445, 447, 542, 605, 922, 1146]. style [305, 398]. sub [230, 279]. sub-networks [279]. sub-pixel [230]. subgraph [248]. subject [1098]. subsection [725]. subsection-aware [725]. subspace [320, 390, 495]. substructure [524]. subtype [1155]. success [1015]. successful [263]. successive [860]. suitable [727]. sum [309]. summarization [171, 725]. summation [1120]. Sunway [562]. super [17, 240, 985, 997]. super-resolution [17, 240, 985, 997]. supercomputer [562]. Supercomputing [602, 915, 1006, 1026]. supercomputing-ready [1026]. supervised [58, 74, 169, 269, 311, 338, 376, 388, 490, 799, 863, 873, 1156]. supervising [1074]. supervision [1004]. supervoxel [640]. support [78, 217, 229, 1127]. supported [192]. surface [40, 155, 788, 974, 1097]. surfaces [362, 415]. surrogate [300, 1011, 1110]. surrogate-assisted [1011]. surrogate-based [1110]. surveillance [297, 404, 833, 838]. survey [141, 291, 310, 516, 543, 614, 651, 763, 865, 922, 932, 958, 1128]. survivability [853]. sustainable [123, 953]. SVDFormer [832]. SVM [271, 878]. SVN [1119]. SVN-YOLO [1119]. SW [842]. SW26010 [47]. SW26010-pro [47]. swap [393, 524, 859]. Swarm [13, 95, 255, 318, 487, 743, 769, 770, 998, 1059, 1165]. SwarmCNN [777]. swarms [607]. sweep [83]. swin [684, 1003]. switch [460, 1125]. Swpmmas [47]. SYCL [452, 625]. symmetric [500, 554, 660]. synchronization [6, 204]. synchronous [514]. synergies [1020]. Syntactic [189, 367, 668]. Syntactic-guided [367]. syntax [200]. synthesis [318]. Synthetic [248]. system [14, 38, 42, 47, 51, 73, 94, 108, 110, 141, 152, 153, 156, 185, 209, 229, 241, 259, 272, 278, 289, 297, 330, 332, 337, 344, 362, 373, 383, 404, 421, 433, 453, 454, 469, 511, 515, 517, 547, 549, 612, 616, 619, 636, 637, 642, 677, 708, 736, 769, 797, 839, 843, 852, 859, 911, 919, 1027, 1031, 1068, 1076, 1081, 1085, 1090]. system-of-systems [852]. Systematic [253, 747, 780, 1023, 1146, 1147]. systems [39, 45, 57, 112, 149, 192, 197, 215, 227, 254, 283, 321, 349, 419, 435, 441, 473, 513, 578, 650, 694, 742, 748, 763, 771, 796, 852, 866, 912, 942, 1002, 1025, 1152]. SZ4IoT [392].

T [378, 583, 879]. T-BFL [378]. T-Sanitation [583]. table [210]. table-based [210]. tabu [501]. tackling [443]. tactical [895]. tags [585]. tailed [407]. tailoring [1135]. tardiness [606]. Target [22, 95, 413, 414, 492, 579, 662, 790, 868, 936, 1036, 1097]. targets [1154]. Task [11, 29, 35, 52, 76, 141, 192, 279, 289, 295, 377, 504, 557, 568, 604, 615, 619, 639, 643, 644, 741, 745, 755, 787, 827, 870, 901, 952, 969, 1022, 1029, 1044, 1151, 1152]. tasking [935]. tasks [4, 99, 121, 226, 564, 770, 1000]. taxonomy [773]. TC [163]. TC-BERT [163]. TCN [1095]. TCP [474]. TCP-based [474]. teacher [754]. tears [526]. technique [248, 360, 415, 462, 672, 825, 1067]. techniques [72, 247, 284, 340, 342, 555, 593, 598, 780, 819, 845, 916, 961]. technologies [452, 568, 681]. technology [163, 167, 566, 646, 843]. temperature [162, 394, 716, 829, 965]. temperature-aware [965]. template [481]. Temporal [18, 71, 129, 281, 311, 389, 488, 627, 698, 720, 733, 798, 808, 826, 848, 864, 873, 928, 938, 986, 1033, 1139]. Temporal-spatial [873]. tenant [774]. Tensor [320, 354, 450]. term [55, 68, 115, 117, 475, 499, 808, 928, 963, 991, 1042, 1083]. terminal [611]. terms [714]. ternary [190, 568]. terrestrial [555]. terrestrial-nonterrestrial [555]. test [597, 728, 944, 987, 1113, 1151]. test-generation [597]. test-time [1113]. testability [1145]. testing [133, 917, 995].

TestLock [1145]. tethered [538]. Text [75, 101, 138, 171, 367, 488, 500, 851, 855, 914, 1106]. text-based [488]. Text-to-image [914]. texts [715]. textual [46, 386, 683]. texture [344]. TGAC [835]. their [53, 106, 173, 294, 571, 740, 1009, 1133]. Theoretical [677]. Theory [228, 401, 403, 487, 544, 590, 829, 1056]. therapeutics [5, 693]. Thermal [534, 553, 603, 776, 779, 978]. thermoplastic [13]. thickness [779]. Things [123, 271, 301, 332, 547, 610, 653, 773, 845, 905, 1087]. Things-cloud [845]. third [1112]. thoughts [50]. threat [142, 335, 628]. threats [246, 547]. Three [9, 139, 680, 806, 845, 883, 1009]. three-dimensional [680, 883, 1009]. Three-stream [139]. threshold [90, 113]. thresholds [426]. throughput [398]. tier [1058]. Time [68, 114, 116, 121, 128, 129, 159, 179, 203, 206, 222, 226, 229, 260, 262, 282, 307, 337, 344, 377, 385, 389, 412, 416, 425, 449, 460, 472, 499, 512, 536, 595, 626, 669, 704, 714, 716, 750, 778, 783, 789, 792, 799, 806, 807, 834, 854, 867, 875, 878, 886, 892, 928, 946, 974– 976, 986, 991, 993, 1010, 1050, 1054, 1055, 1068, 1087, 1102, 1113, 1130, 1134, 1162]. time-aware [377]. time-dependent [203]. Time-frequency [262, 385, 595, 975, 991]. time-series [282, 499, 792, 1130]. time-slice [499]. time-space [714]. TimeAttentionBiLSTM [1083]. timed [247]. times [1061]. timing [268]. timing-driven [268]. tiny [715, 721]. Tiny-ParsBERT [715]. TLNN [96]. TMATrack [724]. TMTC [670]. token [74, 408, 724, 857]. tolerance [711, 891, 1144]. tolerant [564, 626, 672, 753, 1029, 1060]. tomography [1040]. tool [451, 477, 988]. top [876]. top-down [876]. Topic [75, 171, 375, 566, 999]. topic-enhanced [75]. topological [250, 324, 479, 803]. topology [10]. TOPSIS [85, 103, 173, 351, 447, 740]. torus [48]. torus-like [48]. total [309, 606]. touch [685]. Tp [546]. Tp-yolov8 [546]. TPU [60]. traceable [14]. track [494]. tracker [432]. tracking [71, 337, 481, 682, 724, 759, 838, 879, 925, 941, 997, 1077]. tractive [1107]. trade [1038]. trade-off [1038]. trading [14]. Traffic [16, 18, 129, 198, 281, 307, 413, 546, 583, 598, 612, 618, 635, 690, 692, 698, 749, 798, 818, 834, 835, 867, 908, 1024, 1067, 1083, 1117, 1125, 1139, 1164]. traffic-based [834]. Train [494, 617]. trained [75]. training [101, 212, 316, 542, 589, 725, 756, 943]. trajectory [143, 655, 671, 1046, 1121, 1129]. transaction [97, 689]. transaction-released [689]. transactions [77]. Transfer [80, 305, 348, 471, 683, 782, 824, 870, 1084, 1138]. transferability [609]. transform [39, 416, 713, 782, 926, 1028]. transform-enhanced [713]. transformation [648, 915]. Transformer [3, 19, 86, 144, 155, 161, 191, 221, 262, 280, 281, 304, 305, 317, 350, 365, 408, 429, 474, 500, 670, 671, 684, 690, 716, 798, 824, 832, 848, 890, 892, 947, 957, 966, 970, 988, 993, 1003, 1010, 1019, 1032, 1091, 1103, 1111]. transformer-based [350, 1003]. Transformer-cnn [966]. transformer-facilitated [304]. transformers [282, 349, 526, 823, 1136]. transistor [1101]. transit [742]. translation [530]. transmission [241, 858, 976]. transmissions [1058]. transparency [890]. transport [446, 562, 1081]. transportation [45, 192, 637]. trapezoidal [837]. traveling [161]. Treating [243]. treatment [386]. tree [50, 57, 225, 343, 515]. trees [334, 846]. trend [1033]. trends [1031, 1147]. tri [644]. tri-encoder [644]. triage [294]. Triangle [145]. Triangle-induced [145]. triangular [250]. Triaxial [876]. trick [604]. trigonometric [680]. trip [292]. triple [476, 999]. triple-relationship [476]. Triplet [82, 620]. Trojan [794]. TRSBi [1103]. TRSBi-YOLO [1103]. truck [612]. truncated [269, 971, 1040]. Trust [232, 378, 401, 641, 793, 839, 905, 1017, 1025]. trust-aware [905, 1025]. trust-driven [641]. trust-rating [401]. trusted [670, 1069]. trustworthy [542]. TS [705]. TS-guided [705]. TSCANet [139, 311]. TSESRec [304]. TSFN [1130]. tubular [687]. tubular-aware [687]. tumor [235, 345, 502, 656, 684, 1132]. tuning [50, 266, 587, 737, 844, 1152]. tuple [103]. turbine [808, 829, 1097]. turbines [32]. Turbulence [858]. Turbulence-resistant [858]. TVM [172, 214]. tweets [736]. twins [789]. Two [101, 136, 296, 311, 378, 382, 512, 600, 639, 660, 676, 751, 851, 927, 1123]. two-branch [927]. two-dimensional [136, 378, 600, 660]. two-layer [382]. Two-phase [1123]. two-stage [101, 512, 639, 676, 751]. two-stream [311]. type [53, 637, 760]. type-II [760]. types [392, 424, 908].

U [3, 70, 752]. U-Net [70]. U-shaped [3, 752]. UAV [192, 202, 384, 492, 557, 662, 680, 713, 787, 827, 854, 868, 879, 898, 907, 1037, 1126, 1157]. UAV-enabled [827]. UAVs [11, 83, 790, 936]. UCP [741]. UCS [936]. UCS-YOLO [936]. UDS [1117]. UDS-YOLO [1117]. UFS [1124]. Ultra [495, 808, 1070, 1141]. ultra-high-dimensional [495]. ultra-large [1070]. Ultra-lightweight [1141]. ultra-short-term [808]. ultrasound [267]. UML [738]. unbalanced [102, 319]. unbiased [691]. Uncertainty [315, 967, 1122]. Uncertainty-guided [1122]. Underlying [93]. undersampling [979]. understanding [112, 847, 1108]. Underwater [157, 244, 579, 638, 759, 773, 840, 888, 934, 966, 1091, 1111]. UNet [331, 684, 706]. unified [281, 627, 741, 1113]. unimodal [660]. union [970]. unique [160]. unit [1164]. unitary [1028]. universal [815]. unlearning [697, 717]. unlicensed [329]. Unlocking [443]. unmanned [105, 276, 302, 852]. Unpaired [48]. unstructured [909]. Unsupervised [37, 575, 654, 657, 1107, 1124]. unveiled [10]. Unveiling [20, 635]. UOKM [673]. updatable [673]. update [150, 225]. upon [379]. upper [1149]. Urban [221, 223, 297, 404, 637, 1024, 1026]. URL [438]. use [681]. User [27, 72, 89, 93, 122, 328, 329, 369, 540, 736, 745, 881, 1022, 1066]. user-centric [736]. user-friendly [881]. user-specific [93]. user-to-multiple [72]. users [194]. Using [5, 23, 24, 33, 43, 77, 78, 88, 89, 100, 106, 113, 116, 117, 140, 177, 179, 181, 182, 209, 210, 247, 284, 285, 292, 314, 321, 328, 331, 342, 346, 347, 354, 364, 390, 399, 401, 402, 410, 422, 423, 450, 455, 474, 479, 497, 510, 567, 570, 573, 581, 597, 598, 623, 626, 634, 635, 637, 638, 644, 659, 674, 694, 704, 705, 709, 731, 776, 777, 792, 801, 825, 837, 846, 850, 877, 882, 887, 890, 892, 918, 944, 949, 956, 972, 976, 1015, 1051, 1053, 1075, 1091, 1099, 1101, 1112, 1114, 1131, 1136, 1159, 1166]. using [36, 49, 230, 257, 373, 438, 444, 457, 702, 817, 878, 881, 1010]. UST [752]. UST-SU [752]. USV [696]. utility [81, 571]. utilization [289, 292, 931]. utilizing [521, 816]. uTransformer [281]. UTS [868]. UTS-SAM [868]. UWSoS [358].

V [398]. v1.1 [371]. VAE [16, 495]. VAES [567]. validation [1150]. value [57, 259, 343, 400, 812, 832]. valued [204, 216, 425, 778, 985]. values [324]. VANETs [426]. variability [615]. variable [123, 152, 170, 778, 1118]. variable-order [152]. variance [514]. variances [324]. variants [186]. Variational [5, 16, 97, 117, 154, 346, 811, 1096]. various [310]. VASP [484]. VBATS [601]. vector [62, 78, 98, 217, 586, 801, 1069]. vectorization [428]. vectors [80]. Vehicle [29, 105, 241, 275, 290, 302, 314, 379, 430, 475, 576, 655, 748, 837, 1073, 1121]. vehicle-road [430]. vehicle-to-infrastructure [576]. vehicles [276, 600, 749]. vehicular [76, 564]. velocity [170, 426]. velocity-aware [426]. VEM [708]. vendor [206]. ventilation [51]. veracity [285]. verifiable [296, 421, 486, 817]. verification [522, 607, 663, 738]. Verilog [1146]. verse [897]. vertical [697]. vertically [707]. vessel [1158]. vest [727]. VFCkM [707]. VGG [709]. VGG-16 [709]. via [16, 56, 63, 131, 167, 180, 293, 338, 375, 386, 407, 452, 488, 575, 609, 629, 633, 690, 722, 775, 784, 824, 845, 856, 869, 871, 872, 888, 946, 971, 982, 984, 1026, 1049, 1051, 1058, 1061, 1087, 1125, 1143]. vibrant [297, 404]. vibrating [197]. video [144, 297, 404, 408, 522, 670, 699, 752, 833, 985]. view [491, 720]. viewpoint [677]. views [91]. virtual [142, 184, 237, 260, 469, 573, 623]. virtualization [1128]. visibility [579]. visible [22, 56, 220, 333, 765, 849, 978, 1056]. visible-infrared [56, 220, 849]. Vision [155, 172, 262, 280, 305, 349, 350, 526, 549, 569, 642, 773, 805, 847, 925, 988]. vision-based [549]. Vision-language [805, 847, 925]. vista [297, 404]. vista-lite [297, 404]. visual [46, 107, 407, 409, 467, 481, 574, 678, 682, 683, 724, 782, 967]. visual-textual [46, 683]. visualizer [881]. VIT [155, 363, 1104]. vital [1079]. VITS [694]. VLAD [792]. VLIW [172, 667]. VLSI [812, 1003]. VNDN [637]. VolRec [472]. Volterra [731]. volumetric [472]. voting [1059]. voxel [466]. VPSNet [466]. vulnerabilities [731, 780, 871]. vulnerability [508]. VVC [464, 1053].

wafer [761]. wait [606]. walk [543]. walk-based [543]. walks [389]. WAN [291]. warning [394]. WARNs [1050]. Wasserstein [638]. water [195, 251, 625, 1068]. watershed [103]. Wave [270, 642]. WAVECAP [424]. wavelet [424, 713, 782, 872, 926, 1050, 1158]. wavelet-based [872]. wavelet-capsule [424]. WBSN [647]. weak [1077]. weakly [74, 311, 495]. weakly-supervised [311]. weapon [852]. wear [723, 988]. Wearable [1094]. weather [505, 1162]. web [313, 1114]. weight [61, 85, 385, 417, 759]. weighted [238, 606, 665, 730, 899, 1001, 1092, 1125, 1133]. weighting [341, 701, 896]. well [300]. well-based [300]. Wetland [512]. whale [8, 209, 709, 813, 1138]. WHANet [1158]. while [399]. White [236, 410, 746, 1104]. White-faced [236]. Whitney [987]. who [263]. wide [6, 291, 728]. Wiedemann [521]. WiFi [373]. WiFi-based [373]. wild [188]. wind [32, 808, 829, 1095, 1097]. window [213, 321]. wing [108, 1027]. wire [222]. wirelength [218]. wireless [2, 8, 170, 209, 298, 485, 796, 853, 862, 891, 922, 965, 1102]. wise [16, 145, 983, 984, 1105]. within [385]. WK [175]. WK-recursive [175]. WOA [709]. WOA-FMO [709]. wolf [767, 819, 877, 1042]. words [375]. workface [968]. Workflow [364, 994]. workflows [206, 465]. working [1098]. workload [964]. workloads [813]. world [417, 681]. worn [342]. worst [837]. WRF [445]. written [381]. WSN [140, 401]. WTDD [32]. WTDL [926]. WTDL-Net [926]. WTMTOA [432]. WV [201].

X [809, 1096]. X-ray [809, 1096]. XGBoost [16, 630, 1039]. XGBoost-based [16]. XIDINTFL [16]. XIDINTFL-VAE [16].

Yager [1133]. Yannakakis [497]. yarn [239]. yarn-dyed [239]. YOLO [34, 105, 303, 415, 493, 494, 505, 551, 554, 558, 579, 580, 618, 662, 664, 692, 867, 936, 1034, 1035, 1052, 1103, 1117, 1119, 1154, 1157, 1162]. YOLO-based [558]. YOLO-CE [579]. YOLO-EDGE [692]. YOLO-Fast [580, 664]. YOLO-SS [303]. YOLO11 [579]. YOLOv10n [1119]. YOLOv11 [931, 1068]. YOLOv11-based [1068]. YOLOv11n [974]. YOLOv5 [261, 511, 1112].

YOLOv5-based [1112]. YOLOv7 [554, 721]. YOLOv7-tiny-based [721]. YOLOV8 [32, 34, 202, 275, 365, 384, 505, 546, 574, 727, 946, 951]. YOLOv8-based [365]. YOLOv8-WTDD [32]. YOLOv8s [222, 854]. YOLOv8s-GSW [222]. YOLOv9 [492]. YOLOX [605].

zebra [266].  zero [685, 689].  zero-

knowledge [689]. zero-touch [685]. ZNN [1054].

References

Trigui:2025:AIL

[1] Omar Trigui, Sawsan Daoud, Mohamed Ghorbel, Mariem Dammak, Chokri Mhiri, and Ahmed Ben Hamida. Automated identification and localization of interictal epileptiform discharges: leveraging morphological analysis, five-criterion fulfillment, and machine learning approach. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06558-z.

Alimohammadi:2025:EAC

[2] Edris Alimohammadi, Sajad Haghzad Klidbary, and Mohammad Javadian. Energy-aware clustering method for cluster head selection to increasing lifetime in wireless sensor network. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06474-2.

Bai:2025:GGA

[3] Xuefei Bai, Yongsong Wan, Weiming Wang, and Bin Zhou. GUFORMER: a gradient-aware U-shaped transformer neural network for real image denoising. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06576-x.

Duan:2025:MQB

[4] Lintao Duan and Haiying Wang. Multiqueue-based energy-efficient scheduling strategy for tasks with deadline constraints in cloud data center. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06580-1.

Wang:2025:CBC

[5] Chengling Wang, Yuexin Zhang, Yunru Ma, Peng Chen, and Yang Xiang. CNN-based continuous authentication for digital therapeutics using variational autoencoder. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06490-2.

Dehghani:2025:NEK

[6] Abbas Dehghani, Sadegh Fadaei, and Resul Das. A novel effective key synchronization approach based on optimized deep neural networks for IoT-based low-power wide area networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06571-2.

Bailin:2025:ADR

[7] Li Bailin, Chen Ao, Wu Panqi, Zhang Chao, and Fu Wenlong. Application of discrete random forest algorithm in multi-person asynchronous parallel disassembly sequence planning for hydropower station equipment maintenance. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06540-9.

Sheng:2025:EER

[8] Hao Sheng, Chen Jun, Cui Jianqun, Fan Xiying, and Li Zhen. An energy-efficient routing algorithm for dual-energy harvesting-assisted wireless sensor networks based on whale optimization strategy. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06536-5.

Kumar:2025:DAP

[9] Uddeshaya Kumar, Manish Garg, and Dharminder Chaudhary. Design and analysis of a post-quantum secure three party authenticated key agreement protocol based on ring learning with error for mobile device. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06467-1.

Chen:2025:MSS

[10] Hongwei Chen, Jianpeng Wang, and Zexi Chen. Multi-scale spatiotemporal topology unveiled: enhancing skeleton-based action recognition. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06531-w.

Yan:2025:DTA

[11] Shaokun Yan and Yuanqing Xia. A distributed task allocation method for heterogeneous UAVs in dynamic and communication-constrained environments. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06517-8.

Wang:2025:CBS

[12] Ruyan Wang, Qinglin Zeng, Zhigang Yang, and Puning Zhang. Cloudbased secure human action recognition with fully homomorphic encryption. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06512-z.

Ozturk:2025:OMR

[13] Ercüment Öztürk, Ayfer Dönmez Çavdar, and Tugrul Çavdar. Optimization of mixture ratios of raw materials in thermoplastic hybrid composites based on particle swarm optimization algorithm. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06555-2.

Yin:2025:CPP

[14] Luona Yin and Huaqun Wang. Conditional privacy-preserving spectrum trading scheme based on traceable ring signature for DSS system. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06518-7.

Temp:2025:MMA

[15] Daniel C. Temp, Alexandre A. F. da Costa, Angelo N. C. Vieira, Ester S. Oribes, Ivan M. Lopes, Paulo Silas S. de Souza, Marcelo C. Luizelli, Arthur F. Lorenzon, and Fábio D. Rossi. MAPER: mobility-aware energy-efficient container registry migrations for edge computing infrastructures. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06516-9.

Abdulganiyu:2025:XVX

[16] Oluwadamilare Harazeem Abdulganiyu, Taha Ait Tchakoucht, Yakub Kayode Saheed, and Hilali Alaoui Ahmed. XIDINTFL-VAE: XGBoost-based intrusion detection of imbalance network traffic via class-wise focal loss variational autoencoder. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06552-5.

Xiao:2025:DPF

[17] Huanling Xiao, Xintong Chen, Liuhui Luo, and Cong Lin. A dual-path feature reuse multi-scale network for remote sensing image super-resolution. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06569-w.

Chen:2025:TSPa

[18] Hongwei Chen, Hui Han, Yifan Chen, Zexi Chen, Rong Gao, and Xia Li. A traffic speed prediction algorithm for dynamic spatiotemporal graph convolutional networks based on attention mechanism. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06539-2.

Liu:2025:MNA

[19] Chenou Liu, Kangjian He, Dan Xu, and Hongzhen Shi. MDH-Net: advancing 3D brain MRI registration with multi-stage transformer and dualstream feature refinement hybrid network. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06470-6.

Zhu:2025:IAS

[20] Chaoyang Zhu, Mengxia Wang, Mengxing Guo, Jinxin Deng, Qipei Du, Wei Wei, and Yunxiang Zhang. Innovative approaches to solar energy forecasting: unveiling the power of hybrid models and machine learning algorithms for photovoltaic power optimization. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06504-z.

Gokturk:2025:DDS

[21] Gökhan Göktürk and Kamer Kaya. DiFuseR: a distributed sketch-based influence maximization algorithm for GPUs. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06566-z.

Liu:2025:MTD

[22] Qinxiao Liu, Hangyu Chen, and Fen Zhao. Maritime target detection algorithm based on fusion of visible and infrared images. The Journal of Supercomputing, 81(1):

??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06509-8.

Gorgin:2025:EHA

[23] Saeid Gorgin, Malik Zohaib Nisar, and Jeong-A Lee. Efficient hardware accelerators for k-nearest neighbors classification using most significant digit first arithmetic. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06466-2.

Gaonkar:2025:PCA

[24] Manisha Naik Gaonkar, Veena Thenkanidiyoor, and Aroor Dinesh Dileep. A parallel computing approach to CNNbased QbE–STD using kernel-based matching. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06497-
                                        9.

Nayerifard:2025:RPB

[25] Tahereh Nayerifard, Haleh Amintoosi, and Abbas Ghaemi Bafghi. A robust PRNU-based source camera attribution with convolutional neural networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06579-8.

Makhanov:2025:TGP

[26] Nursultan Makhanov, Ho Duc Nhan, Kok-Seng Wong, and Nguyen Anh Tu. Towards good practice for convolution and attention with PANs in federated medical image classification. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06476-0.

Zhang:2025:UPS

[27] Hongxia Zhang, Hao Li, Zeya Li, and Pengyu Chen. User preference and social relationship-aware recommendations base on a novel light graph convolutional network. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06583-y.

Mao:2025:RMH

[28] Junjie Mao, Hanxiao Shi, and Xiaojun Li. Research on multimodal hate speech detection based on selfattention mechanism feature fusion. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06602-y.

Zhang:2025:OVE

[29] Lei Zhang, Miao Wang, Liqiang Wang, Zijian Chen, and Hong Zhang. Optimizing vehicle edge computing task offloading at intersections: a fuzzy decision-making approach. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06477-z.

Qi-liang:2025:DFC

[30] Wu Qi-liang, Wang Xing, Zhang Tong, Miao Zi-shu, Ye Wei-liang, and Li Hao. DiffREE: feature-conditioned diffusion model for radar echo extrapolation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06577-w.

Jin:2025:AER

[31] Zeyu Jin and Wenjiao Zai. Audiovisual emotion recognition based on bi-layer LSTM and multi-head attention mechanism on RAVDESS dataset. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06582-z.

Yu:2025:YWM

[32] Xiaoyan Yu, Peng Yan, Shaokai Zheng, Qinghan Du, and Daolei Wang. YOLOv8-WTDD: multi-scale defect detection algorithm for wind turbines. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06487-x.

Izadi:2025:NFL

[33] Saadat Izadi and Mahmood Ahmadi. New fusion loss function based on knowledge generation using Gumbel– SoftMax for federated learning. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06593-w.

Huang:2025:AYI

[34] Yuqin Huang, Dezhi Han, Bing Han, and Zhongdai Wu. ADV-YOLO: improved SAR ship detection model based on YOLOv8. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06527-6.

Xie:2025:DRL

[35] Bo Xie and Haixia Cui. Deep reinforcement learning-based dynamical task offloading for mobile edge computing. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06603-x.

Yadav:2025:FFA

[36] Rahul Kumar Yadav, Shashi Prakash Tripathi, and Abhay Kumar Rai. Finding future associations in complex networks using multiple network features. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06544-5.

Liu:2025:RDC

[37] Xing Liu and Bo Gao. Reconstructiondriven contrastive learning for unsupervised skeleton-based human action recognition. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06573-
                                        0.

Luo:2025:EHP

[38] Yongtao Luo, Jie Liu, Chunye Gong, and Tun Li. An efficient heterogeneous parallel password recovery system on MT-3000. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06532-9.

Li:2025:MME

[39] Jianing Li, Wenjing Zhang, and Bing Zhao. 3D medical model encryption based on five-dimensional hyperchaotic systems with 3D Arnold transform and selectable multiple spiral arrangements. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06483-1.

Zhang:2025:EMM

[40] Heng Zhang, Wei Fu, Xiaoming Wang, Dong Li, Danchen Zhu, and Xingwang Su. An efficient model for metal surface defect detection based on attention mechanism and multi-scale feature. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06591-y.

Wang:2025:POH

[41] Haifeng Wang, Wenkang Guo, and Ming Zhang. Performance optimization of heterogeneous computing for large-scale dynamic graph data. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06562-3.

Lin:2025:GII

[42] Lieqing Lin, Qi Zhong, Jiasheng Qiu, and Zhenyu Liang. E-GRACL: an IoT intrusion detection system based on graph neural networks. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06471-5.

Keyhanipour:2025:GIR

[43] Amir Hosein Keyhanipour. Graphinduced rank-aggregation using information fusion operators. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06595-8.

Xiangyu:2025:DFG

[44] Wu Xiangyu, Du Xuehui, Yang Qiantao, Liu Aodi, and Wang Wenjuan. Dynamic fine-grained access control for smart contracts based on improved attribute-based signature. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06570-3.

Zhu:2025:DAC

[45] Sifeng Zhu, Zhaowei Song, Changlong Huang, Hai Zhu, and Rui Qiao. Dependency-aware cache optimization and offloading strategies for intelligent transportation systems. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06596-7.

Chen:2025:MGV

[46] Yuzhong Chen, Liyuan Shi, Jiali Lin, Jingtian Chen, Jiayuan Zhong, and Chen Dong. Multi-granularity visualtextual jointly modeling for aspectlevel multimodal sentiment analysis. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06567-y.

Tian:2025:SOP

[47] Min Tian, Chaoshuai Xu, Xiaoming Wu, Jingshan Pan, Ying Guo, Wei Du, and Zhenguo Wei. Swpmmas: an optimized parallel maxmin ant system algorithm based on the SW26010-pro processor. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06581-0.

Park:2025:UDP

[48] Jung-Heum Park. Unpaired disjoint path covers in bipartite toruslike graphs with edge faults. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06572-1.

Zhou:2025:ELF

[49] Heng Zhou, Qingguo Zhou, Xiaorun Tang, Jun Shen, Binbin Yong, and Yuanming Huang. Electrical load forecasting based on the fusion of multi-scale features extracted by using neural ordinary differential equation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06485-z.

Chen:2025:ECC

[50] Songlin Chen, Weicheng Wang, Xiaoliang Chen, Maolin Zhang, Peng Lu, Xianyong Li, and Yajun Du. Enhancing Chinese comprehension and reasoning for large language models: an efficient LoRA fine-tuning and tree of thoughts framework. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06499-7.

Du:2025:RAM

[51] Pan Du, Xinping Wang, Tiezhi Li, Chang Su, and Zhenyu Li. Resilience analysis of mine ventilation cyber-physical fusion system. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06533-8.

Panda:2025:EDP

[52] Sanjaya Kumar Panda, Thanmayee Pounjula, Bhargavi Ravirala, and David Taniar. An energy, delay and priority-aware task offloading algorithm for fog computing incorporating load balancing. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06557-0.

Rather:2025:AAT

[53] Bilal Ahmad Rather, Muhammad Imran, and Fozia Bashir Farooq. Algebraic analysis of p-type brooms and their application in allergic drugs. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06600-0.

Wang:2025:DMO

[54] Yu Wang, Yongjie Ma, Quanxiu Li, and Yan Zhao. A dynamic multiobjective optimization evolutionary algorithm based on classification of environmental change intensity and collaborative prediction strategy. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06480-4.

Uwimana:2025:STL

[55] Eustache Uwimana, Yatong Zhou, and Ndiaye Mareme Sall. A short-term load demand forecasting: Levenberg– Marquardt (LM), Bayesian regularization (BR), and scaled conjugate gradient (SCG) optimization algorithm analysis. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06513-y.

Zhang:2025:LNV

[56] Hongying Zhang and Jiangbing Zeng. Lightweight network for visibleinfrared person re-identification via self-distillation and multi-granularity information mining. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06543-
6.

Huang:2025:CCD

[57] Feixiong Huang, Yubiao Pan, Huizhen Zhang, and Mingwei Lin. CD-NRocks: computable data nodes with RocksDB to improve the read performance of LSM-tree-based distributed key-value storage systems. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06526-7.

Meng:2025:DMS

[58] Jing Meng, Zhenfang Zhu, Jiangtao Qi, and Huaxiang Zhang. DRSS: a multimodal sentiment analysis approach based on dual representation and self-supervised learning strategy. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06524-9.

Sun:2025:CCC

[59] Zhuoran Sun, Ying Ying Liu, and Parimala Thulasiraman. Cooperative, collaborative, coevolutionary multiobjective optimization on CPU–GPU multi-core. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06525-
8.

Villarrubia:2025:BSC

[60] Jorge Villarrubia, Luis Costero, Francisco D. Igual, and Katzalin Olcoz. Balanced segmentation of CNNs for multi-TPU inference. The Journal of Supercomputing, 81(1):

??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06605-9.

Xiao:2025:GWP

[61] Zhengqing Xiao, Youliang Tian, Changgen Peng, Yangyang Long, and Chuanda Cai. Ghost-Weight protocol: a highly efficient blockchain consensus for IoT. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06565-0.

Li:2025:ASD

[62] Jinze Li, Xiangyu Meng, Zichen Qi, Dong Guo, and Cong Fu. Attack stage detection method based on vector reconstruction error autoencoder and explainable artificial intelligence. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06473-3.

Liang:2025:FSL

[63] Xingzhu Liang, Yachen Xu, Yu e Lin, and Chunjiong Zhang. Federated split learning via dynamic aggregation and homomorphic encryption on non-IID data. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06612-w.

Zeng:2025:AAS

[64] Biqing Zeng, Liangqi Xie, Ruizhe Li, Yongtao Yao, Ruiyuan Li, and Huimin Deng. Aspect-aware semantic feature enhanced networks for multimodal aspect-based sentiment analysis. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06472-4.

Anjali:2025:NSE

[65] Anjali and Anjana Gupta. A novel Shannon entropy-based backward cloud model and cloud K-means clustering. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06528-5.

Cai:2025:AKN

[66] Liang Cai, Shijie Zhao, Fanshuai Meng, and Tianran Zhang. Adaptive K-NN metric classification based on improved Kepler optimization algorithm. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06559-y.

Zhang:2025:NLC

[67] Longfei Zhang, Gang Wang, and Wei Zhou. A novel loop closure detection algorithm based on crossroad scenes. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06488-w.

Ma:2025:SSB

[68] Shichao Ma, Shengfa Miao, Shaowen Yao, Xin Jin, Xing Chu, Qian Yu, Yuling Tian, and Ruoshu Wang. SDHNet: a sampling-based dual-stream hybrid network for long-term time series forecasting. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06495-x.

Hong:2025:CPS

[69] Jiwon Hong, Hyeongmin Kim, Suhyeon Oh, Yerin Im, Hyeonseong Jeong, Hyunmin Kim, Eunkueng Jang, and Kyounggon Kim. Combating phishing and script-based attacks: a novel machine learning framework for improved client-side security. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06551-6.

Chen:2025:MSL

[70] Yilin Chen, Xin Hu, Tao Lu, Lu Zou, and Xiangyun Liao. A multi-scale large kernel attention with U-Net for medical image registration. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06489-9.

Zhang:2025:MTC

[71] Xiufeng Zhang, Jinwei Zhou, and Guobin Qi. Multimodal temporal context network for tracking dynamic changes in emotion. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06484-0.

Nooh:2025:EBC

[72] Sameer Abdullah Nooh. Enhancing beyond 5G connectivity and security: optimizing user-to-multiple AP associations with hybrid deep learning and innovative optimization techniques. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06503-0.

Gao:2025:SPF

[73] Lei Gao, Jingfei Jiang, Jinwei Xu, Weijia Wang, and Pengbo Wu. SAP-FIS: a parallel fuzzy inference system for air combat situation assessment. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06521-y.

Luo:2025:EWS

[74] Huilan Luo and Zhen Zeng. Enhancing weakly supervised semantic segmentation through multi-class token attention learning. The Journal of Supercomputing, 81(1):

??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06618-4.

Zeng:2025:PPP

[75] Jiangfeng Zeng, Xinyu Li, and Xiao Ma. PPTopicPLM: plug-and-play topic-enhanced pre-trained language model for short-text rumor detection. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06549-0.

Ullah:2025:OVE

[76] Ihsan Ullah and Youn-Hee Han. Optimizing vehicular edge computing: graph-based double-DQN approaches for intelligent task offloading. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06599-4.

Chin:2025:DCE

[77] Zi Hau Chin, Vishnu Monn Baskaran, Chee Keong Tan, Ian K. T. Tan, and Timothy T. V. Yap. DAP-CBR: enhancing Bitcoin block propagation efficiency using dynamic compact block relay’s prefilling of transactions. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06468-0.

Shahid:2025:EMR

[78] Maida Shahid, Muhammad Awais Hassan, Faiza Iqbal, Ayesha Altaf, Sayyed Wajihul Husnain Shah, Ana Visiers Elizaincin, and Imran Ashraf. Enhancing movie recommendations using quantum support vector machine (QSVM). The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06501-2.

Sun:2025:CFR

[79] Yu Sun and Qicheng Liu. Collaborative filtering recommendation based on K-nearest neighbor and nonnegative matrix factorization algorithm. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06537-4.

Wang:2025:ATS

[80] Liangliang Wang, Lei Wang, Qiaoyong Jiang, Zhaoqi Wang, Wenqian Zhu, and Zhennan Wang. An adaptive transfer strategy guided by reference vectors for many-objective optimization problems. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06547-2.

Yan:2025:HUS

[81] Huizhen Yan, Fengyang Li, Ming-Chia Hsieh, and Jimmy Ming-Tai Wu. High-utility sequential pattern mining in incremental database. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06568-x.

Li:2025:TEN

[82] Dan Li, Hongbin Xia, and Yuan Liu. Triplet extraction network with dual gating mechanism and dependency-oriented attention. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06575-y.

Ozdag:2025:NHP

[83] Recep Özdag. A novel hybrid path planning method for sweep coverage of multiple UAVs. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06574-z.

Chen:2025:RSD

[84] Yanping Chen, Feifan Ran, Xiaomin Jin, Haizhou Liu, and Zhongmin Wang. Robust service deployment for edge computing in industrial internet with joint profit awareness and multi-server collaboration. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06609-5.

Jing:2025:GCL

[85] Li Jing, Liu Ke, Zhang Lei, Yin Xiaoya, Jia Yuanyuan, and Jia Huinan. Geohash coding location privacy protection scheme based on entropy weight TOPSIS. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06511-0.

Hu:2025:MAM

[86] Chengyu Hu, Jin Liu, Xingye Li, Meijing Li, and Huihua He. MST-ARGCN: modality-squeeze transformer with attentional recurrent graph capsule network for multimodal sentiment analysis. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06588-7.

Sahraneshinsamani:2025:MPP

[87] Nima Sahraneshinsamani, Sandra Catalán, and José R. Herrero. Mixed-precision pre-pivoting strategy for the LU factorization. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06523-w.

Fardad:2025:OPP

[88] Hamidreza Fardad, Faramarz Safi-Esfahani, and Behrang Barekatain.

Optimizing production planning and sequencing in hot strip mills: an approach using multi-objective genetic algorithms. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06469-z.

Latifi:2025:MCB

[89] Faride Latifi, Ramin Nassiri, Mehran Mohsenzadeh, and Hamidreza Mostafaei. A Markov chain-based multi-criteria framework for dynamic cloud service selection using user feedback. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06508-9.

Xia:2025:FOC

[90] Huangzhi Xia, Yifen Ke, Riwei Liao, and Yunqiang Sun. Fractional order calculus enhanced dung beetle optimizer for function global optimization and multilevel threshold medical image segmentation. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06592-x.

Wu:2025:CAM

[91] Xianglin Wu, Tianhao Meng, Jingwei Zhang, Qing Yang, and Jintao Chen. Consensus algorithm for maintaining large-scale accesscontrol views of education data. The Journal of Supercomputing, 81(1):

??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06625-5.

Wu:2025:API

[92] Zheqian Wu and Yingmin Li. An adaptive physics-informed deep learning approach for structural nonlinear response prediction. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06586-9.

Dalal:2025:DUO

[93] Sumit Dalal, Sarika Jain, and Mayank Dave. DepressionFeature: Underlying ontology for user-specific depression analysis. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06585-w.

Brahim:2025:NPR

[94] A. Hadj Brahim, H. Ali Pacha, M. Naim, and A. Ali Pacha. A novel pseudo-random number generator: combining hyperchaotic system and DES algorithm for secure applications. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06639-z.

Alqudsi:2025:IOS

[95] Yunes Alqudsi. Integrated optimization of simultaneous target assignment and path planning for aerial robot swarm. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06620-w.

Tan:2025:MMI

[96] Linlin Tan, Yinghong Cao, Santo Banerjee, and Jun Mou. Multi-medical image protection: compression– encryption scheme based on TLNN and mask cubes. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06624-
6.

Singh:2025:VON

[97] Chandra Prakash Singh, Rohita Yamaganti, and Lokendra Singh Umrao. Variational Onsager Neural Networksbased fair proof-of-reputation consensus for blockchain with transaction prioritisation for smart cities. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06626-4.

He:2025:CVA

[98] Maowei He, Hongxia Zheng, Hanning Chen, Zhixue Wang, Xingguo Liu, Yelin Xia, and Haoyue Wang. A clustering and vector angle-based adaptive evolutionary algorithm for multi-objective optimization with irregular Pareto fronts. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06496-w.

Chen:2025:MAR

[99] Lei Chen, Jieru Hou, Yunpeng Ma, and Yikai Zhao. A modified averageroulette cellular automaton algorithm for optimization tasks. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06561-4.

Bai:2025:CFC

[100] Xuefei Bai, Yongsong Wan, and Weiming Wang. CEPDNet: a fast CNNbased image denoising network using edge computing platform. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06646-0.

Lei:2025:GGT

[101] Jianjun Lei, Sida Chen, and Ying Wang. GenerCTC: a general twostage contrastive training framework for text classification. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06628-2.

Zhang:2025:OBFb

[102] Ran Zhang, Jingguo Bi, Lixiang Li, and Haipeng Peng. An optimal bound for factoring unbalanced RSA moduli by solving generalized implicit factorization problem. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06478-y.

Naz:2025:ETL

[103] Sumera Naz, Aqsa Tasawar, Areej Fatima, Shariq Aziz Butt, and Zhoe Comas Gonzalez. An efficient 2-tuple linguistic cubic q-rung orthopair fuzzy CILOS–TOPSIS method: evaluating the hydrological geographical regions for watershed management in Pakistan. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06505-y.

Yang:2025:MIM

[104] Dingyu Yang, Kangpeng Zheng, Shiyou Qian, Qin Hua, Kaixuan Zhang, Jian Cao, and Guangtao Xue. Mitigating interference of microservices with a scoring mechanism in large-scale clusters. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06534-7.

Lyu:2025:LYL

[105] Yifan Lyu, Tianze Zhang, Xin Li, Aixun Liu, and Gang Shi. LightUAV-YOLO: a lightweight object detection model for unmanned aerial vehicle image. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06611-x.

Rajpopat:2025:CPD

[106] Subodh Rajpopat, Sunil Kumar, and Narinder Singh Punn. Cerebral palsy detection from infant using movements of their salient body parts and a feature fusion model. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06520-z.

He:2025:MMS

[107] Yu He, Kang Zhou, and T. Lifang Tian. Multi-modal scene graph inspired policy for visual navigation. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06541-8.

Ding:2025:NCM

[108] Pengfei Ding, Penghui Geng, and Weiwei Hu. A new controllable multiwing chaotic system: applications in high-security color image encryption. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06635-3.

Sayadi:2025:BPE

[109] Ladan Sayadi, Abdolah Amirany, Mohammad Hossein Moaiyeri, and Somayeh Timarchi. Balancing precision and efficiency: an approximate multiplier with built-in error compensation for error-resilient applications. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06563-2.

Shi:2025:ESM

[110] Leyi Shi, Qihang Yang, Luhan Gao, and Haowei Ge. An ensemble system for machine learning IoT intrusion detection based on enhanced artificial hummingbird algorithm. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06475-1.

Zhou:2025:EPP

[111] Sufang Zhou, Jianing Fan, Ke Yuan, Xiaoyu Du, and Chunfu Jia. Efficient privacy-preserving online medical pre-diagnosis based on blockchain. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06486-y.

Li:2025:AHG

[112] Bin Li, Haoyu Wang, Xaoyu Tan, Qiong Li, Jue Chen, and Xihe Qiu. Adaptive heterogeneous graph reasoning for relational understanding in interconnected systems. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06623-7.

Hu:2025:EMT

[113] Kun Hu and Yuanbin Mo. An efficient multi-threshold image segmentation method for COVID-19 images using reinforcement learningbased enhanced sand cat algorithm. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06498-8.

He:2025:CLN

[114] Qi-Qiao He, Xueyuan Gong, and Yain-Whar Si. Collaborative learning with normalization augmentation for domain generalization in time series classification. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06622-8.

Dai:2025:SSS

[115] Zhi-Qiang Dai, Jie Li, Yang-Jie Cao, and Yong-Xiang Zhang. SALSTM: segmented self-attention long shortterm memory for long-term forecasting. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06493-z.

Ijaz:2025:EET

[116] Samia Ijaz, Saima Gulzar Ahmad, Kashif Ayyub, Ehsan Ullah Munir, and Naeem Ramzan. Energy-efficient time and cost constraint scheduling algorithm using improved multi-objective differential evolution in fog computing. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06550-7.

Qi:2025:EQL

[117] Han Qi, Xinyue Lv, Changqing Gong, and Abdullah Gani. Enhanced quantum long short-term memory by using bidirectional ring variational quantum circuit. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06636-2.

Gu:2025:ALD

[118] Runhe Gu and Luchun Lin. Application of latent Dirichlet allocation and autoencoder to real estate datasets. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06659-9.

Yue:2025:FAB

[119] Xiaohan Yue, Gang Yi, Haoran Si, Haibo Yang, Shi Bai, and Yuan He. A face authentication-based searchable encryption scheme for mobile device. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06554-3.

Zhou:2025:SRC

[120] Sizhong Zhou. Spectral radius and component factors in graphs. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06522-x.

Guha:2025:SAD

[121] Krishnendu Guha. Self-aware decentralized security for real time approximate computing tasks in FPGA-based edge platforms. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06538-3.

Zheng:2025:MGL

[122] Jianxing Zheng, Min Li, Suge Wang, Jian Liao, and Xiaoya Wan. Multigranularity label-aware user interest modeling for news recommendation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06502-1.

Ranjan:2025:EES

[123] Rajeev Ranjan, Raj Anwit, and Prabhat Kumar. Energy efficient and sustainable mobile data collection in internet of things: a variable dimension SSO-based approach. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06630-8.

Gan:2025:MOO

[124] Wei Gan, Hongye Li, and Pengpeng Hao. Many-objective optimization algorithm based on the similarity principle and multi-mechanism collaborative search. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06553-4.

Chen:2025:RRD

[125] Jiang Chen and Weijie Ye. Roaen: reversed dependency graph and orthogonalgating strategy attention-enhanced network for aspect-level sentiment classification. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06542-7.

Tang:2025:CEE

[126] Yiou Tang, Yan Ma, Chunling Xiao, Min Wu, and Guoyuan Zeng. Classification of EEG event-related potentials based on channel attention mechanism. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06627-3. See cor-

rection [405].

Peng:2025:DBN

[127] Jianting Peng, Xuyu Xiang, Jiaohua Qin, and Yun Tan. Dualbranch networks for privacy-preserving cross-modal retrieval in cloud computing. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06643-3.

Merdassi:2025:NLT

[128] Imen Merdassi, Cherif Ghazel, and Leila Saidane. A novel location and time privacy-preserving approach for mobile cloud based on blockchain. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06604-w.

Chen:2025:STG

[129] Guihui Chen, Yuli Wei, Jiao Peng, Xinyu Zheng, Kai Lu, and Zhongbing Li. Spatio-temporal graph neural network based on time series periodic feature fusion for traffic flow prediction. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06653-1.

Wang:2025:DIM

[130] Jingwen Wang, Zhoulin Cao, Chunzhi Xie, Yanli Li, Jia Liu, and Zhisheng Gao. DGN: influence maximization based on deep reinforcement learning. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06621-9.

Zhu:2025:ABS

[131] Chao Zhu and Qiang Ding. Aspectbased sentiment analysis via dual residual networks with sentiment knowledge. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06546-3.

Shi:2025:ILC

[132] Meifeng Shi, Guoyan Jia, and Makoto Yokoo. The improved local cost simulation algorithms based on coalition structure generation for solving distributed constraint optimization problems. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06644-2.

Jha:2025:SHT

[133] Vidyapati Jha and Priyanka Tripathi. Selective hypothesis testing in cognitive IoT sensor network.

The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06515-w.

Sadeghi-Nasab:2025:ODP

[134] Alireza Sadeghi-Nasab and Mohsen Rahmani. Optimizing data privacy: an RFD-based approach to anonymization strategy selection. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06642-4.

Liang:2025:MPP

[135] Bo Liang and Jie Gao. A multi-process parallel clustering algorithm for resource reconfiguration in cloud manufacturing. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06607-7.

Guo:2025:BTD

[136] Yan Ru Guo, Xiang Fei Yang, Xiang Yu Hua, and Jun Pan. Bilateral two-dimensional linear discriminant analysis and its applications. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06660-2.

Wang:2025:FSA

[137] Jie Wang, Chaochao Sun, and Yuan Peng. FedBat: a self-adapting bat algorithm-based federated learning approach. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06514-x.

Wang:2025:SGF

[138] Liqin Wang, Pengcheng Yang, Xu Wang, Zhihong Xu, and Yongfeng Dong. Scene graph fusion and negative sample generation strategy for image-text matching. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06652-
2.

Bi:2025:TTS

[139] Yuandong Bi, Hong Zhu, Jing Shi, and Bin Song. TsCANet: Threestream contrastive adaptive network for cross-domain few-shot learning. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06482-2.

Mostafa:2025:EHE

[140] Reham R. Mostafa, Fatma A. Hashim, Ahmed M. Khedr, Zaher Al Aghbari, Imad Afyouni, Ibrahim Kamel, and Naveed Ahmed. EMGODV-Hop: an efficient range-free-based WSN node localization using an enhanced mountain gazelle optimizer. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06616-6.

Hussain:2025:SIC

[141] Adedoyin A. Hussain and Barakat A. Dawood. A survey on IoT-cloud task in healthcare system. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06629-1.

Tripathi:2025:QLB

[142] Sarvapriya Tripathi, Himanshu Upadhyay, and Jayesh Soni. A quantum LSTM-based approach to cyber threat detection in virtual environment. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06615-7.

Zhang:2025:DCB

[143] Qian Zhang, Xing Zhang, Zhiguang Chu, and Xiang Li. Density clustering-based optimization model for trajectory data publication. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06617-5.

Li:2025:CVQ

[144] Hao Li, Xiaohai He, Shuhua Xiong, Haibo He, and Honggang Chen. A compressed video quality enhancement algorithm based on CNN and transformer hybrid network. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06654-0.

Gavagsaz:2025:TID

[145] Elaheh Gavagsaz and Alireza Souri. Triangle-induced and degree-wise sampling over large graphs in social networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06613-9.

Gautam:2025:IMS

[146] Rahul Kumar Gautam, Anjeneya Swami Kare, and S. Durga Bhavani. Interest maximization in social networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06598-5.

Vlasic:2025:FST

[147] Andrew Vlasic, Hunter Grant, and Salvatore Certo. Feature selection through quantum annealing. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06673-x.

Zhao:2025:OBL

[148] Yanfen Zhao and Hao Liu. Oppositionbased learning Harris hawks optimization with steepest convergence for engineering design problems. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06649-x.

Cao:2025:CED

[149] Zhengjun Cao and Huachen Ye. Comment on “Efficient design of an authenticated key agreement protocol for dew-assisted IoT systems”. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06661-1.

Hamdipour:2025:ARO

[150] Ali Hamdipour, Abdolali Basiri, Mostafa Zaare, and Seyedali Mirjalili. Artificial rabbits optimization algorithm with automatically DBSCAN clustering algorithm to similarity agent update for features selection problems. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06606-8.

Tavares-Silva:2025:ASI

[151] Sthéfano Henrique Mendes Tavares-Silva, Sidney Marlon Lopes-Lima, Ricardo Paranhos-Pinheiro, Liosvaldo Mariano Santiago-Abreu, Rafael Diniz Toscano-Lima, and Sérgio Murilo Maciel Fernandes. Antivirus solution to IoT malware detection with authorial next-generation sandbox. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06506-x.

Bashir:2025:CSF

[152] Zia Bashir, M. G. Abbas Malik, and Sadam Hussain. A computational study of fractional variable-order nonlinear Newton–Leipnik chaotic system with radial basis function network. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06492-0.

Li:2025:MCE

[153] Xiaohong Li, Jin Yao, Peng liu, and Yang Han. MC-CRS: enhanced conversational recommender system based on multi-contrastive learning. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06666-w.

Qi:2025:HEV

[154] Han Qi, Sihui Xiao, Zhuo Liu, Changqing Gong, and Abdullah Gani. A high-efficiency variational quantum classifier for high-dimensional data. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06676-8.

Wei:2025:RCV

[155] Hao Wei, Linchang Zhao, Ruiping Li, and Mu Zhang. RFAConv-CBM-ViT: enhanced vision transformer for metal surface defect detection. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06662-0.

Pan:2025:PPB

[156] Yubiao Pan, Jianing Zhao, Yixiang Cai, Huizhen Zhang, and Mingwei Lin. PMCKV: pipeline-based multi-compactions KV stores to improve the system performance. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06680-y.

Liu:2025:IUC

[157] Yiwen Liu, Xiaoyu Zhang, Jinchao Zhu, and Panlong Tan. Improving underwater camouflage object segmentation with dual-decoder attention network. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06584-x.

Li:2025:PFL

[158] Hongjiao Li, Jiayi Xu, Ming Jin, and Anyang Yin. Personalized federated learning with global information fusion and local knowledge inheritance collaboration. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06529-4.

Xiang:2025:NDT

[159] Qiuhong Xiang, Hongfang Gong, and Cheng Hua. A new discretetime denoising complex neurodynamics applied to dynamic complex generalized inverse matrices. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06601-z.

Chang:2025:GLL

[160] Rui Chang, Gang Liu, Yao Qian, Haojie Tang, Gaoqiang Wang, and Durga Prasad Bavirisetti. GDNet: a low-light image enhancement network based on Ghost–Block and unique image decomposition. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06683-9.

Liu:2025:DRL

[161] Chang Liu, Xue-Feng Feng, Feng Li, Qing-Long Xian, Zhen-Hong Jia, Yu-Hang Wang, and Zong-Dong Du. Deep reinforcement learning combined with transformer to solve the traveling salesman problem. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06691-9.

Bai:2025:MTP

[162] Xiangqi Bai, Lingtao Zhang, Yanyan Feng, Haoran Yan, and Quan Mi. Multivariate temperature prediction model based on CNN–BiLSTM and Random-Forest. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06689-3.

Kim:2025:TBL

[163] Taero Kim, Changdae Oh, Hyeji Hwang, Eunkyeong Lee, Yewon Kim, Yunjeong Choi, Sungjin Kim, Hosik Choi, and Kyungwoo Song. TC-BERT: large-scale language model for Korean technology commercialization documents. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06597-6.

Zhang:2025:ECE

[164] Guozhen Zhang, Zhimin Yue, and Dajin Wang. 1-extra 3-component edge connectivity of modified bubble-sort networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06610-y.

Bouleghlimat:2025:LLR

[165] Imene Bouleghlimat, Souheila Boudouda, and Salima Hacini. LSPP: a leakage-resilient security approach for a cloud-assisted big data. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06657-x.

Mishra:2025:DNB

[166] Deepak Kumar Mishra and Pawan Singh Mehra. DiabeticChain: a novel blockchain approach for patient-centric diabetic data management. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06589-6.

Liu:2025:NFE

[167] Suping Liu and Xiaomin Li. New fuzzy entropy via class-consistent technology with applications to attribute reduction for heterogeneous data. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06631-7.

Si:2025:LLC

[168] Yazhong Si, Chen Li, and Fan Yang. LGASR: latent-content guided adversarial sand-dust image reconstruction strategy. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06638-0.

Zhang:2025:SSS

[169] Meijia Zhang, Junzheng Li, and Shengpeng Yu. Semi-supervised singleimage dehazing based on spatialchannel feature enhancement. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06665-x.

Panda:2025:EDH

[170] Subrat Kumar Panda, Debasis Acharya, Dushmanta Kumar Das, and R. Kumar Rajagopal. An enhanced DV–Hop localization algorithm in wireless sensor networks with variable velocity strategy and human conception optimization. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06641-5.

AbdelAziz:2025:CBT

[171] Nabil M. AbdelAziz, Aliaa A. Ali, Soaad M. Naguib, and Lamiaa S. Fayed. Clustering-based topic modeling for biomedical documents extractive text summarization. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06640-6.

Yu:2025:OCV

[172] Meng-Shiun Yu, Hao-Chun Chang, Chong-Teng Wang, Yu-Wei Tien, Tai-Liang Chen, and Jenq-Kuen Lee. Optimizing computer vision algorithms with TVM on VLIW architecture based on RVV. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06530-x.

Karamaz:2025:DMR

[173] Fatih Karamaz and Faruk Karaaslan. Distance measures of r, s, t-spherical fuzzy sets and their applications in MCGDM based on TOPSIS. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06560-5.

Yuan:2025:CCD

[174] Ruosong Yuan, Wenwen Zhang, Xiaokang Dong, and Wanjun Zhang. Crns: CLIP-driven referring nuclei segmentation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06692-

Nathiya:2025:BEW

[175] R. Nathiya, D. Meiyappan, Savari Prabhu, and Sudeep Stephen. Bandwidth of wk-recursive networks and its sparse matrix computation. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06633-5.

Zhao:2025:SSC

[176] Xu Zhao, Xiaohong Wang, and Bingge Cong. SGSLNet: stratified contextual graph pooling for point cloud segmentation with graph structural learning. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06656-y.

Alsuwat:2025:IMM

[177] Emad Alsuwat and Hatim Alsuwat. An improved multi-modal framework for fake news detection using NLP and Bi– LSTM. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06671-z.

Guo:2025:EGC

[178] Rui Guo, Wen Xiong, Yungang Zhang, and Yanfang Hu. Enhancing game customer churn prediction with a stacked ensemble learning model. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06675-9.

Diao:2025:CNT

[179] Jiarong Diao, Kai Cui, Yuling Huang, Chujin Zhou, Jianqing Li, and Haoyan Song. ChebyshevNet: a novel time series analysis model using Chebyshev polynomial. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06672-y.

Huang:2025:ARC

[180] Zhaoman Huang, Yayu Yang, Mingzu Zhang, and Xing Yang. Assessing reliability in Complete Josephus Cube networks via strongly Menger edge-connectivity. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06564-1.

Kapil:2025:SBH

[181] Gayatri Kapil, Neeraj Kumar, Ashish Kumar Mourya, and Vijay Kumar. Securing big healthcare data using attribute and honey-based encryption in cloud environment. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06535-6.

Taibi:2025:CNC

[182] Salaheddine Taibi, Lyazid Toumi, and Salim Bouamama. Complex network community discovery using fast local move iterated greedy algorithm. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06614-8.

Catalan:2025:ENN

[183] Izan Catalán, José Flich, and Carles Hernández. Exploiting neural networks bit-level redundancy to mitigate the impact of faults at inference. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06693-7.

Parsafar:2025:RLB

[184] Parsa Parsafar. A reinforcement learning-based GWO–RNN approach for energy efficiency in data centers by minimizing virtual machine migration. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06510-1.

Jia:2025:RRL

[185] Chao Jia, Tao Yu, and ZiJian Song. Robust reinforcement learning with augmented state for leveling control of multi-cylinder hydraulic system. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06681-x.

Meng:2025:MMS

[186] Guang-Cong Meng, Yong-Xin Dong, and Yue-Hua Feng. A modified multistep splitting iteration and its variants for computing PageRank. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06669-7.

Alammary:2025:IIP

[187] Ali Saleh Alammary. Investigating the impact of pretraining corpora on the performance of Arabic BERT models. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06698-2.

Xi:2025:IWH

[188] Mengyao Xi and Hao Liu. An improved wild horse optimization algorithm based on reinforcement learning for numerical and engineering optimizations. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06651-3.

Wu:2025:IMN

[189] Danqing Wu and Chao Zhu. Interactive memory networks based on syntactic dependencies for aspectlevel sentiment classification. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06594-9.

Dabbagh:2025:CMM

[190] Arsalan Dabbagh, Mehrzad Karamimanesh, Kourosh Hassanli, and Ebrahim Abiri. CD-MAC: mixedsignal binary/ternary in-memory computing accelerator for power-constrained MAC processing. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06700-
x.

Li:2025:INL

[191] Xue-Ning Li, Fang-Jiong Chen, Ye-Ping Lai, Peng Tang, and Xiao-Jun Liang. ICAT-net: a lightweight neural network with optimized coordinate attention and transformer mechanisms for earthquake detection and phase picking. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06664-y.

Du:2025:TPR

[192] Jianbo Du, Jianjun Zhang, Jie Li, Jiaju Lv, Aijing Sun, Jing Jiang, Pengfei Du, and Jing Bai. Task placement and resource allocation for UAV and edge computing supported transportation systems. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06647-z.

Li:2025:GCR

[193] Xingyu Li and Jinglei Liu. Graph convolutional and random Fourier feature mapping for hyperspectral image clustering. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06696-4.

Yin:2025:CEH

[194] Junqi Yin, Jesse Hines, Emily Herron, Tirthankar Ghosal, Hong Liu, Suzanne Prentice, Vanessa Lama, and Feiyi Wang. chatHPC: Empowering HPC users with large language models. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06637-1.

Mirzaie:2025:ERC

[195] Sara Mirzaie and Omid Bushehrian. Efficient root cause localization in IoT-enabled water distribution networks by hierarchical anomaly analysis. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06716-3.

Wu:2025:HHS

[196] Nengwu Wu, Wenjie Zhao, Yuxiang Chen, Jiahong Xiao, Jin Wang, Wei Liang, Kuan-Ching Li, and Nitin Sukhija. HFSL: heterogeneity split federated learning based on client computing capabilities. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06632-6.

Erdag:2025:JPF

[197] Özgür Erdag, James F. Peters, and Ömür Deveci. The Jacobsthal– Padovan–Fibonacci p-sequence and its application in the concise representation of vibrating systems with dual proximal groups. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06608-6.

Che:2025:ISD

[198] Xingliang Che, Wen Xiong, Xian Zhang, and Xitong Zhang. An integrated static and dynamic graph fusion approach for traffic flow prediction. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06670-0.

Zhang:2025:LMS

[199] Xingpeng Zhang, Jing Xu, Dong He, Kaixin Wang, and Liping Wang. Lightweight multi-scale attention group fusion structure for nuclei segmentation. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06710-9.

He:2025:MFL

[200] Jiangtao He. A multichannel fusion learning model with syntax for Chinese-oriented aspectlevel sentiment classification. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06674-w.

Dhibar:2025:ASI

[201] Sibasish Dhibar and Madhu Jain. ANFIS simulation integrated in FM/FM/1(CV + WV) queue with Bernoulli service interruption and metaheuristic optimization for mathematical model. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06481-
3.

Zhang:2025:OYA

[202] Yunjie Zhang, Guofeng Gao, Yadong Chen, and Zhenjian Yang. ODD-YOLOv8: an algorithm for small object detection in UAV imagery. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06703-8. See cor-

rection [384].

Mohapatra:2025:SBH

[203] J. Mohapatra, L. Govindarao, and S. Priyadarshana. A splitting based higher-order numerical scheme for 2D time-dependent singularly perturbed reaction-diffusion problems. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06479-x.

Liu:2025:EQS

[204] Yutang Liu, Qin Zhang, and Ruoxia Li. Exponentially quasi-synchronization control of quaternion-valued memristive neural networks: matrix measure strategies and Frobenius norm methods. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06699-1.

Zhang:2025:IER

[205] Yunyi Zhang, Ye Du, Wei He, and Yu Tang. Inference of evidence reasoning rule with Gaussian distribution reliability and its application in safety assessment. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06648-y.

Jahrmoi:2025:FAR

[206] Mohammad Amin Ghasvari Jahrmoi, Mehrdad Ashtiani, and Fatemeh Bakhshi. FaasFlows: an approach for reducing vendor lock-in and response time in serverless workflows. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06706-5.

Tiwari:2025:LLL

[207] Sambhavi Tiwari, Manas Gogoi, Shekhar Verma, and Krishna Pratap Singh. Learning to learn: a lightweight meta-learning approach with indispensable connections. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06701-w.

Zang:2025:DSA

[208] XiuBo Zang, HongBin Xia, and Yuan Liu. Diffusion social augmentation for social recommendation. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06695-5.

Saeedi:2025:EEC

[209] Ahmad Saeedi, Marjan Kuchaki Rafsanjani, and Samaneh Yazdani. Energy efficient clustering in IoT-based wireless sensor networks using binary whale optimization algorithm and fuzzy inference system. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06556-1.

Huang:2025:HCT

[210] Cheng-Ta Huang, Njabulo Sinethemba Shongwe, Hao-Yu Weng, Chi-Yao Weng, and Shiva Prasad Sirmulwar. Hybrid coding table-based semireversible data hiding using least significant bits and encryption. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06712-7.

Zhu:2025:PID

[211] Wen-Han Zhu, Rong-Xia Hao, Jou-Ming Chang, and Jaeun Lee. Packing internally disjoint Steiner paths of data center networks. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06545-4.

Mahapatra:2025:ADS

[212] Ajit Kumar Mahapatra, Nibedan Panda, and Binod Kumar Pattanayak. Adaptive dimensional searchbased orthogonal experimentation SSA (ADOX-SSA) for training RBF neural network and optimal feature selection. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06507-w.

Zuo:2025:PDC

[213] Zhibin Zuo, Demin Wang, Xiaowei Nie, Xiaoduo Pan, Miaolei Deng, and Mimi Ma. PDCF-DRL: a contention window backoff scheme based on deep reinforcement learning for differentiating access categories. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06634-4.

Castello:2025:EGM

[214] Adrián Castelló, Héctor Martínez, Sandra Catalán, Francisco D. Igual, and Enrique S. Quintana-Ortí. Experienceguided, mixed-precision matrix multiplication with Apache TVM for ARM processors. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06720-
7.

Ga:2025:EGB

[215] Sangmin Ga, Paul Hyunbin Cho, Gordon Euhyun Moon, and Sungwon Jung. Efficient GNN-based social recommender systems through social graph refinement. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06682-w.

Guo:2025:DSC

[216] Zhifeng Guo, Zijun Chen, Xue Sun, and Wenyuan Liu. H-dominant skyline community search in multi-valued networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06679-5.

Suppalap:2025:SSV

[217] Siwakon Suppalap, Dawrawee Makmuang, Vipavee Damminsed, and Rabian Wangkeeree. Smooth support vector machine with rescaled generalized pinball loss for classification. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06697-3.

Vincy:2025:MLA

[218] G. Caroline Vincy and M. David Raj. Minimum linear arrangement and embedding of (K_{11} − C_{11})^{n} into specific graphs with optimal wirelength calculation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06519-6.

Sezer:2025:PNP

[219] Bora Bugra Sezer and Sedat Akleylek. PPLBB: a novel privacy-preserving lattice-based blockchain platform in IoMT. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06650-4.

Liu:2025:MGE

[220] Huilin Liu, Yuhao Wu, Zihan Tang, Xiaolong Li, Shuzhi Su, Xingzhu Liang, and Pengfei Zhang. Multi-granularity enhanced feature learning for visibleinfrared person re-identification. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06731-4.

Zhang:2025:RRM

[221] Jing Zhang, Bing Li, Yao Zhang, Yuguang Xu, and Hongan Li. Research on the recommendation method of urban location point of interest based on DTCN–EFFN-Transformer. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06742-1.

Song:2025:YGR

[222] Limei Song, Shikun Lu, Yu Tong, and Fengyi Han. YOLOv8s-GSW: a real-time detection model for hexagonal barbed wire breakpoints. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06738-x.

Sarin:2025:GLQ

[223] Saket Sarin, Sunil K. Singh, Sudhakar Kumar, and Shivam Goyal. Geospectra: leveraging quantum-SAR and deep learning for enhanced geolocation in urban environments. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06619-3.

Yue:2025:QLB

[224] Xingqi Yue, Lang Li, Qiuping Li, Jiahao Xiang, and Zhiwen Hu. QLW: a lightweight block cipher with high diffusion. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06707-4.

Ren:2025:SKU

[225] Zhengwei Ren, Xiaojuan Li, Pei He, Rongwei Yu, Shiwei Xu, Yan Tong, and Jinshan Tang. In-situ key update and minimal key set of encrypted outsourced data under binary key-derivation tree. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06684-8.

Mohammadi:2025:EHA

[226] Jamal Mohammadi, Mahmoud Shirazi, and Mehdi Kargahi. Energyharvesting-aware federated scheduling of parallel real-time tasks. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06685-7.

Tseng:2025:ESC

[227] Tzu-Lan Tseng, Wen-Yau Liang, and Hung-Lin Huang. Exploring a social-curiosity-based algorithm for group recommender systems. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06590-z.

Tian:2025:RNR

[228] Ran Tian, Shanwei Li, and Guoying Yang. Retraction note: Research on a distributed auto-negotiation model based on Stackelberg game theory. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06777-4.

Pu:2025:CBD

[229] Luoxi Pu and Zhi Quan. A CIMSbased decision support system for ecommerce: boosting real-time and effective digital marketing. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06730-5.

Xiong:2025:PZB

[230] Jie Xiong, Dongsheng Wang, Jian Yin, and Runfang Wu. Precise Z-Block positioning and dimension measurement using improved Canny edge detection and sub-pixel contour fitting. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06769-4.

Gao:2025:MMM

[231] Jiaqi Gao, Mingrui Fan, Yaru He, Daoqi Han, Yueming Lu, and Yaojun Qiao. MACAE: memory moduleassisted convolutional autoencoder for intrusion detection in IoT networks. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06704-7.

Wang:2025:TMI

[232] Liang Wang, Yilin Li, and Lina Zuo. Trust management for IoT devices based on federated learning and blockchain. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06715-
                                        4.

Wang:2025:LLP

[233] Shanshan Wang, Guofeng Zhao, Chuan Xu, Zhenzhen Han, and Shui Yu. LPQAA: a lightweight post-quantum access authentication scheme for satellite network. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06687-5.

Zhou:2025:ABR

[234] Wang Zhou, Ying Tian, Amin Ul Haq, and Sultan Ahmad. An autoencoder-based recommendation framework toward cold start problem. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06721-6.

Zhang:2025:PCM

[235] Hao Zhang, Jiashu Wang, Yunhao Zhao, Lianjie Wang, Wenyin Zhang, Yeh-Cheng Chen, and Neal Xiong. Pocket convolution Mamba for brain tumor segmentation. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06732-3.

Wang:2025:WFC

[236] Yinuo Wang, Huanqi Zheng, Qiang Wu, Shengkun Yang, and Yucheng Zhou. White-faced capuchin optimizer: a new bionic metaheuristic algorithm for solving optimization problems. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06713-6.

Swain:2025:IVM

[237] Smruti Rekha Swain, Anshu Parashar, Ashutosh Kumar Singh, and Chung Nan Lee. An intelligent virtual machine allocation optimization model for energyefficient and reliable cloud environment. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06734-1.

Qin:2025:ICR

[238] Xiwen Qin, Siqi Zhang, Xiaogang Dong, Tingru Luo, Hongyu Shi, and Liping Yuan. An improved conditional relevance and weighted redundancy feature selection method for gene expression data. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06714-5.

Zhang:2025:MCG

[239] Hongwei Zhang, Zhidong Lu, Xiwei Chen, Shuai Lu, and Le Yao. Masked contrastive generative adversarial network for defect detection of yarn-dyed fabric. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06711-8.

Du:2025:RBI

[240] Junkun Du, Mingqing Wang, Xin Wang, Zhipeng Yang, Xiaojie Li, and Xi Wu. Reference-based image super-resolution with attention extraction and pooling of residuals. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06587-8.

He:2025:DLB

[241] Lian ge He, Jun Song, Yan Zhang, and Xin yang Wu. Deep learningbased electric vehicle transmission system speed ratio prediction research. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06748-9.

Ge:2025:CCD

[242] Mengjiao Ge, Wen Su, Jinfeng Gao, and Guoqiang Jia. CDMANet: central difference mutual attention network for RGB-D semantic segmentation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06760-z.

Boujelben:2025:TAR

[243] Abir Boujelben and Ikram Amous. Treating anomalies in rule bases associated to ontologies. The Journal of Supercomputing, 81(1):

??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06688-4.

Hu:2025:RUI

[244] Kuo-Jui Hu, Yi-Tsung Pan, Li-Wei Jiang, Sin-Der Lee, and Sheng-Long Kao. A robust underwater image enhancement algorithm. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06719-0.

Wang:2025:PCP

[245] Jiaqi Wang and Dongqin Cheng. The 3-path-connectivity of pancake graphs. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06702-9.

Kamatchi:2025:SEP

[246] K. Kamatchi and E. Uma. Securing the edge: privacy-preserving federated learning for insider threats in IoT networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06752-z.

NamvariTazehkand:2025:MSE

[247] Leila NamvariTazehkand and Saeid Pashazadeh. Modeling, simulation, and evaluation of causal order assurance techniques in causal broadcast algorithms using timed colored Petri nets. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06754-x.

Zhao:2025:SMO

[248] Ming Zhao. Synthetic minority oversampling technique based on natural neighborhood graph with subgraph cores for class-imbalanced classification. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06655-z.

Li:2025:PLM

[249] Xuefeng Li, Chen Chen, Jian Wei, Chensu Zhao, and Xiaqiong Fan. Precision localization method for fake news detection across multiple domains. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06776-5.

Huang:2025:RNC

[250] Rongbing Huang, Muhammad Farhan Hanif, Muhammad Kamran Siddiqui, Mazhar Hussain, and Muhammad Faisal Hanif. Retraction note: On connection number-based topological indices and entropy measures for triangular γ-graphyne network. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06781-8.

Wang:2025:MIB

[251] Wen-Chuan Wang, Wei-Can Tian, Kwok-Wing Chau, and Hongfei Zang. MSBES: an improved bald eagle search algorithm with multi- strategy fusion for engineering design and water management problems. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06727-0.

Madani:2025:FEH

[252] Bachir Madani, Mohamed Salah Azzaz, Said Sadoudi, and Redouane Kaibou. Fast and efficient hardware architecture of Chebyshev polynomials algorithm for resisting to side channel attacks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06761-y.

Pacheco:2025:SLR

[253] A. Lopez Pacheco and J. Aguilar. Systematic literature review on quantum applications in nanotechnology. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06747-w.

Khah:2025:HML

[254] Yashar Pourardebil Khah, Mirsaeid Hosseini Shirvani, and Homayun Motameni. A hybrid machine learning approach for feature selection in designing intrusion detection systems (IDS) model for distributed computing networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06677-7. See cor-

rection [419].

Ling:2025:MOP

[255] Qinghua Ling, Zexu Li, Wenkai Liu, Jinlong Shi, and Fei Han. Multiobjective particle swarm optimization based on particle contribution and mutual information for feature selection method. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06762-x.

Wei:2025:PLR

[256] Xing Wei, Jiong Xia, Cang Liu, Qi wen He, Jun Chen, Zhen Wei, Chong Zhao, Fan Yang, and Yang Lu. Proposal-level reliable featureguided contrastive learning for SFOD. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06773-8.

Zarei:2025:OES

[257] Sevda Zarei, Sadoon Azizi, and Awder Ahmed. Optimizing edge server placement and load distribution in mobile edge computing using ACO and heuristic algorithms. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06780-9.

Chen:2025:CCR

[258] Fei Chen, Carter Zhang, and Bo Ning. CRKG: combining retrieval knowledge with generative language models. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06728-z.

Nie:2025:CSK

[259] Shiqiang Nie, Jie Niu, Fangxing Yu, Jianqiang Ma, Xingxing Zhu, and Weiguo Wu. Constructing a scalable key-value store engine on multidisk system. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06725-2.

Tan:2025:HPR

[260] Kai Tan, Dongyang Zhan, Lin Ye, Hongli Zhang, Binxing Fang, and Zhihong Tian. A high-performance realtime container file monitoring approach based on virtual machine introspection. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06779-2.

Xie:2025:LIY

[261] Yinggang Xie and Yanwei Zhao. Lightweight improved YOLOv5 algorithm for PCB defect detection. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06739-w.

Wang:2025:IBF

[262] Jingyuan Wang, Yuan Zhao, Wenyan Wang, and Ziheng Wu. Improving bearing fault diagnosis method based on the fusion of time-frequency diagram and a novel vision transformer. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06793-4.

Ardakani:2025:ICS

[263] Saeid Pourroostaei Ardakani, Jianwei Hu, Jing Zhang, Kaifeng Jin, Tianhong Cai, Anthony Graham Bellotti, and Xiuping Hua. Identifying crowdfunding storytellers who deliver successful projects: a machine learning approach. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06785-4.

Zhao:2025:AOE

[264] Yongpeng Zhao, Shengwei Fu, Langlang Zhang, and Haisong Huang. Aitken optimizer: an efficient optimization algorithm based on the Aitken acceleration method. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06709-2.

Shirini:2025:ILR

[265] Kimia Shirini, Meysam Balaneshin Kordan, and Sina Samadi Gharehveran. Impact of learning rate and epochs on LSTM model performance: a study of chlorophyll-a concentrations in the Marmara Sea. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06806-2.

Ren:2025:PPT

[266] Qingxin Ren and Feng Feng. PID parameter tuning optimization based on multi-strategy fusion improved zebra optimization algorithm. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06548-1.

Wang:2025:DNN

[267] Xiaolong Wang, Hedi An, Jinsong Zhang, Dongya Huang, and Junxian Wen. DPSMUNet: a new network based on a dual-pooling selfattention module for carotid artery plaque segmentation in ultrasound images. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06770-x.

Lin:2025:ATD

[268] Zhifeng Lin, Yilu Chen, Yanyue Xie, Chuandong Chen, Jun Yu, and Jianli Chen. An analytical timing-driven placer for modern heterogeneous FP-GAs. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06755-w.

Zhu:2025:FSS

[269] Suxia Zhu, Yunmeng Wang, and Guanglu Sun. Federated semisupervised learning based on truncated Gaussian aggregation. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06798-z.

Lloria:2025:OMW

[270] Diego Lloria, Sandra Roger, Germán León, José M. Badía, Carmen Botella-Mascarell, and Jose A. Belloch. Optimizing millimeter wave MIMO channel estimation through GPU-based edge artificial intelligence. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06795-2.

Khan:2025:BIN

[271] Abdullah Ayub Khan, Asif Ali Laghari, Abdullah M. Baqasah, Rex Bacarra, Roobaea Alroobaea, Majed Alsafyani, and Jamil Abedalrahim Jamil Alsayaydeh. BDLT-IoMT — a novel architecture: SVM machine learning for robust and secure data processing in Internet of Medical Things with blockchain cybersecurity. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06782-7.

Yu:2025:ADS

[272] Dongxian Yu, Xiaoyu Zhou, Ali Noorian, and Mehdi Hazratifard. An AIdriven social media recommender system leveraging smartphone and IoT data. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06722-5. See cor-

rection [383].

Li:2025:PDP

[273] Yifan Li, Shuming Zhou, and Eddie Cheng. Paired 2-disjoint path covers of balanced hypercube under the partitioned edge fault model. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06736-z.

Zaminkar:2025:IKM

[274] Mina Zaminkar. An improved Kmeans and DPC-empowered clustering approach for efficient routing in the FANET. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06724-3.

Zhang:2025:DRL

[275] Lili Zhang, Ke Zhang, Kang Yang, Wei Wei, Jing Li, Hongxin Tan, Pei Yu, Yucheng Han, and Xudong Yang. Driving risks from light pollution: an improved YOLOv8 detection network for high beam vehicle image recognition. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06809-z.

Tang:2025:PPU

[276] Can Tang, Tao Peng, Xingxing Xie, and Junhu Peng. 3D path planning of unmanned ground vehicles based on improved DDQN. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06690-w.

Yang:2025:HPI

[277] Hao Yang, Junyang Yu, and Rui Zhai. High-precision intrusion detection for cybersecurity communications based on multi-scale convolutional neural networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06737-y.

Li:2025:EBN

[278] Shiyong Li, Huan Liu, Wenzhe Li, and Wei Sun. Edge bank: a novel resource pricing and management system for edge service provider. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06578-9.

Yuxin:2025:MMH

[279] Zhang Yuxin, Zhao Enjiao, Liang Hong, and Zhou Wentao. MATD3 with multiple heterogeneous sub-networks for multi-agent encirclement-combat task. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06756-9.

Kamal:2025:EDN

[280] Sadia Kamal, Parth Sharma, P. K. Gupta, Mohammad Khubeb Siddiqui, Ankush Singh, and Abhijit Dutt. DVTXAI : a novel deep vision transformer with an explainable AI-based framework and its application in agriculture. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06494-y.

Li:2025:UUS

[281] Junyan Li, Wenyong Dong, and Xuewen Gui. uTransformer: unified spatial-temporal transformer with external factors for traffic flow forecasting. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06774-7.

Alioghli:2025:EMT

[282] Abdul Amir Alioghli and Feyza Yildirim Okay. Enhancing multivariate timeseries anomaly detection with positional encoding mechanisms in transformers. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06694-6.

Camara:2025:AAS

[283] Jesús Cámara, Javier Cuenca, Victor Galindo, Arturo Vicente, and Murilo Boratto. An autotuning approach to select the inter-GPU communication library on heterogeneous systems. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06794-3.

MohammedJany:2025:DMC

[284] Shaik MohammedJany, Chandra Bhushana Rao Killi, Shaik Rafi, and Syed Rizwana.

Detecting multimodal cyber-bullying behaviour in social-media using deep learning techniques. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06772-
9.

Saini:2025:VAA

[285] Jatinderkumar R. Saini and Shraddha Vaidya. A veracity assessment algorithm for classification of healthcare information using feature bag mash-up approach. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06500-3.

Zhang:2025:OBFa

[286] Bolei Zhang and Lifa Wu. Online budget-feasible mobile crowdsensing with constrained reinforcement learning. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06767-6.

Sun:2025:PCC

[287] Zhuoran Sun, Ying Ying Liu, and Parimala Thulasiraman. Publisher correction: Cooperative, collaborative, coevolutionary multi-objective optimization on CPU–GPU multicore. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06803-5.

Castorena:2025:ECD

[288] Carlos Castorena, Jesus Lopez-Ballester, Juan A. De Rus, Maximo Cobos, and Francesc J. Ferri. Edge computing for driving safety: evaluating deep learning models for cost-effective sound event detection. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06796-1.

Huang:2025:GAI

[289] Haojing Huang, Jiajun Li, Fei Lu, and Jianxin Li. Generative adversarial imitation learning computing task offloading scheme for optimizing of generated sample utilization and system overhead. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06744-z.

Wang:2025:VDA

[290] Yuhai Wang, Shuobo Xu, Peng Wang, Lele Liu, YanShun Li, and Ze Song. Vehicle detection algorithm based on improved RT–DETR. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06766-7.

Ouamri:2025:CSS

[291] Mohamed Amine Ouamri, Turki Alharbi, Daljeet Singh, and Zenadji Sylia. A comprehensive survey on software-defined wide area network (SD-WAN): principles, opportunities and future challenges. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06718-1.

Mat:2025:ESB

[292] Abdullah Ugur Mat and Ayse Nurdan Saran. Enhancing session-based trip recommendations using matrix factorization: a study on algorithm efficiency and resource utilization. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06726-1.

Li:2025:MOF

[293] Yuanhe Li, Wenjian Zhong, and Yuanqing Wu. Multi-objective flexible job-shop scheduling via graph attention network and reinforcement learning. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06741-2.

Ahmad:2025:BFS

[294] Waseem Ahmad, Aurang Zeb, Muhammad Asif, and Muzhou Hou. Bipolar fuzzy soft hamacher aggregations operators and their application in triage procedure for handling emergency earthquake disaster. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06757-8.

Salari-Hamzehkhani:2025:IID

[295] Behnam Salari-Hamzehkhani, Mehdi Akbari, and Faramarz Safi-Esfahani. Introducing an improved deep reinforcement learning algorithm for task scheduling in cloud computing. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06668-8.

Torki:2025:SFV

[296] Omid Torki, Hamid Mala, and Maede Ashouri-Talouki. Secure fully-verifiable outsourcing of modular exponentiation: two servers check each other. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06745-y.

Alasiry:2025:SVL

[297] Areej Alasiry and Mohammed Qayyum. A smart vista-lite system for anomaly detection and motion prediction for video surveillance in vibrant urban settings. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06753-y. See cor-

rection [404].

Ghosh:2025:ECS

[298] Saurav Ghosh, Kanyaka Chakraborty, Piyali Bagchi Khatua, and Utpal Biswas. Efficient charging schedules in a rechargeable wireless sensor network with multiple chargers. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06804-4.

Zhang:2025:MMC

[299] Yaozeng Zhang, Jing Ma, and Yuguang Jia. MCAN: multimodal cross-aware network for fake news detection by extracting semantic-physical feature consistency. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06815-1.

Saberali:2025:EWB

[300] Behzad Saberali, Kai Zhang, Farzad Saberali, Fatna Adinani Said, and Lu Yang. Enhanced well-based surrogate reservoir modeling with integrated streamlines simulation data. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06686-6.

Trivedi:2025:EAA

[301] Chandan Trivedi, Keyur Parmar, and Udai Pratap Rao. ALMASH : an anonymity-based lightweight mutual authentication scheme for Internet of Healthcare Things. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06801-7.

Emami:2025:ALS

[302] Hojjat Emami. An adaptive local search-based arithmetic optimization algorithm for unmanned aerial vehicle placement. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06812-4.

Tang:2025:YSO

[303] Qiang Tang, Chang Su, Yuan Tian, Shibin Zhao, Kai Yang, Wei Hao, Xubin Feng, and Meilin Xie. YOLO-SS: optimizing YOLO for enhanced small object detection in remote sensing imagery. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06765-8.

Liu:2025:TTF

[304] Chen Liu, Tianhao Yu, Xianghong Zhou, Lixin Zhou, and Xiaoyu Gong. TSESRec: A transformerfacilitated set extension model for session-based recommendation. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06814-2.

Yang:2025:LLV

[305] Gaoming Yang, Chenlong Yu, Xiujun Wang, Xianjin Fang, and Ji Zhang. LVAST: a lightweight vision transformer for effective arbitrary style transfer. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06787-2.

Yang:2025:RID

[306] Yixuan Yang, Mingrong Dong, Kai Zeng, and Tao Shen. Research on the improvement of domain generalization by the fusion of invariant features and sharpness-aware minimization. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06797-0.

Wang:2025:RDL

[307] Yiqiao Wang, Jinling Chen, Bo Yang, Yu Chen, Yanlin Su, and Rong Liu. RT-DETRmg: a lightweight realtime detection model for small traffic signs. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06800-8.

Oviesi:2025:QNN

[308] Safura Oviesi, Mohamad Jafar Tarokh, and Mohamad kazem Momeni. Quantum neural network-assisted learning for small medical datasets: a case study in emphysema detection. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06740-3.

Yue:2025:NFS

[309] Zhongzheng Yue, Fei Wen, and Zhijun Li. Neighbor full sum distinguishing total coloring of planar graphs. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06825-z.

Bhattacharya:2025:SVS

[310] Tathagata Bhattacharya, Adithya Vardhan Peddi, Srikanth Ponaganti, and Sai Teja Veeramalla. A survey on various security protocols of edge computing. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06678-6.

Zhang:2025:TTS

[311] Haiping Zhang, Haixiang Lin, Dongjing Wang, Dongyang Xu, Fuxing Zhou, Liming Guan, Dongjing Yu, and Xujian Fang. TSCANet: a two-stream context aggregation network for weaklysupervised temporal action localization. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06810-6.

Yang:2025:GGF

[312] Qimeng Yang, Yi Liu, Lanlan Lu, and Lei Liu. GFIDF: gradual fusion intent detection framework. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06708-3.

Shahba:2025:DFW

[313] Leyla Shahba, Ahmad Heidary-Sharifabad, and Mohammadreza Mollahoseini Ardakani. Detection of fake web pages and phishing attacks with rabbit optimization algorithm. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06658-w.

Khan:2025:MMA

[314] Ajmal Khan, Naveed Iqbal, Zeeshan Kaleem, Zul Qarnain, and Mohammed M. Bait-Suwailam. A multimodel approach for predicting electric vehicle specifications and energy consumption using machine learning. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06820-4.

Guerrero-Pantoja:2025:AUM

[315] David Guerrero-Pantoja, Erick Pautsch, Clara Almeida, Silvio Rizzi, George K. Thiruvathukal, and Maria Pantoja. Accelerating uncertainty methods for distributed deep learning on novel architectures. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06818-y.

Manchala:2025:NSP

[316] Pravali Manchala and Manjubala Bisi. A novel source project and optimized training data selection approach for cross-project fault prediction. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06750-1.

Mnasri:2025:HBF

[317] Sami Mnasri, Dorsaf Salah, and Hanen Idoudi. A hybrid blockchain and federated learning attention-based BERT transformer framework for medical records management. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06816-0.

Mou:2025:PSL

[318] Jianhui Mou, Jian Wang, Yangwei Wang, Ming Xiang, and Lihua Zhang. Pattern synthesis of linear antenna-array for high gain and low sidelobe level based on sand cat swarm optimization algorithm. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06763-w.

Tang:2025:MME

[319] Jiajia Tang, Binbin Ni, Yutao Yang, Yu Ding, and Wanzeng Kong. MECG: modality-enhanced convolutional graph for unbalanced multimodal representations. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06729-y.

Ma:2025:TSL

[320] Fei Ma, Aihua Hou, Feixia Yang, and Guangxian Xu. Tensor subspace learning and folded-concave function regularization for hyperspectral anomaly detection. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06791-6.

Asiamah:2025:SEL

[321] Emmanuel Acheampong Asiamah, Nana Kwadwo Akrasi-Mensah, Prince Odame, Eliel Keelson, Andrew Selasi Agbemenu, Eric Tutu Tchao, Mohammed Al-Khalidi, and Griffith Selorm Klogo. A storage-efficient learned indexing for blockchain systems using a sliding window search enhanced online gradient descent. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06805-3.

Khomami:2025:CGGa

[322] Mohammad Mehdi Daliri Khomami, Mohammad Reza Meybodi, and Alireza Rezvanian. A cellular Goore game-based algorithm for finding the shortest path in stochastic multi-layer graphs. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06786-3.

Li:2025:IPC

[323] Zhe Li, Jinsong Wang, and Yi Li. An improved PBFT consensus algorithm based on reputation and gaming. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06822-2.

Zhang:2025:EVV

[324] Weilin Zhang, Lihua You, Hechao Liu, and Xiaona Fang. The expected values and variances for degree-based topological indices in random spiro chains. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06808-0.

Chen:2025:SRN

[325] Zeqiang Chen, Zhiqing Li, Xu Tang, Lai Chen, and Nengcheng Chen. STP-Net: a recurrent neural network for spatiotemporal processes predictive learning. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06823-1.

Zhu:2025:PPF

[326] Libo Zhu and Xiang Chen. Privacy protection in federated learning: a study on the combined strategy of local and global differential privacy. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06845-9.

Wang:2025:GAE

[327] Fei Wang, Ming Ju, Xianxun Zhu, Qiuyu Zhu, Haiquan Wang, Chunhua Qian, and Rui Wang. A geometric algebra-enhanced network for skin lesion detection with diagnostic prior. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06833-z.

Ramesh:2025:UAB

[328] Parameswaran Ramesh, P. T. V. Bhuvaneswari, V. S. Dhanushree, G. Gokul, and S. Sahana. User associationbased load balancing using reinforcement learning in 5G heterogeneous networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06788-1.

Rastogi:2025:PAU

[329] Eshita Rastogi, Mukesh Kumar Maheshwari, and Ayush Rastogi. Performance analysis of user power consumption in NR-unlicensed DRX. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06839-7.

Wang:2025:OMC

[330] Jian wei Wang, Qing Zhang, and Cheng sheng Pan. Optimization method of C2 system architecture based on AL-CARO. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06768-5.

Li:2025:AUC

[331] Hongzhi Li, Zhanghao Ren, Guoqing Zhu, and Jiaxi Wang. AE-UNet: a composite lung CT image segmentation framework using attention mechanism and edge detection. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06874-4.

Krishna:2025:CCI

[332] Charu Krishna, Divya Kumar, and Dharmender Singh Kushwaha. CovMed-Care: confluence of internet of things, blockchain and machine learning for remote monitoring system of pandemic patients. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06751-0.

Du:2025:RAB

[333] Yu Du, Liwei Yang, and Yuchuan Luo. Resource allocation based on optimized cellular network AP layout for visible light communication heterogeneous network. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06834-y.

He:2025:PCE

[334] Qi He, Yan Wang, Jianxi Fan, and Baolei Cheng. Parallel construction of edge-independent spanning trees in complete Josephus cubes. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06784-5.

Patil:2025:PSC

[335] Rajendra Patil, Sivaanandh Muneeswaran, Vinay Sachidananda, Peng Hongyi, and Mohan Gurusamy. PRI-ORITI: scoring and categorizationbased threat prioritization. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06465-3.

Ahmad:2025:DDP

[336] Farooq Ahmad, Xinfeng Zhang, Zifang Tang, Fahad Sabah, Muhammad Azam, and Raheem Sarwar. Deep deterministic policy gradients with a selfadaptive reward mechanism for image retrieval. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06764-9.

Cakici:2025:DRT

[337] Muhammed Emir Çakici, Feyza Yildirim Okay, and Suat Özdemir. Deepat: a real-time deep learning based model for aircraft tracking system. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06759-6.

Lu:2025:SIR

[338] Zhengyang Lu, Weifan Wang, Tianhao Guo, and Feng Wang. Singleimage reflection removal via selfsupervised diffusion models. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06837-9.

Shang:2025:GIM

[339] Mengying Shang, Mengnan Tian, and Xinduan Wang. GA-DE: an integrated meta-heuristic approach for optimizing feedforward neural networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06799-y.

Ohue:2025:IMM

[340] Masahito Ohue, Nobuaki Yasuo, and Masami Takata. Innovations in mathematical modeling, AI, and optimization techniques. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06861-9.

Yu:2025:DPW

[341] Bengong Yu, Chengwei Cao, and Ying Yang. Dynamic position weighting aspect-focused graph convolutional network for aspect-based sentiment analysis. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06783-6.

Kia:2025:HAR

[342] Zohre Kia, Meisam Yadollahzaeh-Tabari, and Homayun Motameni. Human activity recognition by bodyworn sensor data using bi-directional generative adversarial networks and frequency analysis techniques. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06743-0.

Lin:2025:RRI

[343] Xinwei Lin, Yubiao Pan, Wenjuan Feng, Huizhen Zhang, and Mingwei Lin. RIOKV: reducing iterator overhead for efficient short-range query in LSM-tree-based key–value stores. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06735-0.

Zhang:2025:NHA

[344] Yanjun Zhang, Xin Guo, Hongchen Guo, and Yichen Zhang. A new hardware architecture of highperformance real-time texture classification system based on FPGA. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06705-6.

Liu:2025:CPF

[345] Shiguo Liu, Dejian Wei, Junzhong Zhang, Xurui Ji, and Hui Cao. ConvBiFuseNet: a parallel fusion model with routing attention for MRI brain tumor classification. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06758-7.

Giveki:2025:SIR

[346] Davar Giveki and Sajad Esfandyari. Semantic image representation for image recognition and retrieval using multilayer variational auto-encoder, InceptionNet and low-level image features. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06792-5.

Sinha:2025:BCA

[347] Bam Bahadur Sinha, Ramnish Sinha, and Vishnu Priye. Beyond classical approaches: redefining the landscape of high-accurate movie recommendation using QNN. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06746-x.

Ge:2025:DMO

[348] Fangzhen Ge, Xuan Zhao, Debao Chen, Longfeng Shen, and Huaiyu Liu. A dynamic multi-objective optimization algorithm based on probabilitydriven prediction and correlationguided individual transfer. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06832-0.

Martin-Salinas:2025:EAV

[349] Ignacio Martin-Salinas, Jose M. Badia, Oscar Valls, German Leon, Rocio del Amor, Jose A. Belloch, Adrian Amor-Martin, and Valery Naranjo. Evaluating and accelerating vision transformers on GPU-based embedded edge AI systems. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06807-1.

Zhang:2025:VTB

[350] Enzhi Zhang, Rui Zhong, Xingbang Du, Mohamed Wahib, and Masaharu Munetomo. Vision transformerbased meta loss landscape exploration with actor-critic method. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06867-3.

Kong:2025:RSS

[351] Dekun Kong and Wengaung Yang. Research on site selection of emergency material reserve based on set pair analysis and TOPSIS integration method: a case study of Hebei Province, China. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06813-3. See cor-

rection [447].

Sakano:2025:NNP

[352] Koh Sakano, Kairi Furui, and Masahito Ohue. NPGPT: natural productlike compound generation with GPTbased chemical language models. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06860-w.

Gan:2025:IDM

[353] Lian Gan, YuHong Du, Shuai Wang, WeiJia Ren, ZiQi Rong, and XinLong Li. Intelligent decision making algorithm for path planning based on reference linguistic fuzzy set. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06842-y.

Lopez-Oliva:2025:EQC

[354] Vicente Lopez-Oliva, Jose M. Badia, and Maribel Castillo. Efficient quantum circuit contraction using tensor decision diagrams. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06836-w.

Espinosa:2025:SEG

[355] Elena Espinosa, Ricardo Quislant, Rafael Larrosa, and Oscar Plata. SeqMatcher: efficient genome sequence matching with AVX-512 extensions. The Journal of Supercomputing, 81 (1):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06789-0.

Ghadiri:2025:IBL

[356] Mohammad Javad Ghadiri and Mehri Bagherian. Improved binary linear programming models for finding maximum edge bi-clique in bipartite graphs. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06733-2.

Liang:2025:RMP

[357] Xinyu Liang, Guannan Si, Jianxin Li, Pengxin Tian, Zhaoliang An, and Fengyu Zhou. Relational message passing with mutual information maximization for inductive link prediction. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06749-8.

Wang:2025:MPG

[358] Kexin Wang, Dingrui Xue, Yingdong Gou, Wanlong Qi, Bo Li, Jiancheng Liu, Yinglong Feng, and Yuqing Lin. Meta-path-guided causal inference for hierarchical feature alignment and policy optimization in enhancing resilience of UWSoS. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06848-6.

Dehkordi:2025:EIC

[359] Afsaneh Banitalebi Dehkordi. EDBLSD-IIoT: a comprehensive hybrid architecture for enhanced data security, reduced latency, and optimized energy in industrial IoT networks. The Journal of Supercomputing, 81 (2):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06872-6.

Jha:2025:AOR

[360] Mukesh Kumar Jha and Mohit Kumar. An autonomic offloading and resource allocation technique for IoT applications in edge computing. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06491-1.

Chung:2025:FGG

[361] Wu-Chun Chung, Jyun-Sen Tong, and Zhi-Hao Chen. A fine-grained GPU sharing and job scheduling for deep learning jobs on the cloud. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06849-5.

Abideen:2025:AIS

[362] Syed Zain Ul Abideen, Abdul Wahid, Mian Muhammad Kamal, Nouman Imtiaz, Nabila Sehito, Yousef Ibrahim Daradkeh, Mahmoud Ahmad Al-Khasawneh, Abdullah Alwabli, and Inam Ullah. Advancements in IoT system security: a reconfigurable intelligent surfaces and backscatter communication approach. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06819-x.

Guo:2025:ROP

[363] Yuepeng Guo, ZhenPing Lan, Yanguo Sun, Yuheng Sun, Xinxin Li, Yuru Wang, and Bo Li. Research on occlusion pedestrian reidentification based on ViT model. The Journal of Supercomputing, 81 (2):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06831-1.

Diaz-Cano:2025:AMW

[364] Roberto Díaz-Cano, Francesc Folch, Enrique S. Quintana-Ortí, and Pedro Alonso-Jordá. Acceleration of the MVS workflow using graphics processors. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06835-x.

Xiao:2025:LGL

[365] Senyue Xiao, Jianhua Liu, Zeming Pan, Shaoze Wang, Yang Yang, Zilong Song, and Anni Fan. LiteYOLO-GHG: a lightweight YOLOv8-based algorithm for transformer bushing fault detection. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06852-w.

Mehtarizadeh:2025:SPP

[366] Homa Mehtarizadeh, Najme Mansouri, Behnam Mohammad Hasani Zade, and Mohammad Mehdi Hosseini. Stock price prediction with SCA–LSTM network and statistical model ARIMA– GARCH. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06775-6.

Wu:2025:SGO

[367] Di Wu, Le Zhang, and Yao Chen. Syntactic-guided optimization of imagetext matching for intra-modal modeling. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06840-0.

Wang:2025:CCCb

[368] Zhen Wang, Dong Zhao, Ali Asghar Heidari, Huiling Chen, and Guoxi Liang. CGWRIME: collaboration and competition-boosted RIME optimizer for engineering optimization problems. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06817-z.

Lei:2025:LPP

[369] Lei Zhang, Yongbo Bai, Shiyi Lin, Shuaishuai Lian, Yijia Geng, and Zhili Liu. Location privacy protection scheme of user collaborative probabilistic indistinguishability based on Hyperledger Fabric. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06854-8.

Ewees:2025:OFS

[370] Ahmed A. Ewees, Mohammed M. Alshahrani, Abdullah M. Alharthi, and Marwa A. Gaheen. Optimizing feature selection and remote sensing classification with an enhanced machine learning method. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06790-7.

Sravan:2025:SLS

[371] S. S. Sravan, Susmita Mandal, and P. J. A. Alphonse. SDSMS-LoRa: secure dynamic session key management scheme for LoRaWAN v1.1. The Journal of Supercomputing, 81 (2):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06802-6.

Jiang:2025:FMP

[372] Rong Jiang, Yulin Li, Xuetao Pu, Xueke Wang, Wenyu Niu, and Zhiming Song. A fair multi-party contract signing scheme based on off-chain protocols and on-chain smart contracts. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06844-w.

Wang:2025:NAW

[373] Chenlu Wang, Xiaoyi Fu, Ziyi Yang, and Shenglin Li. NeuralWiGait: an accurate WiFi-based gait recognition system using hybrid deep learning framework. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06878-0.

Xu:2025:MCH

[374] Shuangfei Xu, Zhanjun Huang, Wenhao Bi, and An Zhang. A Monte Carlo hyper-heuristic algorithm with low-level heuristics reward prediction for missile path planning. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06771-w.

Wang:2025:LLA

[375] Peipeng Wang, Xiuguo Zhang, and Zhiying Cao. LogSD: log anomaly detection via topic words awareness semantic augmentation and categoryguided Mixup data augmentation. The Journal of Supercomputing, 81 (2):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06850-y.

Wu:2025:DDA

[376] Heng Wu, Zijun Zheng, Laishui Lv, Changchun Zhang, Dalal Bardou, Shanzhou Niu, and Gaohang Yu. Dara: distribution-aware representation alignment for semi-supervised domain adaptation in image classification. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06886-0.

Qi:2025:IMO

[377] Mingjun Qi, Xiaochun Wu, Keke Li, and Fenghao Yang. IPAQ: a multiobjective global optimal and timeaware task scheduling algorithm for fog computing environments. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06853-9.

Jiang:2025:BMB

[378] Rong Jiang, Hejiao Zhang, Zhiming Song, Shenghu Tian, and Wenlu Lou. T-BFL model based on twodimensional trust and blockchainfederated learning for medical data sharing. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06873-5.

Darabkh:2025:NGR

[379] Khalid A. Darabkh, Mamoun F. Al-Mistarihi, and Mera Ismail Al-Maaitah. Next-generation routing for autonomous vehicle networks based on innovative clustering: integrating SDN and fog computing along with AODV upon failure. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06880-6.

Xing:2025:EEE

[380] Hongjia Xing, Feng Yang, Xu Qiao, Fanruo Li, and Xinxin Huang. Enhanced end-to-end regression algorithm for autonomous road damage detection. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06871-7.

Lin:2025:HMD

[381] Ching-Sheng Lin. A hybrid model for the detection of multi-agent written news articles based on linguistic features and BERT. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06882-4.

Chen:2025:KNI

[382] Bo Chen, Yulin Zhang, Yunming Wang, Rui Tong, Yufeng Chen, Lingdong Sun, and Wenxue Xie.

Key node identification method for two-layer fusion networks based on improved structural holes. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06843-x.

Yu:2025:CAD

[383] Dongxian Yu, Xiaoyu Zhou, Ali Noorian, and Mehdi Hazratifard. Correction to: An AI-driven social media recommender system leveraging smartphone and IoT data. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06857-5. See [272].

Zhang:2025:COY

[384] Yunjie Zhang, Guofeng Gao, Yadong Chen, and Zhenjian Yang. Correction: ODD–YOLOv8: an algorithm for small object detection in UAV imagery. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06829-9. See [202].

Geng:2025:QEM

[385] Renxuan Geng, Yuang Guo, Guotao Wang, Yuansong Liu, Bingze Lv, Hui Wang, and Songyi Yu. A qualitative evaluation method for the weight of moving objects within a sealed cavity based on time-frequency spectrogram features. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06717-2.

Yang:2025:ETT

[386] Zhimi Yang and Bo Shen. Estimating textual treatment effect via causal disentangled representation learning. The Journal of Supercomputing, 81 (2):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06887-z.

Xia:2025:MMS

[387] Pusen Xia, Shengwei Tian, Long Yu, Xin Fan, Zhezhe Zhu, Hualong Dong, Na Qu, Tong Liu, and Xiao Yuan. Mdcsnet: multi-scale dynamic spatial information fusion with criticality sampling for point cloud classification. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06838-8.

Simen:2025:EEQ

[388] Anton Simen, Rodrigo Bloot, Otto M. Pires, and Erick G. Sperandio Nascimento. Evolutionary-enhanced quantum supervised learning model. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06875-3.

Sheng:2025:CTD

[389] Jinfang Sheng, Yifan Zhang, and Bin Wang. Continuous-time dynamic graph learning based on spatio-temporal random walks. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06881-5.

Li:2025:DSC

[390] Meng Li, Bo Yang, Tao Xue, and Shaowei Han. Deep subspace clustering using dual self-expressiveness and convolutional fusion. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06885-1.

Sultana:2025:ROQ

[391] Afrin Sultana and Edgard Muñoz-Coreas. Resource optimized quantum squaring circuit. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06876-
2.

Idrees:2025:SAL

[392] Sara Kadhum Idrees, Joseph Azar, Raphaël Couturier, Ali Kadhum Idrees, and Franck Gechter. SZ4IoT: an adaptive lightweight lossy compression algorithm for diverse IoT devices and data types. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06667-9.

Zhang:2025:NSD

[393] Yunsong Zhang, Yuejuan Han, Jianfeng Jiang, and Lantao You. Node-to-set disjoint paths problem in divideand-swap cube. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06895-
z.

Khala:2025:EML

[394] Mohamed Khala, Naima El Yanboiy, Ismail Elabbassi, Omar Eloutassi, Mohammed Halimi, Youssef El Hassouani, and Choukri Messaoudi. Enhancing machine learning model for early warning in PV plants: air temperature prediction informed by power temperature coefficient. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06909-w.

Osmani:2025:FSD

[395] Nooshin Osmani, Erfan Esmaeeli, and Sorayya Rezayi. Fusion strategies for deep convolutional neural network representations in histopathological image classification. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06663-z.

Wang:2025:CCCa

[396] Bo Wang, Rong Jiang, Xuetao Pu, and Hejiao Zhang. An on-chain and off-chain collaborative data sharing and access control model for electronic medical records. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06884-2.

Li:2025:DLP

[397] Benchao Li, Ting Wang, and Ruisheng Ran. Discriminant locality preserving projection on Grassmann Manifold for image-set classification. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06904-1.

Zong:2025:PHT

[398] Jixiang Zong, Jiulong Wang, Guirun Li, Ruopu Wu, and Di Zhao. Polaris 23: a high throughput neuromorphic processing element by RISC-V customized instruction extension for spiking neural network (RV-SNN 2.0) and SIMD-style implementation of LIF model with backpropagation STDP. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06826-y.

Rafieinejad:2025:IAC

[399] Mohammadreza Rafieinejad, Mohammadreza Binesh Marvasti, Seyyed Amir Asghari, and Kimiya Shahbakhti. Improve accuracy in CNNs while using approximate computing methods. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06901-4.

Lu:2025:DDH

[400] Jinkang Lu, Meng Lv, Peixuan Li, Zhu Yuan, and Ping Xie. Dhcache: a dual-hash cache for optimizing the read performance in key–value store. The Journal of Supercomputing, 81 (2):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06828-w.

Tourani:2025:STR

[401] Azadeh Navaei Tourani, Hamid Haj Seyyed Javadi, Hamidreza Navidi, and Arash Sharifi. A study on trust-rating mechanism for WSN node sensors using evolutionary game theory. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06824-0.

Kumar:2025:FRB

[402] Shashi Shekhar Kumar, Ritesh Chandra, Anurag Harsh, and Sonali Agarwal. Fuzzy rule-based intelligent cardiovascular disease prediction using complex event processing. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06911-2.

Chahardoli:2025:ECP

[403] Meysam Chahardoli, Nafiseh Osati Eraghi, and Sara Nazari. An energy consumption prediction approach in smart cities by CNN–LSTM network improved with game theory and Namib Beetle Optimization (NBO) algorithm. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06811-5.

Alasiry:2025:CSV

[404] Areej Alasiry and Mohammed Qayyum. Correction: A smart vista-lite system for anomaly detection and motion prediction for video surveillance in vibrant urban settings. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06877-1. See [297].

Tang:2025:CCE

[405] Yiou Tang, Yan Ma, Chunling Xiao, Min Wu, and Guoyuan Zeng. Correction: Classification of EEG eventrelated potentials based on channel attention mechanism. The Journal of Supercomputing, 81(2):

??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06830-2. See [126].

Jin:2025:RGN

[406] Yan Jin, Haoyu Shi, and Huaiye Meng. Robust graph neural networks based on feature fusion. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-06917-4.

Zhang:2025:MLT

[407] Qian Zhang, Chenghao Ji, Mingwen Shao, and Hong Liang. MEKF: longtailed visual recognition via multiple experts with knowledge fusion. The Journal of Supercomputing, 81 (2):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-06920-9.

Zhang:2025:EVT

[408] Qian Zhang, Zuosui Yang, Mingwen Shao, and Hong Liang. An efficient video transformer network with token discard and keyframe enhancement for action recognition. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-06927-2.

Liang:2025:MLN

[409] Hong Liang, Xian Li, Mingwen Shao, and Qian Zhang. Multi-level navigation network: advancing finegrained visual classification. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06933-4.

Ferreira:2025:UDL

[410] Fernando Rodrigues Trindade Ferreira and Loena Marins do Couto. Using deep learning on microscopic images for white blood cell detection and segmentation to assist in leukemia diagnosis. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06903-2.

Wang:2025:DSM

[411] Yan Wang, Zheng Gong, Dayu Jia, Aiping Tan, and Minchao Liu. Dynamic sharding model and performance optimization method for consortium blockchain. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06870-8.

Huang:2025:MTS

[412] Nana Huang, Hongwei Ding, Ruimin Hu, Pengfei Jiao, Zhidong Zhao, Bin Yang, and Qi Zheng. Multitime-scale with clockwork recurrent neural network modeling for sequential recommendation. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06925-4.

Liu:2025:TTD

[413] Tao Liu, Chenyoukang Lin, Yunteng Hu, Ruyi Cao, and Wendong Zhang. Traffic target detection based on context enhancement and feature purification. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06944-1.

Sun:2025:LLA

[414] Yuheng Sun, Zhenping Lan, Yanguo Sun, Yuepeng Guo, Xinxin Li, Yuru Wang, and Bo Li. Ldstd: lowaltitude drone aerial small target detector. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06950-3.

Li:2025:SDD

[415] Biao Li, Bing Wang, Xiong Hu, Jianhui Zhai, and Changping Ji. A small defect detection technique for industrial product surfaces based on the EA–YOLO model. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06929-0.

Chen:2025:MMS

[416] Wuqi Chen, Junjie Ye, Chunna Zhao, and Yaqun Huang. MFFCNN: multiscale fractional Fourier transform convolutional neural network for multivariate time series forecasting. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06888-y.

Xue:2025:EOW

[417] Wenjin Xue, Guowei Xu, Nan Yang, and Jian Liu. Enhancing open-world object detection with AIGC-generated datasets and elastic weight consolidation. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06910-3.

Cardama:2025:RIR

[418] F. Javier Cardama, Jorge Vázquez-Pérez, César Piñeiro, Juan C. Pichel, Tomás F. Pena, and Andrés Gómez. Review of intermediate representations for quantum computing. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06892-2.

Khah:2025:CHM

[419] Yashar Pourardebil Khah, Mirsaeid Hosseini Shirvani, and Homayun Motameni. Correction: A hybrid machine learning approach for feature selection in designing intrusion detection systems (IDS) model for distributed computing networks. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06869-1. See [254].

Huang:2025:PPL

[420] Qiongyan Huang, Yuhan Xia, Yunfei Long, Hui Fang, Ruiwei Liang, Yin Guan, and Ge Xu. Prompt4LJP: prompt learning for legal judgment prediction. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06945-0.

Peng:2025:TPP

[421] Tianqi Peng, Bei Gong, and Pengxuan Sun. Toward privacy-preserving verifiable DSSE for attribute-based cloud computing system. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06912-1.

Eufracio:2025:RBF

[422] Alfredo Cisneros Eufracio, Roberth Saenz Perez Alvarado, Jimmy Aurelio Rosales Huamani, Uwe Rojas Villanueva, Jose Luis Castillo Sequera, and Jose Manuel Gomez Pulido. Rock block fall prediction prototype by structural control applied to slopes using quantum machine learning (QML).

The Journal of Supercomputing, 81 (2):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06913-0.

Alahmed:2025:ADL

[423] Hiba A. Alahmed and Ghaida A. Al-Suhail. AlzONet: a deep learning optimized framework for multiclass Alzheimer’s disease diagnosis using MRI brain imaging. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06924-5.

Mishra:2025:CWG

[424] Sanjukta Mishra, Samarjit Kar, and Parag Kumar Guhathakurta. Cloud-WAVECAP: Ground-based cloud types detection with an efficient wavelet-capsule approach. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06941-4.

Li:2025:LSQ

[425] Ruoxia Li, Linli Si, and Jinde Cao. Lagrange stability of quaternion-valued memristive neural networks on time scales: linear optimization method. The Journal of Supercomputing, 81 (2):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06914-z.

Darabkh:2025:LFC

[426] Khalid A. Darabkh, Mamoun F. Al-Mistarihi, and Bayan Abdallah Odat. Leveraging fog computing and software-defined networking for a novel velocity-aware routing protocol with election and handover thresholds in VANETs. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06883-3.

Lopez:2025:IEP

[427] Marta López, Esteban Stafford, and Jose Luis Bosque. Intelligent energy pairing scheduler (InEPS) for heterogeneous HPC clusters. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06907-y.

Campos:2025:EDF

[428] Cristian Campos, Rafael Asenjo, and Angeles Navarro. Exploring data flow design and vectorization with oneAPI for streaming applications on CPU+GPU. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06891-3.

Liang:2025:MMS

[429] Hong Liang, Cuiping Wang, Mingwen Shao, and Qian Zhang. MAQT: multiscale attention and query-optimized transformer for end-to-end pose estimation. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06923-6.

Liu:2025:CPS

[430] Jianhang Liu, Chunxing Xia, Xuerong Cui, and Haibo Wu. CPCS: a perception sharing scheme of vehicleroad cooperation based on cybertwin. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06939-y.

Li:2025:DDM

[431] Pei Li, Tong Bai, Xiaoying Pan, and Chengyu Zuo. Defocus deblur method of multi-scale depth-of-field cross-stage fusion image based on defocus map forecast. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06934-3.

Safi-Esfahani:2025:LWH

[432] Faramarz Safi-Esfahani, Leili Mohammadhoseini, Habib Larian, and Seyedali Mirjalili. LEVYEFO-WTMTOA: the hybrid of the multitracker optimization algorithm and the electromagnetic field optimization. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06856-6.

Yang:2025:MRS

[433] Xiaonan Yang, Zuoxi Zhao, Kai Yuan, Can Xiao, and YangFan Luo. Mobile recognition system for multinational currencies based on CA–DSC-RepVGG algorithm. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-06936-1.

Zhang:2025:IJA

[434] Qiwen Zhang and Tian Zhen. Improved Jaya algorithm for energyefficient distributed heterogeneous permutation flow shop scheduling. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-06938-z.

Zangeneh:2025:NSC

[435] Iman Zangeneh, Amir Massoud Bidgoli, and Ardeshir Dolati. Novel solution of the controller and backup controller placement problem for improving reliability in IoT-based data monitoring systems. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06846-8.

Nascimento:2025:AAS

[436] Erick Nascimento, Luan Lins, Eduardo Tavares, Paulo Pereira, Jamilson Dantas, Sokol Kosta, and Paulo Maciel. Availability assessment of SDN–ICN service for multi-access edge computing. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06879-z.

Prasad:2025:BDG

[437] Chandrabhushan Prasad, Sri Khetwat Saritha, and Sweta Jain. Bilinear diffusion graph convolutional network model for social recommendation. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06930-7.

Zaimi:2025:EMM

[438] Rania Zaimi, Khouloud Safi Eljil, Mohamed Hafidi, Mahnane Lamia, and Farid Nait-Abdesselam. An enhanced mechanism for malicious URL detection using deep learning and DistilBERT-based feature extraction. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06908-x.

Wu:2025:GHF

[439] Chengmao Wu and Siyu Zhou. Generalized harmonic fuzzy partition C-means clustering. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06723-4.

Li:2025:MOE

[440] Wei Li, Wenhao Tang, and Lei Wang. Many-objective evolutionary algorithm based on dynamic mating and strengthened fitness selection mechanism. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06821-3.

Dong:2025:APB

[441] Jian Dong, Wei Bao, Xiaoqi Cao, Yang Xu, Yuze Yang, Binbin Li, Qi Zhang, and Heng Ye. AISBench: an performance benchmark for AI server systems. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06778-3.

Vazquez-Perez:2025:IIC

[442] Jorge Vázquez-Pérez, F. Javier Cardama, César Piñeiro, Juan C. Pichel, Tomás F. Pena, and Andrés Gómez. Inqasm: InQuIR compiler to NetQASM. The Journal of Supercomputing, 81 (2):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06955-y.

Aymaz:2025:UPO

[443] Samet Aymaz. Unlocking the power of optimized data balancing ratios: a new frontier in tackling imbalanced datasets. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06919-2.

Xiao:2025:MDD

[444] Junbi Xiao, Ruifeng Sun, and Jianhang Liu. MLDDoS: a distributed denial of service attack detection method using multi-level sketch. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06942-3.

Tondreau:2025:SDH

[445] Pierre-Simon Callist Yannick Tondreau, Juan C. Perez, Juan P. Díaz, Vicente Blanco Pérez, and Jonatan Felipe. Singularity to deploy HPC applications: a study case with WRF. The Journal of Supercomputing, 81 (2):??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06893-1.

Alkhudhayr:2025:MCR

[446] Hanadi Alkhudhayr and Hanin Ardah. Mitigating cyberphysical risks in IoT-enabled smart transport infrastructure. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06948-x.

Kong:2025:CRS

[447] Dekun Kong and Wenguang Yang. Correction: Research on site selection of emergency material reserve based on set pair analysis and TOP-SIS integration method: a case study of Hebei Province, China. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06971-y. See [351].

Huang:2025:LHD

[448] Hong Huang, Wengang Luo, Yunfei Wang, Yinghang Zhou, and Weitao Huang. LogCTBL: a hybrid deep learning model for log-based anomaly detection. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06926-3.

Fu:2025:MMS

[449] Youjia Fu, Zihao Xu, Junsong Fu, Huixia Xue, Shuqiu Tan, Lei Li, and Shaoxun Qing. Monomm: a Multiscale Mamba-Enhanced network for real-time monocular 3D object detection. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06946-z.

Pastor:2025:CDB

[450] Alfred M. Pastor, Jose M. Badia, and Maribel Castillo. A community detection-based parallel algorithm for quantum circuit simulation using tensor networks. The Journal of Supercomputing, 81(3): ??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06918-3.

Suarez:2025:KCB

[451] Daniel Suárez, Francisco Almeida, Vicente Blanco, and Pedro Toledo. KubePipe: a container-based highlevel parallelization tool for scalable machine learning pipelines. The Journal of Supercomputing, 81(3): ??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06956-x.

Raul:2025:CGC

[452] Nozal Raúl and Jose Luis Bosque. CPU-GPU co-execution through the exploitation of hybrid technologies via SYCL. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06963-y.

Yanamala:2025:FAH

[453] Rama Muni Reddy Yanamala and Muralidhar Pullakandam. FPGAaccelerated hybrid CNN–LSTM system for efficient EEG-based drowsiness recognition. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06947-
                                        y.

Zhang:2025:FIE

[454] Xuncai Zhang, Mengrui Liu, and Ying Niu. Facial image encryption scheme based on improved 4-D hyperchaotic system. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-06943-2.

Roh:2025:HQC

[455] Emily Jimin Roh, Joo Yong Shim, Joongheon Kim, and Soohyun Park. Hybrid quantum-classical 3D object detection using multi-channel quantum convolutional neural network. The Journal of Supercomputing, 81 (3):??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-06968-7.

Zhou:2025:SRF

[456] Sizhong Zhou, Zhiren Sun, and Yuli Zhang. Spectral radius and kfactor-critical graphs. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06902-3.

Zhao:2025:PSM

[457] Shengsheng Zhao, Yantao Yu, Tiancong Huang, Guojin Liu, and Yucheng Wu. Pricing strategies in mobile crowdsensing: an enhanced MAPPO approach using a behavior network. The Journal of Supercomputing, 81 (3):??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06957-w.

Liu:2025:MSE

[458] Tianbao Liu and Zhe Feng. Multistrategy enhanced dandelion optimizer based on elliptic approximation strategy and adaptive fitnessdistance-similarity balance for solar photovoltaic parameter estimation. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06899-9.

Xiao:2025:JCO

[459] Qimu Xiao and Mingyu Xiao. Joint computation offloading and resource allocation in multi-cell MEC networks. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06921-8.

Xiong:2025:ORT

[460] Zhi Xiong, Linhui Tan, Jianlong Xu, and Lingru Cai. Online real-time energy consumption optimization with resistance to server switch jitter for server clusters. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06827-x.

Zhang:2025:NSS

[461] Shengcai Zhang, Zhiying Fu, Dezhi An, and Huiju Yi. Network security situation assessment based on BKA and cross dual-channel. The Journal of Supercomputing, 81(3): ??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06932-5.

Nguyen:2025:CBA

[462] Da-Thao Nguyen, Thanh-Phuong Nguyen, and Ming-Yuan Cho. Cloudbased AIoT intelligent infrastructure for firefighting pump fault diagnosis-based hybrid CNN–GRU deep learning technique. The Journal of Supercomputing, 81(3): ??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06965-w.

Ismail:2025:OSB

[463] Nabil A. Ismail, Shaimaa Abu Khadra, Gamal M. Attiya, and Salah Eldin S. E. Abdulrahman. Optimizing SIKE for blockchain-based IoT ecosystems with resource constraints. The Journal of Supercomputing, 81 (3):??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06906-z.

Atencia:2025:PQO

[464] Javier Ruiz Atencia, Otoniel Mario López Granado, Manuel Pérez Malumbres, and Miguel Onofre Martínez-Rach. Perceptual QP optimization for VVC with dual hybrid neural networks. The Journal of Supercomputing, 81 (3):??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06954-z.

Nunez-Chongo:2025:CDD

[465] Osiris Núñez-Chongo, Hernán Asorey, Antonio Juan Rubio-Montero, Mauricio Suárez-Durán, Rafael Mayo-García, and Manuel Carretero. Convergent data-driven workflows for open radiation calculations: an exportable methodology to any field. The Journal of Supercomputing, 81(3): ??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06894-0.

Wen:2025:VOD

[466] Jia Wen, Qi Zhang, and Guanghao Zhang. VPSNet: 3D object detection with voxel purification and fully sparse convolutional networks. The Journal of Supercomputing, 81 (3):??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06890-4.

Yang:2025:SVI

[467] Qimeng Yang, Hao Meng, Yuanbo Yan, Shisong Guo, and Qixing Wei. SFVE: visual information enhancement metaphor detection with multimodal splitting fusion. The Journal of Supercomputing, 81(3): ??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-06958-9.

Kim:2025:PEI

[468] Suhri Kim, Youngdo Lee, and Kisoon Yoon. Performance evaluation of isogeny-based digital signature algorithms: introducing FIBSfast isogeny-based digital signature*. The Journal of Supercomputing, 81 (3):??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-06970-z.

Roselin:2025:DDL

[469] J. Roselin and Israelin J. Insulata. Decentralized dynamic load balancing for virtual machines in cloud computing: a blockchain-enabled system with state channel optimization. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-06922-7.

Zheng:2025:CGG

[470] Zhiwen Zheng, Xiaoyun Chen, and Musheng Huang. Cluster-guided graph attention auto-encoder. The Journal of Supercomputing, 81(3): ??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06953-0.

Zhou:2025:PLA

[471] Zhuguo Zhou, Yujun Lu, and Liye Lv. Pallet localization algorithm based on improved human pose estimation with transfer learning. The Journal of Supercomputing, 81(3): ??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06973-w.

Vicente-Jaen:2025:VRT

[472] Arturo Vicente-Jaén, Juan Mompeán, Juan L. Aragón, and Pablo Artal. VolRec: 4D real-time volumetric reconstruction of OCT data. The Journal of Supercomputing, 81 (3):??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06969-6.

Fang:2025:DRC

[473] Juan Fang, Haoyu Cheng, Yuening Wang, and Ran Zhai. DRCD: a regional-contention-driven arbitration policy for CPU–GPU heterogeneous systems. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07001-7.

Kirubavathi:2025:DMT

[474] G. Kirubavathi, I. R. Sumathi, J. Mahalakshmi, and Durgesh Srivastava. Detection and mitigation of TCPbased DDoS attacks in cloud environments using a self-attention and intersample attention transformer model. The Journal of Supercomputing, 81 (3):??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-06940-5.

Tolun:2025:CBM

[475] Ömer Can Tolun, Kasim Zor, and Onder Tutsoy. A comprehensive benchmark of machine learning-based algorithms for medium-term electric vehicle charging demand prediction. The Journal of Supercomputing, 81 (3):??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-06975-8.

Wang:2025:ERD

[476] Youwei Wang, Lizhou Feng, and Yan Zhang. Early rumor detection method based on stage sampling and triple-relationship graph. The Journal of Supercomputing, 81 (3):??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-06959-8.

Cui:2025:PES

[477] Mengqi Cui, Jinchen Xu, Yuchang Zhou, Hongru Yang, Liguang Ji, and Bei Zhou. PESA: error sensitivity analysis tool for floatingpoint computational programs. The Journal of Supercomputing, 81(3): ??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06962-z.

Shao:2025:RFF

[478] Changshun Shao, Zhenglin Yu, Hongchang Ding, Guohua Cao, Jingsong Duan, and Bin Zhou. A random flight-follow leader and reinforcement learning approach for flexible job shop scheduling problem. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06935-2.

Hasani:2025:MQP

[479] Mehri Hasani, Masoud Ghods, Sourav Mondal, Muhammad Kamran Siddiqui, and Imran Zulfiqar Cheema. Modeling QSPR for pyelonephritis drugs: a topological indices approach using MAT-LAB. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06967-8.

Cao:2025:CFM

[480] Zhipeng Cao, Nianmin Yao, Linqi Meng, Jingyi Fang, and Jian Zhao. Coarse-to-fine medical image registration with landmarks and deformable networks. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07000-8.

Zhang:2025:SSA

[481] Yuxin Zhang, Jiazheng Wen, Ran Wu, Huanyu Liu, and Junbao Li. Siamsdt: a self-adaptive dynamic template siamese network for airborne visual tracking of MAVs on heterogeneous FPGA–SoC. The Journal of Supercomputing, 81(3): ??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-06928-1.

Mashayekhi:2025:DEE

[482] Negin Mashayekhi, Ghassem Jaberipur, Mohammad Reza Reshadinezhad, and Shekoofeh Moghimi. Design of energyefficient and high-speed hybrid decimal adder. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06897-x.

He:2025:IDM

[483] Fuchun He, Chunming Fu, Youwei He, Shaoyong Huo, Jiachang Tang, and Xiangyun Long. Improved dwarf mongoose optimization algorithm based on hybrid strategy for global optimization and engineering problems. The Journal of Supercomputing, 81 (3):??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-06931-6.

Nieves-Perez:2025:CEE

[484] Isidoro Nieves-Pérez, Alfonso Muñoz, Francisco Almeida, and Vicente Blanco. Correction: Energy efficiency and performance analysis of a legacy atomic scale materials modeling simulator (VASP). The Journal of Supercomputing, 81(3): ??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06976-7.

Narwaria:2025:RER

[485] Abhishek Narwaria, Varsha Kumari, and Arka Prokash Mazumdar. RL-EAR: reinforcement learning-based energy-aware routing for softwaredefined wireless sensor network. The Journal of Supercomputing, 81(3): ??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06998-1.

Lan:2025:EIA

[486] Yu Lan, Hui Huang, Zhenjie Huang, Qunshan Chen, and Shuaike Wu. ERBFT: improved asynchronous BFT with erasure code and verifiable random function. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06995-
4.

Qiu:2025:MRC

[487] Hong Qiu, Wentao Yu, Gan Zhang, Xuan Xia, and Kun Yao. Multirobot collaborative 3D path planning based on game theory and particle swarm optimization hybrid method. The Journal of Supercomputing, 81 (3):??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06960-1.

Zhang:2025:TBF

[488] Dabin Zhang, Zehui Yu, Zhimei Zeng, Boting Zhang, Ruibin Lin, and Huanling Hu. A text-based framework for carbon price forecasting via multivariate temporal graph neural network. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06974-9.

Tang:2025:CFS

[489] Yu Tang, Qi Dai, Ye Du, Tian shuai Zheng, and Mei hong Li. Capsule feature selector for software defect prediction. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06949-w.

Shao:2025:SSS

[490] Yunxue Shao, Zhiyang Wang, and Lingfeng Wang. SSIM: self-supervised learning method based on spatially selected shifts and irregular image masking. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07013-3.

Meng:2025:GAM

[491] Yi Meng, Nurbol Luktarhan, Xiaotong Yang, and Guodong Zhao. GBADroid: an Android malware detection method based on multi-view feature fusion. The Journal of Supercomputing, 81 (3):??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06977-6.

Zhang:2025:UAP

[492] Heng Zhang, Yang Peng, and Yan li Liu. UAV aerial photography target detection based on improved YOLOv9. The Journal of Supercomputing, 81 (3):??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06991-8.

Dang:2025:EPP

[493] Lanxue Dang, Zan Li, Shilong Li, Baojun Qiao, and Liming Zhou. Effective plug-and-play lightweight modules for YOLO series models. The Journal of Supercomputing, 81(3): ??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07026-y.

Ma:2025:TTF

[494] Siwei Ma, Ronghua Li, and Henan Hu. Train track fastener defect detection algorithm based on MGSF– YOLO. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07024-0.

Wan:2025:VFA

[495] Junhang Wan, Yanping Chen, and Cong Gao. The VAE–FastGA anomaly detection model based on subspace and weakly correlated ultrahigh-dimensional data. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06988-3.

Zhang:2025:EQG

[496] Yu Zhang, Lu Lu, Rong Zhao, Yijie Guo, and Zhanyu Yang. An efficient quantized GEMV implementation for large language models inference with matrix core. The Journal of Supercomputing, 81(3): ??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06993-6.

Chen:2025:OCA

[497] Zigang Chen, Zhenjiang Zhang, Tao Leng, Haihua Zhu, and Yuhong Liu. Outsourcing collaboration analysis of multiparty privacy data using the improved Yannakakis. The Journal of Supercomputing, 81(3): ??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06994-5.

Wu:2025:QRB

[498] Faguo Wu, Bo Zhou, Jiale Song, and Lijia Xie. Quantum-resistant blockchain and performance analysis. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07018-y.

Bao:2025:GCG

[499] Qingda Bao, Shengfa Miao, Yulin Tian, Xin Jin, Puming Wang, Qian Jiang, Shaowen Yao, Da Hu, and Ruoshu Wang. GDRNet: a channel grouping based time-slice dilated residual network for long-term time-series forecasting. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07011-5.

Ma:2025:GAT

[500] Minyi Ma, Hongfang Gong, and Yingjing Ding. Graph affine transformer with a symmetric adaptation strategy for text classification. The Journal of Supercomputing, 81 (3):??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06992-7.

Luo:2025:AFD

[501] Mao Luo, Xianhong Liu, Xinyun Wu, Caiquan Xiong, and Yuanzhi Ke. An adaptive focal distance tabu search approach for the minimum 2-connected dominating set problem. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07014-2.

Basarslan:2025:MCB

[502] Muhammet Sinan Basarslan. M-C&M-BL: a novel classification model for brain tumor classification: multi-CNN and multi-BiLSTM. The Journal of Supercomputing, 81(3): ??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06964-x.

Esfandiari:2025:IIN

[503] Shima Esfandiari and Seyed Mostafa Fakhrahmad. Identifying influential nodes in complex networks by adjusted feature contributions and neighborhood impact. The Journal of Supercomputing, 81(3): ??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06645-1.

Zhong:2025:PPR

[504] Jie Zhong, Juan Chen, Zongling Wu, Peng Chen, Di Tian, and Xi Li. P2PPO: parallel residual network and prioritized experience replay enhanced PPO for task offloading and resource allocation in SatEC. The Journal of Supercomputing, 81(3): ??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06900-5.

Zhou:2025:AYS

[505] Wentao Zhou, Chengtao Cai, Sutthiphong Srigrarom, Pengfei Wang, Zijian Cui, and Chenming Li. ADH-YOLO: a small object detection based on improved YOLOv8 for airport scene images in hazy weather. The Journal of Supercomputing, 81(3): ??, February 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06999-0.

AlDawi:2025:ABA

[506] Abdelaziz Al Dawi, Necmi Serkan Tezel, Javad Rahebi, and Ayhan Akbas. An approach to botnet attacks in the fog computing layer and Apache Spark for smart cities. The Journal of Supercomputing, 81 (4):??, March 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06915-y.

Yu:2025:MSM

[507] Changhong Yu and Mingxuan Zhang. M^{2}SSCENet: a multi-branch multiscale network with spatial-spectral cross-enhancement for hyperspectral and LiDAR data classification. The Journal of Supercomputing, 81(4): ??, March 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07096-y.

Zhu:2025:HHS

[508] Heming Zhu, Hao Li, and Gehao Lu. HSVDetector: a heterogeneous semantic graph-based method for smart contract vulnerability detection. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-06951-2.

Jiang:2025:EEC

[509] Chenyang Jiang, Zhendong Li, Jun Yang, Yiqiang Wu, and Shuai Li. Effective and efficient conditional contrast for data-free knowledge distillation with low memory. The Journal of Supercomputing, 81(4): ??, March 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07115-y.

Niu:2025:MFF

[510] Yakun Niu, Xiangru Chen, and Hongjian Yin. Median filtering forensics using spatial and frequency domain residuals. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07099-9.

Tong:2025:SYS

[511] Haoyang Tong, Dongyang Gao, Zhixu Wang, Longlong Feng, Yue Li, and Xuewei Bai. SGB-YOLOv5: straw granulator blockage monitoring system. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07079-z.

Xia:2025:STS

[512] Wenyuan Xia, Qing Zhou, Dayu Wu, Siyuan Wang, and Mengshuang Zhou. A scalable two-stage model for real-time wetland bird recognition. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07061-9.

Li:2025:MNM

[513] Peng Li, Zhanao Huang, Yan Li, Yi Kou, and Xiaojie Li. Mfe-net: a multiscale feature enhanced network for mesoscale convective systems identification. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07065-5.

Qin:2025:LSM

[514] Chuandong Qin, Yiqing Zhang, and Yu Cao. Large-scale machine learning with synchronous parallel adaptive stochastic variance reduction gradient descent for high-dimensional blindness detection on Spark. The Journal of Supercomputing, 81(4): ??, March 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07046-8.

Ye:2025:GIG

[515] Pengdong Ye, Yanhua Liang, Yutao Bie, Guihe Qin, Jiaru Song, Yingqing Wang, and Wanning Liu. GDT-IDS: graph-based decision tree intrusion detection system for controller area network. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07116-x.

Gheibi-Fetrat:2025:SSS

[516] Atiyeh Gheibi-Fetrat, Fatemeh Serajeh Hassani, Masoud Mohammadi-Lak, Amir Mirzaei, Negar Akbarzadeh, Mahmoud Reza Kheyrati-Fard, Mohammad Hosseini, Ahmad Javadi Nezhad, Arash Tavakkol, Jeong A. Lee, and Hamid Sarbazi-Azad. A survey of SSD simulators and emulators. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07078-0.

Li:2025:HPA

[517] Runhua Li, Qinglin Wang, and Jie Liu. A heterogeneous parallel algorithm for the Cartesian discrete ordinates for multizone heterogeneous system. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07087-z.

Wang:2025:AAP

[518] Ziheng Wang, Zixuan Shao, Baowei Wang, and Xu Cheng. APPBoost: an adaptive parameter pair boosting algorithm for enhanced robustness against noise and imbalance. The Journal of Supercomputing, 81 (4):??, March 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07053-9.

Li:2025:DSS

[519] Shouliang Li, Rudong Min, Jilong Zhang, Jiale Han, Yulin Shen, Zhen Yang, and Yi Yang. A digital signature scheme based on general Chebyshev polynomial. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07074-4.

Sun:2025:KEN

[520] Pengfei Sun, Zhiping Wang, Liyan Jia, and Xiaoxi Wang. KSIPF: an effective noise filtering oversampling method based on k-means and iterative-partitioning filter. The Journal of Supercomputing, 81(4): ??, March 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07081-5.

Rezaei:2025:MNH

[521] Azita Rezaei, Ali Broumandnia, and Seyed Javad Mirabedini. MLS: a novel hybrid security framework utilizing the Wiedemann algorithm and chaotic mapping for MQTT. The Journal of Supercomputing, 81(4):

??, March 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06858-4.

Lawrence:2025:DSH

[522] Linju Lawrence and R. Shreelekshmi. Double salted HMAC signature with blockchain for faster and secure video integrity verification. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-06996-3.

Zhang:2025:DMO

[523] Zhibo Zhang, Huiqiang Wang, Rongqiang Li, Hongwu Lv, and Dongmiao He. DQN-MSRA: an online SFC deployment method based on multistep reinforcement learning. The Journal of Supercomputing, 81(4): ??, March 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07042-y.

Turkmen:2025:SSC

[524] Muhammed Türkmen, Canan Çiftçi, and Gülnaz Boruzanli Ekinci. Structure and substructure connectivity of folded divide-and-swap cube. The Journal of Supercomputing, 81(4): ??, March 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07100-5.

delaRosa:2025:QSP

[525] Miguel Sánchez de la Rosa, Gabriel Gomez-Lopez, Francisco J. Andújar, Jesús Escudero-Sahuquillo, José L. Sánchez, Francisco J. Alfaro-Cortés, and Pierre-Axel Lagadec. Qualityof-service provision for BXIv3-based interconnection networks. The Journal of Supercomputing, 81(4): ??, March 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07069-1.

Genc:2025:IAC

[526] Hasan Genç, Canan Koç, Esra Yüzgeç Özdemir, and Fatih Özyurt. An innovative approach to classify meniscus tears by reducing vision transformers features with elasticnet approach. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07103-2.

Gao:2025:SCL

[527] Jianbang Gao, Yuxiao Guo, and Guowang Gao. Sim-ConvFormer: a lightweight fault diagnosis framework incorporating SimAM and external attention. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07075-3.

Huang:2025:CFS

[528] Weidi Huang, Zhiqiang Yao, Biao Jin, Zheyu Chen, and Yue Wang. Controllable face soft-biometric privacy enhancement based on attribute disentanglement. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07134-9.

Yan:2025:SCB

[529] Xiao Yan, Fan Xiumei, Kok-Lim Alvin Yau, Xie Zhixin, and Men Rui. Semantic communication based on bilevel routing attention in IoT environment. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07062-8.

Fu:2025:DAD

[530] Honghao Fu and Yidong Chen. Data augmentation and debiasing for signers in signer-independent sign language translation. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07119-8.

Yao:2025:DEG

[531] Ruiyang Yao. Dual-enhanced graph convolutional networks for aspectbased financial sentiment analysis. The Journal of Supercomputing, 81 (4):??, March 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06972-x.

Qin:2025:SID

[532] Qiurong Qin, Yueqin Li, Yajie Mi, Jinhui Shen, Kexin Wu, and Zhenzhao Wang. Sql injection detection algorithm based on Bi–LSTM and integrated feature selection. The Journal of Supercomputing, 81(4): ??, March 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07109-w.

Yang:2025:MSG

[533] Wenbiao Yang, Tingfeng Lai, and Yuhui Fang. Multi-Strategy Golden Jackal Optimization for engineering design. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07106-z.

Zhang:2025:SGR

[534] Zhi-Yuan Zhang, Hao Ren, Hao Li, Kang-Hui Yuan, and Chu-Feng Zhu. Static gesture recognition based on thermal imaging sensors. The Journal of Supercomputing, 81 (4):??, March 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07140-x.

Khanal:2025:DDG

[535] Bikram Khanal and Pablo Rivas. Datadependent generalization bounds for parameterized quantum models under noise. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-06966-9.

Siqueira:2025:APM

[536] Hadley Siqueira and Marcio Kreutz. AccPRET: a proposal of a multicore architecture with reconfigurable accelerators and time predictability. The Journal of Supercomputing, 81 (4):??, March 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07131-y.

Shrestha:2025:SAO

[537] Susav Shrestha, Aayush Gautam, and Narasimha Reddy. Storage access optimization for efficient GPUcentric information retrieval. The Journal of Supercomputing, 81(4): ??, March 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07118-9.

Saif:2025:ERC

[538] Abdu Saif, Nor Shahida Mohd Shah, Yahya M. Al-Moliki, Weiwei Jiang, and Saeed Hamood Alsamhi. Enhancing resilience communication in B5G: optimal deployment of tethered networked flying platforms for disaster recovery. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07098-w.

Li:2025:AMA

[539] Yi Li, Wenying Zheng, and Sai Ji. AI model auditing scheme towards cloud-edge high-performance computing. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07121-0.

Li:2025:RMP

[540] Peng Li, Xinru Zhu, Xiaoshan Li, and Baofeng Huo. Research on mitigating popularity bias in federal recommendation based on users’ behavior. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07144-7.

Chuang:2025:ENI

[541] Po-Jen Chuang and Pang-Yu Huang. Enhancing network intrusion detection by employing Mondrian forests to achieve multiple attack classification. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07123-y.

Pasini:2025:STT

[542] Massimiliano Lupo Pasini, Jong Youl Choi, Kshitij Mehta, Pei Zhang, David Rogers, Jonghyun Bae, Khaled Z. Ibrahim, Ashwin M. Aji, Karl W. Schulz, Jordà Polo, and Prasanna Balaprakash. Scalable training of trustworthy and energy-efficient predictive graph foundation models for atomistic materials modeling: a case study with HydraGNN. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07029-9.

Bozorgi:2025:SRR

[543] Elika Bozorgi, Sakher Khalil Alqaaidi, Afsaneh Shams, Hamid Reza Arabnia, and Krzysztof Kochut. A survey on the recent random walkbased methods for embedding graphs. The Journal of Supercomputing, 81 (4):??, March 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07019-x.

Anjali:2025:DCM

[544] Anjali Anjali and Anjana Gupta. Dynamic cloud model based on decision field theory. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-06989-2. See cor-

rection [590].

Mestre:2025:DLI

[545] Jose I. Mestre, Sergio Barrachina, Darwin Quezada, and Manuel F. Dolz. Deep learning inference optimisation for IoT: Conv2D-ReLU-BN layer fusion and quantisation. The Journal of Supercomputing, 81 (4):??, March 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07107-y.

Ning:2025:TYL

[546] Zhaole Ning, Tianze Zhang, Xin Li, Aiying Wu, and Gang Shi. Tpyolov8: a lightweight and accurate model for traffic accident recognition. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07129-6.

Tang:2025:ADA

[547] Yayuan Tang, Suchi Mishra, Noha Alduaiji, Piyush Kumar Shukla, Mohammad Yahya, and Tao Pang. An advanced data analytics approach to a cognitive cyber-physical system for the identification and mitigation of cyber threats in the medical internet of things (MIoT). The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07093-1.

Zhang:2025:PIM

[548] Yulei Zhang, Guangli Zhu, Yuanyuan Ding, Zhongliang Wei, Lei Chen, and Kuan-Ching Li. A progressive interaction model for multimodal sarcasm detection. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07110-3.

Xu:2025:VBI

[549] Haixia Xu, Yuting Xu, and Kaiyu Hu. A vision-based inspection system for pharmaceutical production line. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07135-8.

Wang:2025:ACR

[550] Zhihang Wang, Jiafei Liu, Chia-Wei Lee, Jingli Wu, and Gaoshi Li. An analysis on component reliability of (n, k)-star networks. The Journal of Supercomputing, 81(4): ??, March 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07128-7.

Wu:2025:AFF

[551] Jun Wu, Yajing Zhang, Tengfei Shan, Zhiwei Xing, Jiusheng Chen, and Runxia Guo. An additive feature fusion attention based on YOLO network for aircraft skin damage detection. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07148-3.

Rafiee:2025:TMS

[552] Mojtaba Rafiee. Toward more secure constructions of flexible multiclient functional encryption schemes. The Journal of Supercomputing, 81 (4):??, March 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07112-1.

Hoorfar:2025:TIE

[553] Hamid Hoorfar, Adam C. Puche, and Istvan Merchenthaler. Thermal image edge detection for AIpowered medical research imaging. The Journal of Supercomputing, 81 (4):??, March 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07050-y. See pub-

lisher correction [603] about missing third author.

Zhou:2025:SYS

[554] Weina Zhou, Yuqi Yang, Ming Zhao, and Wenhua Hu. SCNet-YOLO: a symmetric convolution network for multi-scenario ship detection based on YOLOv7. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07120-1.

Khalid:2025:DLT

[555] Maira Khalid, Jehad Ali, Ahmed Raza Mohsin, Byeong hee Roh, and Mohammed J. F. Alenazi. Deep learning techniques for enhanced security and privacy in 6G terrestrialnonterrestrial network architecture. The Journal of Supercomputing, 81 (4):??, March 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07097-x.

Chai:2025:CCL

[556] Baoqiang Chai, Ruisheng Zhang, Xinyue Li, and Jianxin Tang. CLDE: a competitive learning-driven differential evolution optimization for the influence maximization problem in social networks. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07175-0.

Zhang:2025:CTA

[557] Xiaoyong Zhang and Wei Yue. Collaborative task allocation for heterogeneous UAV coalitions with resource requirements based on multi-genotype genetic algorithm. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07193-y.

Song:2025:IYB

[558] Jian Song, Jie Xie, Qingwang Wang, and Tao Shen. An improved YOLO-based method with lightweight C3 modules for object detection in resource-constrained environments. The Journal of Supercomputing, 81 (5):??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07187-w.

Ying:2025:MSI

[559] Chenghao Ying, Zhiping Zhou, Kewei Li, Zhaozhong Zhang, and Qingshuang Yang. Multi-stream information complementarity network for RGB-D camouflaged object detection. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07201-1.

Aghajani:2025:ESE

[560] Khadijeh Aghajani and Mahbanou Zohrevandi. Enhancing speech emotion recognition: a deep learning approach with self-attention and acoustic features. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07166-1.

Bezerra:2025:CBS

[561] Diego de Freitas Bezerra, Elisson da Silva Rocha, Guto Leoni Santos, André Luis Cavalcanti Moreira, Judith Kelner, Djamel F. H. Sadok, Glauco Estácio Gonçalves, Maria Valéria Marquezini, and Patricia Takako Endo. A cluster-based solution for service function chain allocation in large-scale infrastructure. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07218-6.

Zhang:2025:POM

[562] Zhongliang Zhang, Tao Liu, Chengzhi Wang, Ying Guo, Jingshan Pan, Dawei Zhao, Xiaoming Wu, and Meihong Yang. Parallel optimization of Monte Carlo neutron transport method based on Sunway Bluelight II supercomputer. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07190-1.

Chiyonobu:2025:HSC

[563] Miho Chiyonobu, Masami Takata, Kinji Kimura, and Yoshimasa Nakamura. High-speed computation method for condition numbers in the range restricted general minimum residual method. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06868-2.

Babaiyan:2025:DRL

[564] Vahide Babaiyan and Omid Bushehrian. A deep-reinforcement-learning-based strategy selection approach for faulttolerant offloading of delay-sensitive tasks in vehicular edge-cloud computing. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07196-9.

Qi:2025:IQA

[565] Han Qi, Liyuan Wang, Dianshuo Fu, Abdullah Gani, and Changqing Gong. Improved qARM algorithm for frequent itemsets search in large-scale data sets. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07209-7.

Hou:2025:TMF

[566] Ke Hou, Mengying Wu, Wenjun Wu, and Linhao Huang. Topic mining and forecasting on patent map for GPU technology. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07215-9.

Perez-Sarmiento:2025:PIO

[567] Luis A. Pérez-Sarmiento and Cuauhtemoc Mancillas-López. Parallel implementation of OCB using VAES and GPUs. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07179-w.

Li:2025:KTT

[568] Shuang Li, Wen Yang, Luqun Li, Wenjing Li, Dazhi Li, and Zhehe Wang. Key technologies for task processing in ternary optical computer. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07210-0.

Song:2025:EDP

[569] Jin Song, Bowei Zeng, and Jun Ma. Ellipse detection and positioning in complex vision measurement environments. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07213-x.

Ibrahimoglu:2025:FSN

[570] Nadir İbrahimoğlu, Amjad Osmani, Ali Ghaffari, Faruk Baturalp Günay, Tuğrul Çavdar, and Furkan Yıldız. FootprintNet: a Siamese network method for biometric identification using footprints. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07170-
5.

Bui:2025:CEA

[571] Thiloan Bui, Vanha Tran, Thaigiang Do, Hoangan Le, and Truongminh Ngo. A common and efficient algorithm for discovering high utility co-location patterns and their concise representation from massive spatial data. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07203-z.

Xu:2025:LSC

[572] Zhijie Xu, Yuanyuan Qiu, and Jianqin Zhang. Learnable semantic completion and distance-induced graph matching for domain adaptive object detection. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07208-8.

Kawai:2025:SER

[573] Shunya Kawai, Eriko Maeda, Kazuki Yaguchi, Yasunori Osana, Takefumi Miyoshi, and Hironori Nakajo. Simulation environment for reconfigurable virtual accelerators using a field programmable gate array development environment. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06865-5.

Li:2025:DSE

[574] Cun Li, Shuhai Jiang, and Kangqian Zhou. DYR-SLAM: enhanced dynamic visual SLAM with YOLOv8 and RTAB–Map. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07191-0.

Wei:2025:UDA

[575] Xing Wei, Wenhao Jiang, Fan Yang, Chong Zhao, Yang Lu, Benhong Zhang, and Xiang Bi. Unsupervised domain adaptation via causal-contrastive learning. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07202-0.

Ahmed:2025:JOB

[576] Manzoor Ahmed, Wali Ullah Khan, Mohammad Alamgeer, Eatedal Alabdulkreem, Shouki A. Ebad, Ali M. Al-Sharafi, Ashit Kumar Dutta, and Tahir Khurshaid. Joint optimization for 6G beyond diagonal IRSassisted multi-carrier NOMA vehicleto-infrastructure communication. The Journal of Supercomputing, 81(5): ??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07181-2.

Zhong:2025:CDE

[577] Rui Zhong, Yang Cao, Enzhi Zhang, and Masaharu Munetomo. Competitive differential evolution with knowledge inheritance for singleobjective human-powered aircraft design. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06859-3.

Sun:2025:ERS

[578] Dawei Sun, Yinuo Fan, Chengjun Guan, Jia Rong, Shang Gao, and Rajkumar Buyya. An elastic reconfiguration strategy for operators in distributed stream computing systems. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07216-8.

Chen:2025:YCU

[579] Ruolan Chen, Huibo Zhou, Hui Xie, and Bingyang Wang. YOLO-CE: an underwater low-visibility environment target detection algorithm based on YOLO11. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07205-x.

Song:2025:YFL

[580] Zijing Song, Xiaoyv Zhang, and Panlong Tan. YOLO-Fast: a lightweight object detection model for edge devices. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07172-3. See cor-

rection [664].

Takata:2025:SMR

[581] Masami Takata and Miho Chiyonobu. Similar music recommendation method using Spotify API. The Journal of Supercomputing, 81 (5):??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06864-6.

Liu:2025:ISS

[582] Sicheng Liu, Sichen Tao, Kaiyu Wang, Zhenyu Lei, and Shangce Gao. Improved spherical search algorithm with memory-based dynamic population for optimization. The Journal of Supercomputing, 81 (5):??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07095-z.

Sun:2025:SCM

[583] Jianwen Sun, Bin Zhang, Hongyu Li, Lu Yuan, and Heyu Chang.

T-Sanitation: contrastive masked auto-encoder-based few-shot learning for malicious traffic detection. The Journal of Supercomputing, 81 (5):??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07104-1.

Fang:2025:DDH

[584] Juan Fang, Jingjing Li, and Ziyi Teng. Dhcm: a dynamic hierarchy coordination mechanism for memory optimization. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07224-8.

Kumar:2025:BBF

[585] Ashish Kumar and Kakali Chatterjee. A blockchain-based framework with cryptographic tags for healthcare security. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07183-0.

Tanaka:2025:PES

[586] Tomoaki Tanaka, Michiya Kato, Yasunori Osana, Takefumi Miyoshi, Jubee Tada, Kiyofumi Tanaka, and Hironori Nakajo. Preliminary evaluation of SHAVER: sharing vector registers with an accelerator. The Journal of Supercomputing, 81 (5):??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06863-7.

Yong:2025:SCG

[587] Chen Yong, Hu Defeng, Xu Chao, Chen Nannan, and Liu Jianbo. Smart contract generation model based on code annotation and AST–LSTM tuning. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07186-x.

Li:2025:ASF

[588] Hongjian Li, Gangfan Tan, Chenzi Wang, Yuan Gao, Tao Zhou, and Xiaolin Duan. Adaptive scheduling framework of streaming applications based on performance-to-cost ratio in heterogeneous cloud environment. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07207-9.

Wang:2025:ITA

[589] Xun Wang, Xianxi Zhu, Xiangyu Meng, Zeyang Zhu, Siyu Zhang, and Tao Song. Inference and training acceleration of deep learning partial differential equation solver. The Journal of Supercomputing, 81 (5):??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07204-y.

Anjali:2025:CDC

[590] Anjali and Anjana Gupta. Correction: Dynamic cloud model based on decision field theory. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07286-8. See [544].

Wang:2025:DND

[591] Nianhao Wang and Han Wang. Dcsca-Net: a dual-branch network with enhanced cross-fusion and spatial-channel attention for precise medical image segmentation. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07239-1.

Baawi:2025:HSC

[592] Salwa Shakir Baawi, Ekram Hakem, Abdulkareem A. Al-Hamzawi, Dhiah Al-Shammary, Ayman Ibaida, and Ahmed M. Mahdi. Hilbert similarity convex for efficient EEG feature selections. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07199-6.

Topazal:2025:IDD

[593] S M Topazal, Shayla Islam, Raenu A./L. Kolandaisamy, Mohammad Kamrul Hasan, Ahmad Fadzil Ismail, Nur Hanis Sabrina Suhaimi, Huda Saleh Abbas, Muhammad Attique Khan, and Kamal Ali Alezabi. Intelligent device to device handover management techniques for 5G/ 6G and beyond. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07182-1.

Du:2025:DDC

[594] Yongping Du, Qi Zhang, Shuyi Fu, Ying Hou, and Honggui Han. DCLMD: dynamic clustering and label mapping distribution for constructing in-context learning demonstrations. The Journal of Supercomputing, 81 (5):??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07214-w.

Huang:2025:TFD

[595] Yibo Huang, Weidong Qin, Zhiyong Li, and Qiuyu Zhang. Time-frequency dual-domain attention for acoustic echo cancellation. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07200-2.

Du:2025:DRP

[596] Jinyang Du, Renyun Liu, Du Cheng, Qingliang Li, and Fanhua Yu. Deep recognition of partial differential equations based on reinforcement learning and genetic algorithm. The Journal of Supercomputing, 81(5): ??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07162-5.

Arasteh:2025:AST

[597] Bahman Arasteh, Keyvan Arasteh, and Ali Ghaffari. An automatic software test-generation method to discover the faults using fusion of machine learning and horse herd algorithm. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07219-5.

Malekzadeh:2025:DPI

[598] Saba Malekzadeh, Saleh Yousefi, and Mir Saman Tajbakhsh. DDos prevention in IoT networks by analyzing source-side inter-bot traffic using deep learning techniques. The Journal of Supercomputing, 81 (5):??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07236-4.

Peng:2025:PML

[599] Chenchen Peng, Zeyu Xia, Shengbo Tang, Yifei Guo, Canqun Yang, Tao Tang, and Yingbo Cui. PVGwfa: a multi-level parallel sequenceto-graph alignment algorithm. The Journal of Supercomputing, 81(5): ??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07184-z.

Wang:2025:ADS

[600] Yingqing Wang, Yanhua Liang, Yue Huang, and Guihe Qin. An automated data stream analysis framework for Internet of Vehicles based on online ensemble learning and two-dimensional fractal dimension. The Journal of Supercomputing, 81 (5):??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07227-5.

Shang:2025:VAS

[601] Jiandong Shang, Zhuxin Wen, Haobo Hua, Hengliang Guo, Gang Wu, Wenlong Fan, Yang Guo, and Guangsheng Qin. VBATS: an adaptive strategy for grouped GEMM on GPUs. The Journal of Supercomputing, 81 (5):??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07195-w.

Ranilla-Cortina:2025:EHP

[602] Sandra Ranilla-Cortina, Pedro Alonso-Jordá, Jesús Vigo-Aguiar, and José Ranilla. The evolution of highperformance computing: how AI and quantum computing are reshaping supercomputing. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07241-
7.

Hoorfar:2025:PCT

[603] Hamid Hoorfar, Adam C. Puche, and Istvan Merchenthaler. Publisher correction: Thermal image edge detection for AI-powered medical research imaging. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07257-z. See [553].

Azarkasb:2025:MTL

[604] Seyed Omid Azarkasb and Seyed Hossein Khasteh. The mapping trick: leveraging RoboSoccer obstacle avoidance and navigation for advanced task scheduling solutions in foggy IoE ecosystems. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07147-4.

Zhang:2025:AYS

[605] Hongying Zhang, Wentao Liu, and Enyao Chen. ASOD-YOLOX: a study on small object detection in aerial images based on YOLOX. The Journal of Supercomputing, 81 (5):??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07243-5.

Yin:2025:IDH

[606] Jie Yin, Shuning Zhang, and Guanlong Deng. An improved discrete Harris Hawks optimization algorithm for the no-wait job shop problem to minimize total weighted tardiness. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07234-6.

Meng:2025:HBV

[607] Lei Meng, Yueqiang Xu, and Fuhong Lin. How to balance the verification burden: a multi-hierarchical aggregate signature for drone swarms. The Journal of Supercomputing, 81 (6):??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07222-w.

Verma:2025:PML

[608] Gurudatta Verma and Tirath Prasad Sahu. PSI-MFS: lightweight multiobjective feature selection for enhanced multi-label classification. The Journal of Supercomputing, 81(6): ??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07163-4.

Ding:2025:ITA

[609] Zhiyi Ding, Lei Sun, and Ruiyang Ding. Improving the transferability of adversarial examples via the high-level interpretable features for object detection. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07225-7.

Jain:2025:SAR

[610] Sarika Jain, Priyanka Sukul, and Sven Groppe. A scientometric analysis of reviews on the Internet of Things. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07230-w.

Cao:2025:DNB

[611] Buqing Cao, Yating Yi, and Bing Tang. A deep Q-network-based edge service offloading in cloud-edge-terminal environment. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07228-4.

Yang:2025:TEP

[612] Kang Yang, Lili Zhang, and Pei Yu. Traffic environmental protection edge computing: a monitoring algorithm and system of truck black smoke emission in complex scene. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07276-w.

Xue:2025:SRB

[613] Peng Xue, Qian Gao, and Jun Fan. Social recommendation based on contrastive learning of hypergraph convolution. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07143-8.

Dhrir:2025:MLD

[614] Hanen Dhrir, Maha Charfeddine, and Habib M. Kammoun. Machine learning- and deep learning-based anomaly detection in firewalls: a survey. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07212-y.

Ayyadi:2025:ACR

[615] Armin Ayyadi and Arezoo Jahani. Addressing cost and resource variability for big data task scheduling in heterogeneous cloud environments. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07248-0.

Park:2025:ABA

[616] Ju-Won Park and Taeyoung Hong. Application I/O behavior analysis on leadership cluster system. The Journal of Supercomputing, 81(6): ??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07235-5.

Koiso:2025:PMA

[617] Norie Koiso, Yuki Takemoto, and Masami Takata. Proposed method of acquiring train data for early-modern Japanese printed character recognizers. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06866-4.

Deng:2025:HAY

[618] Yan Deng, Lidong Huang, and Shunxiang Shi. A heterogeneous attention YOLO model for traffic sign detection. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07256-0.

Sui:2025:MAE

[619] Xiao-Fei Sui, Si-Wen Zhang, and Xue-Lian Bai. Mayfly algorithm with elementary functions and mathematical spirals for task scheduling in cloud computing system. The Journal of Supercomputing, 81(6): ??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07231-9.

Wang:2025:PPS

[620] Guangchao Wang, Zhaojun Li, and Changyou Guo. PDPA: a part-of-speech-driven prompt and attention method for aspect sentiment triplet extraction. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07188-9.

Bai:2025:RMA

[621] Fenhua Bai, Hongye Xu, and Chi Zhang. RBC-MSS: asynchronous broadcasting protocol based on multisecret sharing. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07211-
z.

Li:2025:MMO

[622] Wei Li, Zeming Ping, and Lei Wang. Multimodal multiobjective optimization algorithm with fine-grained special crowding distance and dualspace selection mechanism. The Journal of Supercomputing, 81(6): ??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07221-x.

Cuevas:2025:NMA

[623] Erik Cuevas, Oscar A. González-Sánchez, and Alma Rodríguez-Vazquez. A novel metaheuristic algorithm using structured population and virtual particles. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07194-x.

Yang:2025:ISO

[624] Yahao Yang, Yu Liu, and Zhenlong Zhao. An improved snake optimization algorithm based on hybrid strategy. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07258-y.

Buttner:2025:APP

[625] Markus Büttner, Christoph Alt, and Vadym Aizinger. Analyzing performance portability for a SYCL implementation of the 2D shallow water equations. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07063-7.

Serajgah:2025:PFT

[626] Saman Amini Serajgah and Ali-Akbar Ahmadi. Prescribed finitetime adaptive fuzzy-based faulttolerant control of robotic manipulators using dynamic scaling factor. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07270-2.

Li:2025:UTL

[627] Min Li, Shuming Zhou, and Gaolin Chen. A unified temporal link prediction framework based on nonnegative matrix factorization and graph regularization. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07217-7.

Nazir:2025:EEE

[628] Ahsan Nazir, Jingsha He, and Muhammad Salman Pathan. Empirical evaluation of ensemble learning and hybrid CNN–LSTM for IoT threat detection on heterogeneous datasets.

The Journal of Supercomputing, 81 (6):??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07255-1.

Alsofyani:2025:ECS

[629] May Alsofyani and Liqiang Wang. Evaluating ChatGPT’s strengths and limitations for data race detection in parallel programming via prompt engineering. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07237-3.

Kang:2025:EFX

[630] Sung Won Kang and Cheong Hee Park. Effective federated XG-Boost learning for multi-class classification in Non–IID environments. The Journal of Supercomputing, 81 (6):??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07245-3.

Li:2025:FDF

[631] Boyang Li, Huihuang Zhao, and Leyi Li. FDPNet: Deep forgery detection by leveraging multi-scale self-forgery images generating. The Journal of Supercomputing, 81(6): ??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07229-3.

Gao:2025:MSF

[632] Jialin Gao, Liqiang Zhu, and Yao Wang. Multi-scale feature enhanced detection of foreign object intrusions on railways. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07254-2.

Zhang:2025:ESI

[633] Yong Cai Zhang, Yonghua Lin, and Yih Bing Chu. The ensemble of self-information-based feature selection for heterogeneous data via knearest neighborhood rough set model. The Journal of Supercomputing, 81 (6):??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07244-4.

Uzelaltinbulat:2025:DSU

[634] Selin Uzelaltinbulat, Yasemin Kucukciloglu, and Boran Sekeroglu. Diagnosis of sacroiliitis using MR images with a simplified custom deep learning model. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07280-0.

Lashkari:2025:UEM

[635] Arash Habibi Lashkari, Mohammad-Moein Shafi, and Ashley Barkworth. Unveiling evasive malware behavior: toward generating a multi-sources benchmark dataset and evasive malware behavior profiling using network traffic and memory analysis. The Journal of Supercomputing, 81 (6):??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07267-x.

Salehpour:2025:OID

[636] Arash Salehpour, Mohammad Ali Balafar, and Alireza Souri. An optimized intrusion detection system for resource-constrained IoMT environments: enhancing security through efficient feature selection and classification. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07253-3.

Javeed:2025:UCD

[637] Muhammad Awais Javeed, Dawei Li, and Muhammad Awais Ashraf. Urban city data delivery optimization in VNDN using a route-based caching approach with content-type awareness in intelligent transportation system. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07250-6.

Tsai:2025:MSD

[638] Yu-Shiuan Tsai, Chia-Tung Tsai, and Jian-Hong Huang. Multi-scale detection of underwater objects using attention mechanisms and normalized Wasserstein distance loss. The Journal of Supercomputing, 81 (6):??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07251-5.

Li:2025:ATS

[639] Wenjuan Li, Genyuan Yang, and Qiwen Ni. Adaptive two-stage task offloading based on meta reinforcement learning for mobile edge computing. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07274-y.

Meng:2025:MFF

[640] Chao Meng, Yuhong Du, and Guangyu Dong. A multi-feature fusion supervoxel clustering segmentation method based on energy function. The Journal of Supercomputing, 81(6): ??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07220-y.

Battat:2025:TDO

[641] Nadia Battat and Abdallah Makhoul. A trust-driven optimization model for reliable authorization in Hadoop Environment. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07268-w.

Luo:2025:VLM

[642] Wenyu Luo, Changxing Hou, and Annan Xuan. Vision and LiDAR multimodal fusion beam prediction method for millimeter-wave communication system. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07284-w.

Emara:2025:MOT

[643] Farouk A. Emara, Ahmed A. A. Gad-Elrab, and M. A. Abd El-Baky. Multiobjective task scheduling algorithm for load balancing in cloud computing based on improved Harris hawks optimization. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07091-3.

Cai:2025:MTL

[644] Yuanyuan Cai, Fei Tong, and Haitao Xiong. Multi-task learning framework using tri-encoder with caption prompt for multimodal aspect-based sentiment analysis. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07252-4.

Sun:2025:NID

[645] Xiaolong Sun, Zhengyao Gu, and Junwei Ye. Network intrusion detection based on feature fusion of attack dimension. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07271-1.

Yuan:2025:RAS

[646] Yan Yuan, Jinwei Xu, and Yong Wang. Research and application of smart insole assisted gait recognition technology. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07226-6.

Shankani:2025:FMN

[647] Rana Shankani, Maedeh Khalifavi, and Amirhossein Nikoofard. F^{2} multisense: a novel approach to fuzzy fusion in multisensor data to improve saving energy in WBSN. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07233-7.

Almeida:2025:AIH

[648] Fernando Almeida and Edet Okon. Assessing the impact of high-performance computing on digital transformation: benefits, challenges, and size-dependent differences. The Journal of Supercomputing, 81(6): ??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07281-z.

Yang:2025:ANC

[649] Tianyu Yang, Yanmei Hu, and Xiangtao Li. An adaptive network construction for single-cell clustering. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07263-1.

Souza:2025:MAE

[650] Fillipe Souza, Eduardo Tavares, and Carlos Araújo. A modelling approach for estimating energy consumption of NoSQL-based storage systems. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07298-4.

Dhiman:2025:SAE

[651] Shalini Dhiman, Sumitra Nayak, and Swarnendu Kumar Chakraborty. A survey analyzing encryption schemes for IoT security measures. The Journal of Supercomputing, 81(6): ??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07285-9.

Lu:2025:SPL

[652] Hangqing Lu, Quan Yan, and Li Wang. A single-point and lengthrepresentation-based model for nested named entity recognition. The Journal of Supercomputing, 81(6): ??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07240-8.

Chen:2025:PSC

[653] Yulei Chen, Dongwei Guo, and Jianhua Chen. A provably secure certificateless online/offline signature scheme for internet of things. The Journal of Supercomputing, 81 (6):??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07294-8.

Wang:2025:SFS

[654] Zhicheng Wang, Longge Wang, and Jinhu Wu. Semantic feature space construction for unsupervised fewshot image classification. The Journal of Supercomputing, 81(6): ??, April 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07315-6.

Xue:2025:RRR

[655] Qifan Xue, Feng Yang, and Weigong Zhang. Rethink Reynolds’ rules: flockinspired network for vehicle trajectory prediction. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07265-z.

Diao:2025:LCD

[656] Zhaoshuo Diao, Yan Zhang, and Zongzhe Sun. Learnable confidencedriven asymmetric attention fusion mechanism for PET/oCT tumor segmentation. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07313-8.

Hashemi:2025:AMA

[657] Amin Hashemi, Mohammad Bagher Dowlatshahi, and Parham Moradi. AE-MCDM: an autoencoder-based multi-criteria decision-making approach for unsupervised feature selection. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07316-5.

Singh:2025:CMM

[658] Kamal Singh, Nitin Singha, and Swati Bhalaik. CCLNet: multiclass motor imagery EEG decoding through extended common spatial patterns and CNN-LSTM hybrid network. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07319-2.

Cam:2025:FCD

[659] Nguyen Tan Cam and Dinh Hoai Hiep. Fraud calls detection using classimbalanced learning on graph structures. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07259-x.

Chen:2025:TPF

[660] Wei Chen, Jiao Li, and Zhenfeng Chen. Two peak-finding algorithms for twodimensional unimodal symmetric signals based on mirroring and interpolating. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07264-0.

Ren:2025:HBA

[661] Shuyang Ren and Choonhwa Lee. A hierarchical blockchain architecture for federated learning in edge computing networks. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07262-2.

Liu:2025:DYD

[662] Houhong Liu, Fei Tan, and Yuxin Jin. Dual-YOLO: dual-path UAV aerial image target detection algorithm. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07290-y.

Barina:2025:IVL

[663] David Barina. Improved verification limit for the convergence of the Collatz conjecture. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07337-0.

Song:2025:CYF

[664] Zijing Song, Xiaoyu Zhang, and Panlong Tan. Correction: YOLO-Fast: a lightweight object detection model for edge devices. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07321-8. See [580].

Zhang:2025:AIP

[665] Lieping Zhang, Yao Chen, and Cui Zhang. An algorithm for indoor positioning based on LightGBM + ExtraTrees chaotic weighted ensemble: evaluation and comparison. The Journal of Supercomputing, 81 (7):??, May 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07291-x.

Thakkar:2025:MME

[666] Riddhi Thakkar and Madhuri Bhavsar. MeghMesa: a multilevel elasticity for streaming applications in cloud. The Journal of Supercomputing, 81 (7):??, May 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07232-8.

Liu:2025:ISO

[667] Fangjun Liu, Huifu Zhang, and Shangfeng Mo. Instruction selection optimization for VLIW architecture based on classification node merging. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07323-6.

Zhou:2025:EAL

[668] Jinfeng Zhou, Xiaoqin Zeng, and Haoran Zhu. Enhancing aspect-level sentiment analysis through the integration of local context interdependencies and syntactic quality compensation. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07301-y.

Eldakhly:2025:OML

[669] Nabil M. Eldakhly. Optimized machine learning for real-time, noninvasive blood pressure monitoring. The Journal of Supercomputing, 81 (7):??, May 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07330-7.

Feng:2025:TTM

[670] Chunhui Feng, Yongxiang Zhong, and Xiaolong Liu. TMTC: trusted multimodal transformer classification framework for video frame deletion detection. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07307-6.

Liu:2025:MMI

[671] Quankai Liu, Haifeng Sang, and Wangxing Chen. MGFormer: a multiinformation-based GRU-transformer network for pedestrian trajectory prediction. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07304-9.

Liu:2025:NFT

[672] Wenfei Liu, Jiafei Liu, and Qi Wang. A novel fault-tolerant technique for star graph-based interconnection networks. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07272-0.

Hong:2025:PUP

[673] Hanshu Hong, Yibo Sun, and Zhixin Sun. PB-UOKM: a policy-based updatable oblivious key management scheme for secure and practical data sharing in remote storage. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07332-5.

Dixit:2025:QBC

[674] Vimal Dixit and Krishnan Rajkumar. QENN: breast cancer prediction using quantum-enhanced neural network. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07331-6.

BenTaher:2025:GFQ

[675] R. Ben Taher and M. Lassri. On the generalized Fibonacci quaternions and the linear difference equations with periodic coefficients. The Journal of Supercomputing, 81(7): ??, May 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07283-x.

Xia:2025:CCL

[676] Limin Xia and Qiyue Xiao. Comprehensive context learning for twostage human-object interaction detection. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07308-5.

Erho:2025:TPJ

[677] Joseph Agaroghenefuoma Erho, Bunakiye Richard Japheth, and Juliana Iworikumo Consul. Theoretical perspective to the juggling sequence rotation performance pattern from system memory locality reference viewpoint. The Journal of Supercomputing, 81(7): ??, May 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07261-3.

Ao:2025:AVM

[678] Wenxiu Ao, Zhongmei Wang, and Liang Zheng. An audio-visual multimodal adaptive balanced learning method based on gradient modulation. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07311-w.

Misra:2025:QQE

[679] Shreshtha Misra and Poonam Rani.

QEKLR: quantum-enhanced kernel logistic regression for classification. The Journal of Supercomputing, 81 (7):??, May 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07266-y.

Xu:2025:ETO

[680] Tao Xu, Chaoyue Chen, and Dongdong Ma. Exponential-trigonometric optimization algorithm with multi-strategy fusion for UAV three-dimensional path planning. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07335-2.

Osaba:2025:EAQ

[681] Eneko Osaba, Esther Villar-Rodriguez, and Izaskun Oregi. Exploring the application of quantum technologies to industrial and real-world use cases. The Journal of Supercomputing, 81 (7):??, May 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07320-9.

Liu:2025:SAD

[682] Tianwei Liu and Jingwen Luo. Semantic-aware dynamic scene tracking and reconstruction for mobile robots leveraging visual SLAM. The Journal of Supercomputing, 81(7): ??, May 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07343-2.

Wang:2025:DAT

[683] Yuan Wang, Turdi Tohti, and Askar Hamdulla. Domain-adaptive transfer network for visual-textual crossdomain sentiment classification. The Journal of Supercomputing, 81(7): ??, May 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07273-z.

Li:2025:DUC

[684] Xinxin Li, Zhenping Lan, and Aixia Yuan. DPF-Unet: a CNN-swin transformer fusion network for 3D brain tumor segmentation in MRI images. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07339-y.

Khan:2025:AZT

[685] Talha Ahmed Khan, Khizar Abbas, and Wang-Cheol Song. Accelerating zero-touch automation and optimization of beyond 5G services: deep learning and intent-based networking fusion. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07260-4.

Khomami:2025:CGGb

[686] Mohammad Mehdi Daliri Khomami, Ali Mohammad Saghiri, and Mohammad Reza Meybodi. Cellular goore game with multiple learning automata in each cell and its applications. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07249-z.

Chen:2025:ITA

[687] Jialong Chen, Jijun Tong, and Shudong Xia. An implicit tubular-aware network for coronary artery segmentation. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07312-9.

Dong:2025:OKE

[688] Yuyu Dong, Fang’ai Liu, and Xuejian Gao. Optimizing keyphrase extraction with dependency relationaware attention graph convolutional networks. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07303-w.

Xu:2025:LTR

[689] Nengxiang Xu, Yuxiang Chen, and Nam Ling. N-Lock: a transactionreleased shard reconfiguration protocol with zero-knowledge proof. The Journal of Supercomputing, 81 (7):??, May 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07317-4.

Wang:2025:MET

[690] Bing Wang, Chaoqi Cai, and Youming Zhang. MapsTSF: efficient traffic prediction via hybrid Mamba 2transformer spatiotemporal modeling and cross adaptive periodic sparse forecasting. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07328-1.

Sun:2025:MID

[691] Bo Sun, Zhuo Hao, and Jun He. Mutual introspective distillation for unbiased scene graph generation. The Journal of Supercomputing, 81 (7):??, May 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07246-2.

Li:2025:YEO

[692] Yanshun Li, Quanfeng Zheng, and Mengwei Guo. YOLO-EDGE: an object detection algorithm for traffic scenarios. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07275-x.

Bain:2025:OAT

[693] Debanjali Bain and Biswanath Dutta. Ontological approach towards discovering and recommending COVID-19 therapeutics, risk factors, and drug interactions. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07302-x.

Bakhoda:2025:MOS

[694] Saeedeh Bakhoda, Mohammad Abdollahi Azgomi, and Mohammad Reza Ebrahimi Dishabi. Multiobjective scheduling of heterogeneous parallel systems using the VITS algorithm. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07293-9.

Zhang:2025:ACD

[695] Zheqing Zhang, Hongzhi Li, and Kuan-Ching Li. AAMB: a crossdomain identity authentication scheme based on multilayered blockchains in IoMT. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07309-4.

Wang:2025:DDF

[696] Feng Wang, Chenglong Wang, and Yuxuan Zhang. Data-driven fuzzy logic control method for improved USV path planning. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07318-3.

Li:2025:ICO

[697] Lang Li, Li Hu, and Jin Li. Inverse correction-optimized vertical federated unlearning. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07310-x.

Li:2025:SST

[698] Xiguang Li, Yuhan Zhu, and Liang Zhao. STGEN: spatio-temporal generalized aggregation networks for traffic accident prediction. The Journal of Supercomputing, 81(7): ??, May 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07288-6.

Wang:2025:SIC

[699] Jun Wang, Chenhao Sun, and Xiaoli Li. Spatiotemporal information cooperative interaction network for video salient object detection. The Journal of Supercomputing, 81 (7):??, May 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07314-7.

Czarnul:2025:ORA

[700] Paweł Czarnul, Marcel Antal, and Gordana Rakić. Optimization of resource-aware parallel and distributed computing: a review. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07295-
7.

He:2025:OMB

[701] Yunbin He, Chenglong Li, and Fuwei Guo. An oversampling method based on the weighting of minority class clusters. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07325-4.

Yaqoubi:2025:MCS

[702] Zahra Yaqoubi, Behrouz Shahgholi Ghahfarokhi, and Mojtaba Mahdavi. Mobile crowd sensing based spectrum monitoring with privacy protection and malicious behavior detection using hyperledger fabric and identity mixer. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07300-z.

Gharehbagheri:2025:APS

[703] Parviz Gharehbagheri, Hamid Haj Seyyed Javadi, and Hamidreza Bakhshi. An alternative proposed single error detection algorithm for the single parity check algorithm. The Journal of Supercomputing, 81(7): ??, May 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07305-8.

Ashi:2025:IDG

[704] H. A. Ashi and Noufe H. Aljahdaly. Investigation of the damped geophysical KdV equation using the explicit exponential time differencing method. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07289-5.

Saad:2025:EMO

[705] Abla Saad, Osama Abdel-Raouf, and Ahmed Kafafy. Enhanced multi-objective scheduling for heterogeneous computing platforms using hybrid MOEAD with SA and TS-guided initialization strategies. The Journal of Supercomputing, 81 (7):??, May 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07278-8.

Zhan:2025:CDR

[706] MengJun Zhan, XiaoYao Xie, and ZhongYuan Guo. Change detection of remote sensing building images based on dual-domain strip attention DML–UNet network. The Journal of Supercomputing, 81(7): ??, May 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07344-1.

Alfawaz:2025:VFC

[707] Oruba Alfawaz, Ali A. El-Moursy, and Ahmed M. Khedr. VFCkM: a federated clustering framework based on k-means algorithm for vertically partitioned data with shared attributes. The Journal of Supercomputing, 81 (7):??, May 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07296-6.

Verma:2025:AVL

[708] Priyanka Verma, Donna O’Shea, and John G. Breslin. ABIDS-VEM: leveraging an equilibrium optimizer and data ramification in association with ensemble learning for anomalybased intrusion detection system. The Journal of Supercomputing, 81 (7):??, May 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07292-w.

Masaud:2025:DCS

[709] Mohamed Ahmed Ali Masaud, Selçuk Alparslan Avci, and Javad Rahebi. Detecting cyberattacks in smart grids using VGG-16 and whale–fisher mantis optimization algorithm (WOA-FMO). The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07345-0.

Florez:2025:LCP

[710] Arantzazu Florez, Itsaso Rodríguez-Moreno, and Basilio Sierra. Lynxsight: change-point detection through different distance-based common spatial patterns. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07282-y.

Liang:2025:CPC

[711] Bo Liang, Fujiang Yuan, and Jie Gao. Cs-pbft: a comprehensive scoring-based Practical Byzantine Fault Tolerance consensus algorithm. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07342-3.

He:2025:NCD

[712] Jiali He and Mideth B. Abisado. A novel conditional discrimination index approach for feature selection in partially labeled hybrid data. The Journal of Supercomputing, 81 (7):??, May 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07333-4.

Shang:2025:HWT

[713] Lei Shang, Huan Lei, and Wenyuan Yang. HIFNet: wavelet transformenhanced UAV object detection in complex conditions. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07338-
                                        z.

Derakhshan:2025:HNM

[714] Mohammad Hossein Derakhshan, Yadollah Ordokhani, and J. F. Gómez-Aguilar. A hybrid numerical method with high accuracy to solve a timespace diffusion model in terms of the Caputo and Riesz fractional derivatives. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07287-7.

Nooraee:2025:TPO

[715] Mohsen Nooraee, Hamidreza Ghaffari, and Fateme Zarisfi Kermani. Tiny-ParsBERT: an optimized hybrid model for efficient sentiment analysis in Persian texts. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07297-5.

Zhang:2025:MSE

[716] Mengyuan Zhang, Huifang Ma, and Qibin Zhang. Multiscale sessionenhanced long time series modeling for power transformer oil temperature prediction. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07269-9.

Leng:2025:FFE

[717] Yuhe Leng, Lei Xu, and Chungen Xu. FedSSU: flexible and efficient decentralized unlearning for federated learning. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07478-2.

Zhao:2025:MBS

[718] Huaqi Zhao, Zhengguang Lu, and Guojing Li. A multi-branch semantic segmentation method for autonomous driving. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07453-x.

Wu:2025:LGA

[719] Jin Wu, Zhengdong Su, and Wenxuan Luo. Leech Growth Algorithm: a new meta-heuristic algorithm. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07387-4.

Wu:2025:MVT

[720] Yasu Wu, Changlong Fu, and Cheng Xie. Multi-view temporal graph neural network for numerous miniature cascade popularity prediction. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07484-4.

Tiwari:2025:YTB

[721] Satyarth Tiwari and Sanjay Kumar Sharma. YOLOv7-tiny-based lightweight and efficient algorithm for photovoltaic cell crack detection. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07493-3.

Sun:2025:SNC

[722] Hao Sun, Xiaochuang Bai, and Ziyu Hu. Solving non-cyclic dynamic multi-objective optimization problems via GRU prediction and multi-information hybrid exploration. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07471-9.

Tang:2025:NSM

[723] Zhenhua Tang, Haiyan Yin, and Chao Ding. Numerical simulation and multi-objective optimization of braking wear of multi-piston brake pads. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07501-6.

Chen:2025:TTM

[724] Jinguang Chen, Hongxiao Yao, and Lili Ma. TMATrack: token merging for autoregressive visual object tracking. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07498-y.

Wang:2025:MDS

[725] Xianchuan Wang, Ben Lu, and Xianchao Wang. Multi-document summarization through subsectionaware pre-training objectives. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07504-3.

Zhu:2025:FCS

[726] Linying Zhu, Shanshan Wang, and Yi Zhang. FastMS-CDA: speeding up adversarial sample generation with multistage diffusion model. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07480-8.

Chen:2025:AYL

[727] Xi Chen and Guohui Wang. AF-YOLOv8: a lightweight algorithm suitable for helmet and vest detection. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07496-0.

Sheng:2025:CTC

[728] Yunlong Sheng, Xuye Zhuang, and Siting Yang. A combinatorial test case prioritization method based on the quadratic network with wide multilayer kernels. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07459-5.

Yang:2025:MBC

[729] Han Yang, Siwei Hu, and Xue-Jing Zhao. Motif-based community detection in multiplex networks with edge deletion optimization. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07489-z.

Zhang:2025:ARW

[730] Tuyi Zhang, Sanmin Liu, and Wentao Yu. Adaptive resampling and weighted ensemble method for dynamic imbalance data stream classification. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07482-6.

Atici:2025:DAM

[731] Sinan Atici and Gurkan Tuna. A deterministic approach to modelling the dynamics of vulnerabilities, cyberattacks, and security hardening using Lotka–Volterra equations. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07393-6.

Wan:2025:MNL

[732] Miao Wan, Y. C. Lin, and Yu-Liang Qiu. MHAED-Net: a lightweight multiscale hybrid attention encoderdecoder network for the efficient segmentation of industrial forging images. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07456-8.

Cai:2025:BLL

[733] Ye Cai, Jinghui Zhong, and Hao Liao. BPTCN: a low-latency branch prediction model based on temporal convolutional networks. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07490-6.

Daemdoost:2025:GOA

[734] Amin Daemdoost, Asadollah Shahbahrami, and Reza Hassanpour. GPU optimizations to accelerate an image enhancement algorithm. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07447-9.

Jiang:2025:MMM

[735] Yuncong Jiang, Weifei Jia, and Quanlin Yu. MAFFNet: a multimodal adaptive feature fusion net for signal modulation recognition. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07492-4.

Jain:2025:OBN

[736] Kirti Jain and Rajni Jindal. Optimizationbased noise filtering among usercentric tweets to improve predictions in recommendation system. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07457-7.

Xu:2025:EPO

[737] Caixu Xu, Hui Guo, and Jie He. Efficient program optimization through knowledge-enhanced LoRA fine-tuning of large language models. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07378-5.

Wu:2025:RLF

[738] Runfang Wu, Ye Du, and Yu Tang. A research landscape on formal verification of UML dynamic modeling. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07464-8.

Gohel:2025:HPG

[739] Prashant Gohel and Manjunath Joshi. Hybrid and parallel GAN architecture with non-IID noise input. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07495-1. See cor-

rection [1030].

Xu:2025:DMS

[740] Changlin Xu, Rong Yang, and Juhong Shen. Divergence measure of spherical fuzzy sets and their applications in MAGDM based on TOPSIS method. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07468-4.

Wen:2025:UUF

[741] Yongjun Wen, Zhihao Cui, and Lijun Tang. UCP: a unified framework for code generation with pseudocode-based multi-task learning and reinforcement alignment. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07487-1.

Ma:2025:OEE

[742] Changxi Ma, Mingxi Zhao, and Tao Wang. Optimization of energyefficient control in rail transit systems under event impact. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07513-2.

Liu:2025:NPS

[743] Peiwen Liu, Mengnan Tian, and Xinduan Wang. A novel particle swarm optimization with individual-based adaptive learning strategy for numerical optimization. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07352-1.

Xu:2025:NDP

[744] Yuanhang Xu, Yan Wang, and Qi He. Node-disjoint paths in k-ary n-cube with optimal maximum path length. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07454-w.

Tian:2025:MUM

[745] Xianzhong Tian, Yucheng Li, and Ziyi Lin. Multi-user multi-exit DNN inference partitioning and task scheduling strategy. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07508-z.

Al-Shalabi:2025:FAO

[746] Mohammed Al-Shalabi, Mohammed A. Mahdi, and Sawsan A. Saad. Feasibility analysis and opposition white shark optimizer for optimizing modified EfficientNetV2 model for road crack classification. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07509-y.

Lin:2025:KDF

[747] Xiaogang Lin, Xiaoli Zhao, and Yincan Shu. Knowledge-driven federated learning: a systematic literature review on approaches, challenges, and prospects. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07516-z.

Shin:2025:DFA

[748] Yeonghun Shin, Geon Yu, and Taeshik Shon. Digital forensic analysis for vehicle infotainment systems based on packet fingerprinting. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07552-9.

Zhao:2025:SAF

[749] Peng Zhao, Zhuhua Liao, and Aiping Yi. A semi-asynchronous federated learning method integrating personalization and staleness awareness for traffic flow prediction in dynamic Internet of Vehicles. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07523-0.

Meng:2025:RTI

[750] Kun Meng, Shengjie Yang, and Li Yuan. A real-time image dehazing algorithm based on adaptive routing. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07549-4.

Chen:2025:DAN

[751] Yilin Chen, Bo Gao, and Qiong Wang. A dual-archive niche with twostage directed differential evolution for multimodal multi-objective optimization. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07517-y.

Cui:2025:UUS

[752] Zhaojun Cui, Wei Tian, and Shengqin Jiang. UST-SU: a U-shaped video prediction network based on partial autoregression. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07486-2.

Ouyang:2025:FLL

[753] Yiming Ouyang, Zhuoxuan Ji, and Huaguo Liang. Fabdb: a low-latency fault-tolerant architecture based on dynamic bypass for network-on-chip. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07530-1.

Chen:2025:ADF

[754] Haowen Chen, Hao Liu, and Ruyu Wang. AFF-DSTnet: fabric anomaly detection based on adaptive feature fusion dual-student-teacher network. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07506-1.

BahraniPour:2025:EMO

[755] Fatemeh BahraniPour, Mohammad Farshi, and Sepehr Ebrahimi Mood. Enhanced multi-objective cuckoo search with migration operator for benchmark optimization and IoT task scheduling in cloud-fog computing. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07531-0.

Ding:2025:ATB

[756] Ning Ding and Knut Möller. Adversarial training with borderline samples. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07477-3.

Chai:2025:MNR

[757] Jinhao Chai and Jingfang Shen. MixeRF: neural radiance fields with multidimensional feature mixing capability. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07441-1.

Dai:2025:OMA

[758] Enhui Dai, Zhenxue He, and Xiang Wang. An optical microscope algorithm with precise focusing strategy and migration strategy with application in 3D path planning. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07460-y.

Xu:2025:DFB

[759] Haiyan Xu, Jing Ren, and Guanying Huo. Dynamic frame-based weight estimation and joint position-scale model for underwater object tracking. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07537-8.

Wu:2025:LPI

[760] Shuo-Jye Wu, Yao-Ting Tseng, and Coskun Kus. Likelihood and pivotal inference for Kumaraswamy parameters under progressive type-II censoring. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07514-1.

Zhang:2025:SSM

[761] Haibin Zhang, Tingting Lang, and Yanjun Ji. State-space models and depth separable convolution for highprecision hybrid wafer defect detection. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07521-2.

Li:2025:MNC

[762] Yuan Li, Jue Wang, and Jinzhang Li. MEDU-Net: channel-based fullresolution contour refinement feedback for patch-aware gland image segmentation. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07529-8.

Ki:2025:CSM

[763] Cibin Ki, Ramah Sivakumar, and Tony Jan. A comprehensive survey of machine learning and deep learning approaches for anomaly detection in high-performance computing systems. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07503-4.

Roh:2025:CEQ

[764] Emily Jimin Roh, Chaemoon Im, and Soohyun Park. Computation-efficient quantum convolutional neural networks for autonomous driving applications. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07507-0.

Cui:2025:DND

[765] Yekai Cui, Peng Duan, and Jinjiang Li. DFGIC-Net: diffusion featureguided information complementary network for infrared and visible light fusion. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07476-4.

Ajirlou:2025:EEC

[766] Mohammad Farrokhzadeh Ajirlou, Abbas Jalilvand, and Madjid Khalilian. ECDOAGM: an elephant collective defense optimization algorithm with a gray model for managing distributed massive data streams in IoT architectures. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07546-7.

Lai:2025:DDW

[767] Yuanda Lai, Husheng Wu, and Shi Cheng. A distance determination wolf pack algorithm for solving highdimensional complex functions and its application. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07306-7.

He:2025:CPC

[768] Xinfeng He and Ruifeng Miao. CNSD-Par: container network isolation scheme based on behavioral similarity. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07329-0.

Zhang:2025:OBP

[769] Jie Zhang and Zhijun Zhao. Oppositionbased particle swarm optimization algorithm integrating neighborhood rough set for feature selection in heterogeneous information system. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07324-5.

Safi-Esfahani:2025:DSI

[770] Faramarz Safi-Esfahani, Habib Larian, and Seyedali Mirjalili. Dynamic scheduling of independent tasks in cloud computing environment applying improved chicken swarm optimization and differential evolution. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07206-w.

Chakravarty:2025:DIN

[771] Sourajit Chakravarty, Soham Satpati, and Arun Kumar Majumdar. DLC-IDS: a novel democratic leadership classification model for intrusion detection systems. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07336-1.

Barba-Toscano:2025:CNO

[772] Oscar Barba-Toscano, Erik Cuevas, and Miguel Islas-Toski. Cellular neighbors optimizer: a novel metaheuristic approach inspired by the cellular automata and agent-based modeling for global optimization. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07334-3.

Bhattacharjya:2025:FLB

[773] Kamalika Bhattacharjya and Debashis De. Federated learning-based privacypreserving Internet of underwater things: a vision, architecture, computing, taxonomy, and future directions. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07322-7.

Chen:2025:TSPb

[774] Tian Chen, Yu an Tan, and Ning Shi. Toward secure program execution in multi-tenant cloud FPGA environments. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07340-5.

Heng:2025:FEB

[775] Zhang Heng, Chen Jinlian, and Liu Yanli. Feature extraction via B-spline pyramids and adaptive QuadTree optimization. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07346-z.

Shahzadi:2025:FOC

[776] Maryam Shahzadi, Muhammad Zeeshan Khan, and Muhammad Arshad Shehzad Hassan. Fusion of optimal cross-stage partial net with context-aware segmentation for enhanced damage detection in civil infrastructures using thermal images. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07349-w.

Inik:2025:SEM

[777] Özkan Inik. SwarmCNN: An efficient method for CNN hyperparameter optimization using PSO and ABC metaheuristic algorithms. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07347-y.

Zhao:2025:ETS

[778] Qinghe Zhao, Fang Lai, and Xingye Mo. Exploring time series analysis in frequency domain with complexvalued spectral attention and bidirectional variable mamba. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07277-9.

Wang:2025:AAN

[779] Lei Wang, Ning Zhao, and Jinran Li. Application of artificial neural networks for predicting friction coefficient and film thickness considering non-Newtonian and thermal effects. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07367-8.

Touqir:2025:SEF

[780] Asma Touqir, Faisal Iradat, and Olivier Haas. Systematic exploration of fuzzing in IoT: techniques, vulnerabilities, and open challenges. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07371-y.

Qian:2025:ENE

[781] Miao Qian, Yatong Yu, and Xudong Hu. EGAL-Net: an efficient gradient aggregation lightweight network for fabric defect detection. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07355-y.

Shah:2025:TLD

[782] Syed Shakir Hameed Shah, Norziana Jamil, and Atta ur Rehman Khan.

Transfer learning with dual-stage discrete wavelet transform for enhanced visual malware image compression and classification. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07358-9.

Liu:2025:SDS

[783] Yanan Liu, Fei Yan, and Yunqing Liu. SMSERT-DETR: spatial multi-scale efficient real-time detector for maize disease detection. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07353-0.

Zhang:2025:CLC

[784] Qingling Zhang, Peng Wang, and Haoyuan Shan. CoMP-LG: coordinateaware multi-agent pathfinding via learnable communication graphs. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07377-6.

Zhang:2025:LCB

[785] Zhe Zhang, Yuming Chen, and Jie Song. Learning code better through structural information of data flow. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07348-x.

Chen:2025:GEI

[786] Weiqiang Chen, Xusheng Lin, and Dong Yu. GAPPO-ERW: an imitation reinforcement learning approach for AGV path planning. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07361-0.

Li:2025:TOC

[787] JunFei Li, QingTao Zeng, and Wen-Jing Li. Task offloading and computing resource allocation of the joint UAV with computing power network. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07351-2.

Li:2025:LMS

[788] Mingchun Li, Yang Liu, and Xin Li. Lightweight metal surface defect segmentation method based on multi-scale feature fusion and knowledge distillation. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07397-2.

Hornik:2025:LRT

[789] Jacob Hornik, Matti Rachamim, and Chezy Ofir. Leveraging real-time digital twins for smart livestreaming platforms to enhance consumers’ experience. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07386-5.

Wu:2025:GDL

[790] Peng Wu, Tao Yao, and Linliang Zhang. GA-DETR: a lightweight and efficient infrared target detection method for UAVs. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07247-1.

Ji:2025:MDF

[791] Lixia Ji, Jingmei Feng, and Han Zhang. Multi-dimensional feature collaborative fusion networks for multimodal fake news detection. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07365-w.

Naz:2025:EFD

[792] Sidra Naz, Muhammad Hassan Khan, and Muhammad Shahid Farid. Enhancing fall detection using multimodal time-series sensory data and VLAD encoding: a framework. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07370-z.

Nejad:2025:SRR

[793] Fatemeh Vafaei Nejad, Mohammad Mehdi Gilanian Sadeghi, and Mohammad Hossein Rezvani. Secure routing in RPL-based IoT networks considering behavioral trust and energy requirements. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07299-3.

Nie:2025:SEF

[794] Tingyuan Nie, Jingjing Nie, and Kun Zhao. Stacking ensemble featureattention LOFs for detecting hardware Trojan. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07357-w.

Liu:2025:CSA

[795] Siyuan Liu, Hongkun Zhao, and Kang Li. CU-SEMLP: All-MLP-based multimodal interaction model for multimodal sentiment analysis. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07364-x.

Ren:2025:PPB

[796] Chao Ren, Jiayin Song, and Yupei Li. Periodic prediction-based integrated solutions for wireless communication and edge computing in smart railway systems. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07399-0.

Tiwari:2025:HDM

[797] Kajal Tiwari and Sanjay Kumar. A healthcare data management system: blockchain-enabled IPFS providing algorithmic solutions for increased privacy-preserving scalability and interoperability. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07400-
                                        w.

Wang:2025:SGS

[798] Zhanchi Wang, Yi Hu, and Jianrong Tan. St-Graphormer: spatio-temporal graph transformer for end-to-end traffic forecasting. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07354-z.

Liu:2025:SSS

[799] Hanyu Liu and Kezhen Zhong. SS-MIM: Semi-supervised Multimodal Interaction Model for real-time conversational emotion recognition. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07373-w.

Zhao:2025:PPS

[800] Yuqi Zhao, Zhiming Song, and Rong Jiang. A privacy-preserving and secure editable blockchain model for government data sharing. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07388-3.

Moghani:2025:EHS

[801] Samira Moghani, Hossein Marvi, and Zeynab Mohammadpoory. Enhanced heart sound analysis through hierarchical spectral basis vector extraction using deep orthogonal nonnegative matrix factorization. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07350-3.

Huang:2025:SRA

[802] Yonglin Huang, Yi Zheng, and Lei Zhou. Stepwise regression algorithm based on the ordered pair of normalized real numbers framework. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07369-6.

Saif:2025:DHI

[803] Amal Saif, Omar Alhuniti, and Basel A. Mahafzah. Diamond hypercube interconnection network: topological structure and properties. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07379-4.

Lv:2025:CDA

[804] Jialun Lv, Qimeng Yang, and Long Yu. CLIP-driven attention network for multimodal sentiment analysis. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07411-7.

Gu:2025:SCS

[805] Haoxin Gu, Ruiping Xiong, and Yu Tian. SF-CLIP3D: spatialfrequency-enhanced vision-language models for multi-modal 3D object detection. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07384-7.

Chen:2025:PTD

[806] Shiwei Chen, Kaiyuan Wang, and Ensheng Chen. Practical-time decompositions against three kinds of S-boxes with different structures. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07375-8.

Liu:2025:DLT

[807] Gaofeng Liu. DARVFL-LSTM: a time series prediction model integrating dynamic regularization and attention mechanism. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07360-1.

Yi:2025:MSS

[808] Lingzhi Yi, Biao Chen, and Tao Sun. Multi-scale spatial-temporal interactive network for wind turbine ultra-short-term power prediction. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07376-7.

Zheng:2025:MME

[809] Zijun Zheng, Heng Wu, and Gaohang Yu. MERGE: multimodal-enhanced representation and guided ensemble for pneumonia recognition in chest X-ray images. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07409-1.

Wang:2025:LCR

[810] Jingwen Wang, Peng Hu, and Zhihong Wang. Low-cost ReLU search space compression, replacement, and recovery for efficient private network inference. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07382-9.

Qi:2025:VQN

[811] Han Qi, Long Zhang, and Abdullah Gani. Variational quantum neural networks based on dynamic layerwise strategy. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07394-5.

Basiri:2025:EVI

[812] M. Mohamed Asan Basiri. Efficient VLSI implementations of singular value decomposition. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07420-6.

Du:2025:SDS

[813] Nana Du, Yudong Ji, and Weike Nie. Scheduling DAG-structured workloads based on whale optimization algorithm. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07415-3.

Lu:2025:EKG

[814] Chunyu Lu, Tianran Chen, and Ruhui Shi. Enhancing knowledge graph density through graph relation attention and contrastive learning. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07424-2.

Zhang:2025:UFE

[815] Yu Zhang, Daoqi Huang, and Huajun Zhang. A universal framework for enhancing imperceptibility in latent diffusion-based adversarial attacks. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07433-1.

Zhan:2025:CBB

[816] De-Pei Zhan, Jian-Ping Yu, and Hang Yao. A consortium blockchain-based sealed electronic auction scheme utilizing attribute-based hybrid encryption. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07436-y.

Wang:2025:MVP

[817] Huiyong Wang, Jielian Feng, and Changsong Yang. MVFL: verifiable privacy-preserving federated learning using multi-key homomorphic encryption. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07419-z.

Qaffas:2025:ADD

[818] Alaa A. Qaffas. AI-driven distributed IoT communication architecture for smart city traffic optimization. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07426-0.

Biswas:2025:IHM

[819] Saptadeep Biswas, Binanda Maiti, and Uttam Kumar Bera. An innovative hybrid method combining grey wolf and marine predator optimization techniques for global and constrained problem-solving. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07410-8.

Wang:2025:OHS

[820] Lingfei Wang, Maria A. Rodriguez, and Nir Lipovetzky. Optimizing HPC scheduling: a hierarchical reinforcement learning approach for intelligent job selection and allocation. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07396-3.

Fei:2025:HPS

[821] Zhengshun Fei, Xiangyu Qin, and Xinjian Xiang. Hybrid personalized sequence recommendation based on LSTM and filter enhancement.

The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07385-6.

Rajabi-Alni:2025:EAL

[822] Fatemeh Rajabi-Alni, Behrouz Minaei-Bidgoli, and Alireza Bagheri. An efficient algorithm for the limited-capacity many-to-many point matching in one dimension. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07408-2.

Lekshmi:2025:BTH

[823] C. R. Lekshmi and Rajeev Rajan. Beyond transformers: hierarchical contextualization and gated aggregation for multiple predominant instrument recognition in polyphonic music. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07279-7.

Yang:2025:KHC

[824] Bingqing Yang, Maoli Wang, and Jianlei Liu. Kacformer: a hybrid CNN-transformer framework for image dehazing via knowledge transfer. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07429-x.

Jha:2025:ECO

[825] Mukesh Kumar Jha and Mohit Kumar. An efficient computation offloading and resource allocation technique using deep reinforcement learning in edge-cloud collaborated environment. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07366-9.

Chang:2025:PLP

[826] Chia-Ling Chang, Yen-Liang Chen, and Li-Ting Lin. PIFTA4Rec: leveraging personalized item frequency and temporal attention for enhanced next-basket recommendation. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07417-1.

Liu:2025:MRL

[827] Yanpei Liu, Yanqiang He, and Chuang Han. Meta-reinforcement learningbased task offloading method for UAV-enabled mobile edge computing. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07437-x.

Kim:2025:ODB

[828] Jin Hyuk Kim and Jonghwan Choi. OCSAug: diffusion-based optical chemical structure data augmentation for improved hand-drawn chemical structure image recognition. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07406-4.

Huang:2025:HPW

[829] Cheng Huang, Shaojuan Ma, and Xinyi Xu. The hybrid prediction of wind turbine gearbox oil temperature based on chaotic theory. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07438-w.

Teng:2025:SRS

[830] Xiaoxi Teng, Wendong Zhang, and Mengnan Ma. Stff-rtdetr: a small object detection algorithm based on drone aerial photography. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07362-z.

Yuan:2025:EGA

[831] Jiangtao Yuan, Jing Yang, and Shijie Zhou. Efficient group authentication with multiple authentications on resource-limited devices. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07404-6.

Ren:2025:SLT

[832] Xiangyu Ren, Jian Chen, and Fanlong Zhang. SVDFormer: lightweight transformer based on singular value decomposition. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07430-4.

Gao:2025:RDS

[833] Jie Gao, Xiaohong Zhang, and Shaojiang Zhong. Research on data storage and sharing of intelligent video surveillance in steel plants based on blockchain. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07405-5.

Mimi:2025:DFR

[834] Mousumi Ahmed Mimi, Hu Ng, and Timothy Tzen Vun Yap. Directional features and rule-based labeling for real-time network trafficbased Android spyware classification. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07391-8.

Ren:2025:TTG

[835] Chunni Ren, Jinsong Wang, and Zening Zhao. TGAC: traffic graph adaptive convolutional neural networkbased decentralized application encrypted traffic classification. The Journal of Supercomputing, 81(8):

??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07403-7.

Jia:2025:DCN

[836] Yihan Jia, Shengnan Hao, and Ivan Ganchev. DenLsNet-C: a novel model for breast cancer classification in pathology images based on DenseNet and LSTM. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07383-8.

Irvanizam:2025:MCD

[837] Irvanizam Irvanizam, Novi Zahara, and Herry Junaidi. A multicriteria decision-making framework using spherical trapezoidal fuzzy harmonic mean aggregation and the bestworst method for selecting an electric vehicle charging station location. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07372-x.

Zhao:2025:MOT

[838] Runxing Zhao, Jianxin Gong, and Zhiwen Wang. Multi-object tracking in surveillance scenarios based on joint detection and multi-feature fusion. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07381-w.

Langaroudi:2025:CRT

[839] Amir Hossein Langaroudi and Mohammad Abdollahi Azgomi. A collusion-resistant trust management system in fog computing. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07402-8.

Gu:2025:FGF

[840] Hangfan Gu, Jue Wang, and Jinzhang Li. FDAG-GAN: frequency-domain attention-guided GAN with feature restoration for underwater image enhancement. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07407-3. See cor-

rection [934].

Hao:2025:EGC

[841] Yuqi Hao, Yu Sun, and Xiaodong Yu. Enhanced graph convolutional and attention fusion network for hyperspectral image classification. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07445-x.

Sayed:2025:HSC

[842] Ratshih Sayed, H. H. Draz, and Haytham Azmi. HW-SW co-design of image classification accelerator on FPGA. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07380-x.

Noh:2025:MPT

[843] Kangjun Noh, Hyeji Hwang, and Kyungwoo Song. Multi-purpose technology commercialization recommender system with large-scale Korean language model. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07341-4.

Zhou:2025:HED

[844] Xiaoyan Zhou, Yu Lin, and Xiaojian Zheng. A highly efficient distributed database knobs tuning method based on multi-agent deep reinforcement learning. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07416-2.

Farshadinia:2025:SEE

[845] Hojjat Farshadinia, Ali Barati, and Hamid Barati. A secure and energyefficient architecture in Internet of things-cloud computing network by enhancing and combining three cryptographic techniques via defining new features, areas, and entities. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07390-9.

Chuang:2025:MEP

[846] Po-Jen Chuang and Jung-Chun Yao. MLTree: an efficient packet classification algorithm using multiple layered trees in software defined networks. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07425-1.

Deng:2025:AAD

[847] Zhiyong Deng, Ronggui Wang, and Juan Yang. Adpl: attentive dual-modality prompt learning for vision-language understanding. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07392-7.

Fan:2025:DAD

[848] Yifan Fan, Hao Ma, and Ben Wang. A DDoS attack detection method based on improved transformer and temporal feature enhancement. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07440-2.

Xi:2025:DCC

[849] Zhengjie Xi, Jin Liu, and Jing Liu. Dual-constraint capsule network for visible-infrared cross-modality pedestrian re-identification. The Journal of Supercomputing, 81(8):

??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07443-z.

Hu:2025:PSL

[850] Honglang Hu, Chingfang Hsu, and Long Li. Provably secure and lightweight authentication protocol using PUF and blockchain for smart grids. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07422-4.

Ortakci:2025:OSL

[851] Yasin Ortakci and Burak Borhan. Optimizing SBERT for long text clustering: two novel approaches with empirical insights. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07414-4.

Han:2025:MRM

[852] Zaikun Han, Siwen Wei, and Kexin Wang. MPD-RL: Meta-path-driven reinforcement learning for enhancing resilience in unmanned weapon systemof-systems. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07435-z.

Seyedkolaei:2025:DSP

[853] Ali Abdi Seyedkolaei. Dynamic sink placement in wireless sensor networks for enhancing network lifetime and improving network survivability. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07395-4.

Su:2025:CYO

[854] Bing Su, Jie Zhang, and Yifeng Lin. CD-YOLOv8s: an optimized high-altitude real-time UAV recognition method based on image detection. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07446-w.

Liu:2025:ATD

[855] Kaixin Liu, Guofu Feng, and Ming Chen. ACIMNet: a text detection method for chip package drawings based on dilated channel interaction and multidimensional scale fusion. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07450-0.

Zhou:2025:DSO

[856] Huaping Zhou, Xu Cao, and Bin Deng. Dense small object detection via multi-scale fusion and context information enhancement. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07444-y.

Liu:2025:DNM

[857] Yaodi Liu, Kun Zhang, and Rong Tong. A discontinuous NER model based on token prediction and contrastive learning to enhance span. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07442-0.

Zhang:2025:TRS

[858] Ping Zhang, Jingfeng Jie, and Zhi Liu. Turbulence-resistant secure transmission scheme for color images in FSOC based on hyperchaotic encryption and DNA encoding. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07439-9.

Nie:2025:DBI

[859] Shiqiang Nie, Yang Guo, and Weiguo Wu. DPUSwap: building an infinite swap with DPU for cloud computing system. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07421-5.

Song:2025:ASS

[860] Chunli Song, Linzhang Lu, and Chengbin Zeng. Adaptive smoothed successive projection algorithm for data representation. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07423-3.

Li:2025:PPM

[861] Zhengyan Li, Linxi Li, and Jie Hu. Pmsafe: parallel multi-scale attention fusion encoder for medical image segmentation. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07462-w.

Joshi:2025:EEM

[862] Divya Joshi and Vivekanand Jha. Energy-efficient multi-static sink model for wireless sensor network. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07418-0.

Hussain:2025:SSC

[863] Akram Hussain and Muhammad Fahad. Semi-supervised clustered federated learning based indoor localization with non-IID data. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07463-9.

Meng:2025:MMF

[864] Xin Meng, Yuan Zhou, and Shifeng Wang. MFTFF: multi-frame 3D object detection with temporal feature fusion. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07469-3.

Yang:2025:BCO

[865] Aiqiang Yang, Jie Liu, and Keqin Li. Balancing communication overhead and accuracy in compression integration: a survey. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07451-z.

Yu:2025:ABP

[866] Daoming Yu, Shaowen Wang, and Jiaxin Zou. Adaptive bidirectional planning framework for enhanced safety and robust decision-making in autonomous navigation systems. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07389-2.

Ling:2025:FYE

[867] Hanhuang Ling, Ling Li, and Yuxiang Zhang. FDL-YOLO: efficient real-time traffic detection network based on frequency domain learning. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07398-1.

Feng:2025:USA

[868] Zhiwei Feng, Benyi Yang, and Baosong Deng. UTS-SAM: adapting segment anything model for UAV target segmentation based on a complementary dual encoder. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07427-z.

Xu:2025:END

[869] Shoukun Xu, Yihang Shi, and Huajun Zhang. Efficient network defense policies via GNN-enhanced reinforcement learning. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07431-3.

Yao:2025:EEC

[870] Yifan Yao, Juan Zhang, and Yongbin Gao. ECKT: enhancing crosstask knowledge transfer in continual few-shot relation extraction. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07475-5.

Jain:2025:DRM

[871] Richa Jain and Ashok Kumar. Development of a robust model for identifying vulnerabilities in IoT devices via binary and multiclass deep neural network classification. The Journal of Supercomputing, 81(8):

??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07458-6.

Liu:2025:SMA

[872] Xianchang Liu, Peng Wang, and Haoyuan Shan. Scalable multiagent path planning via wavelet-based KAN and enhanced feature extraction. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07470-w.

Chen:2025:TSS

[873] Hongwei Chen, Shikai Kou, and Liya Xia. Temporal-spatial skeleton sequence recognition with selfsupervised representation learning. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07479-1.

Nagarajan:2025:DKR

[874] Susila Nagarajan, Shobana Mahalingam, and Esther Daniel. Deep Kronecker ResNeXt forward fractional network-driven clustering for enhanced load balancing in cloud computing. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07359-8.

Zhang:2025:DGC

[875] Anqin Zhang, Pengzhou Chen, and Ting Zhang. Dynamic graph contrastive learning for multivariate time series anomaly detection. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07455-9.

Feng:2025:TMA

[876] Cheng Feng, Hai Yang, and Xue Li. Triaxial modality attention fusion with top-down mask generation for enhanced multimodal sentiment analysis. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07474-6.

Chaudhary:2025:ACF

[877] Rajesh Kumar Chaudhary, Ravinder Kumar, and Nitin Saxena. Ameliorating clustered federated learning using grey wolf optimization algorithm for healthcare IoT devices. The Journal of Supercomputing, 81 (8):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07374-9.

Zhang:2025:OCS

[878] Xiaodan Zhang, Lu Zhang, and Yichong She. Optimal channels selection based on ABC-SVM in Parkinson’s disease detection using short-time resting state EEG. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07461-x.

Zhao:2025:DSS

[879] Hui Zhao and Lei Zhang. Dualstream siamese network for RGB-T dual-modal fusion object tracking on UAV. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07519-w.

Kalson:2025:NRL

[880] Shweta Kalson and Anjana Gupta. A novel reciprocal linguistic cloud model and its integration with machine learning. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07428-y.

Xia:2025:EUF

[881] Ze Xia, Dinghan Hu, and Jiuwen Cao. EAViz: a user-friendly deep learning-based epilepsy analysis visualizer using multimodal data. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07494-2.

Huang:2025:APL

[882] Juan Huang, Yanhua Zhong, and Xian Chen. Adaptive and personalized learning in STEM education using highperformance computing and artificial intelligence. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07481-7.

Fang:2025:ATD

[883] Jie Fang, Nian Guo, and Yang Xiang. Application of three-dimensional Rubik’s cube scrambling and genetic information diffusion in multiplecolor image encryption. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07363-
y.

Yang:2025:MSI

[884] Tengfei Yang and Hao Liu. A multi-strategy improved exponential distribution optimizer for numerical optimizations and engineering problems. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07452-y.

Chen:2025:FRE

[885] Xi Chen and Guohui Wang. FP-RTDETR: enhancing infrared ship detection with multi-scale feature fusion and lightweight design. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07485-3.

Fan:2025:SME

[886] Haoran Fan, Bin Li, and Shoujun Zhou. Small but mighty: enhancing time series forecasting with lightweight LLMs. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07491-5.

Ferrer:2025:PCC

[887] Joaquín Ferrer, Juan M. Cebrian, and Manuel E. Acacio. Precise characterization of coherence activity in multicores using gem5. The Journal of Supercomputing, 81(8): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07434-0.

Li:2025:SGU

[888] Jinzhang Li, Jue Wang, and Yuan Li. SER-GAN: underwater image enhancement via spatially enhanced residual blocks and generative adversarial network. The Journal of Supercomputing, 81(8):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07356-x.

Zhu:2025:RMD

[889] Weizhi Zhu, Lianfang Tian, and Juanhong Xie. The 3D reconstruction method for driving scenes based on improved neural radiance fields.

The Journal of Supercomputing, 81 (9):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07538-7.

Alyoubi:2025:IME

[890] Adel A. Alyoubi and Bader A. Alyoubi. Interpretable multimodal emotion recognition using optimized transformer model with SHAP-based transparency. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07515-0.

Ktari:2025:IFT

[891] Mouna Ktari, Raïda Ktari, and Ahmed Hadj Kacem. iAgent fault-tolerance approach (iAFTA) based on optimization algorithms in wireless sensor networks. The Journal of Supercomputing, 81(9): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07525-y.

Selvam:2025:FBS

[892] Immaculate Joy Selvam, Srinivasan Chandran, and Saravanakumar Chandrasekaran. FPGA-based SqueezeNet transformer with multi-scale attention fusion for real-time respiration rate estimation using PPG signals. The Journal of Supercomputing, 81 (9):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07543-w.

Alba:2025:PPG

[893] Vincent Alba, Olivier Aumage, and Amina Guermouche. Performance portability of generated cardiac simulation kernels through automatic dimensioning and load balancing on heterogeneous nodes. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07510-5.

Asadianfam:2025:ASO

[894] Shiva Asadianfam and Hojjat Emami. An adaptive seasons optimization algorithm for global optimization. The Journal of Supercomputing, 81 (9):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07511-4.

Guo:2025:MLL

[895] Wenjie Guo, Jingfeng Xue, and Jingjing Hu. Malgta: large language model-based guided malware tactical analysis. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07545-8.

Mao:2025:CEF

[896] Aonan Mao, Di Wu, and Wushour Slamu. A classifier expansion framework with dual knowledge distillation and dynamic weighting for continual relation extraction. The Journal of Supercomputing, 81(9): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07488-0.

Tian:2025:AMO

[897] Huixian Tian and Yingjie Song. An augmented multi-objective multi-verse optimizer algorithm for solving dynamic economic emission dispatch problems. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07527-w.

Zhang:2025:EDE

[898] Chaofeng Zhang and Jingmin Yang. Emsd-detr:efficient small object detection for UAV aerial images based on enhanced RT-DETR model. The Journal of Supercomputing, 81(9): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07524-z.

Sun:2025:LWE

[899] Yuan Sun, Lahuan Li, and Binyao Ma. Locally weighted ensemble clustering based on grain distance. The Journal of Supercomputing, 81 (9):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07548-5.

Wang:2025:PCF

[900] Yi Wang and Dongqin Cheng. The 3-path connectivity of the folded hypercube. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07558-3.

Wang:2025:HMT

[901] Hailin Wang, Song Zhang, and Ao Ma. Harnessing multi-task learning to improve overlapping relation extraction. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07533-y.

Zhang:2025:QDE

[902] Hongke Zhang and Zheng Yan. QuickEn: a data encryption device for high performance storage. The Journal of Supercomputing, 81 (9):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07532-z.

Simsek:2025:FSL

[903] Hakan Simsek and Ömer Faruk Öncel. Fuzzy security level metric of hybrid cryptosystem algorithms. The Journal of Supercomputing, 81 (9):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07547-6.

Lu:2025:OIS

[904] Shizhao Lu, Tanny Chavez, and Alexander Hexemer. Optimizing inference of segmentation on highresolution images in MLExchange. The Journal of Supercomputing, 81 (9):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07413-5.

Dwivedi:2025:BEE

[905] Abhishek Dwivedi, Ratish Agarwal, and Piyush Kumar Shukla. A blockchain-enabled encrypted neural network framework for trust-aware key management and node authentication in Industrial Internet of Things. The Journal of Supercomputing, 81 (9):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07566-3.

Cai:2025:EPS

[906] Guangying Cai, Liujia Cai, and Haoyuan Xue. Enhancing privacy and security in IoT: a CoAP protocol analysis and improvement approach. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07401-9.

Lei:2025:PDE

[907] Huan Lei, Lingfei Ren, and Wenyuan Yang. PAF-DETR: enhancing UAV image detection with partial attention and dynamic feature integration. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07561-8.

Shi:2025:FSD

[908] Kai Shi, Penghao Ding, and Jinsong Wang. A few-shot detection method for new types of network traffic attacks based on meta-learning with crossattention. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07534-x.

Cao:2025:PRU

[909] Jianhao Cao, Hao Cai, and Ning Xu. PCB routing on unstructured meshes with conflict-based search. The Journal of Supercomputing, 81 (9):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07569-0.

Sood:2025:IPQ

[910] Sandeep Kumar Sood, Manmohan Singh, and Munish Bhatia. Industrial progress with quantum algorithms: an in-depth review. The Journal of Supercomputing, 81(9): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07555-6.

Siyuan:2025:APM

[911] Shang Siyuan, Liu Aodi, and Tan Ming. ABAC policy mining method for heterogeneous access control system. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07539-6.

Dehghani:2025:ENB

[912] Abbas Dehghani. An efficient NoCbased architecture for convolutional neural networks on multicore systems. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07526-x.

Huang:2025:MPF

[913] Jinjie Huang and Wenhe Li. Multipath feature fusion and constrained boundary optimization for open intent detection. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07499-x.

Shui:2025:TIP

[914] Xiangyu Shui, Zhenfang Zhu, and Huaxiang Zhang. Text-to-image person retrieval with implicit relation alignment and contrastive learning. The Journal of Supercomputing, 81 (9):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07568-1.

Nikolic:2025:FAD

[915] Sanja Nikolic, Luka Filipovic, and Milica Vukotic. FIT4HPC? — accelerating digital transformation by supercomputing opportunities. The Journal of Supercomputing, 81(9): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07559-2.

Guitouni:2025:QCT

[916] Zied Guitouni and Hafedh Mahmoud Zayani. Quantum chaotic techniques for medical image encryption in next-generation IoMT applications. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07574-3.

Park:2025:SST

[917] Soohyun Park, Jae Hyun Cho, and Joongheon Kim. SQUAD: software testing for quantum distributed learning software. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07556-5.

Saeedi:2025:PEE

[918] Nima Saeedi, Deniz Baharvand, and Sina Samadi Gharehveran. Prediction of electrical energy consumption using principal component analysis and independent components analysis. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07505-2.

Tan:2025:DSS

[919] Jiawei Tan, Li Yang, and Xiong Xiao. A distributed skewed stream processing system based on scoring high-frequency key perception. The Journal of Supercomputing, 81 (9):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07465-7.

Yang:2025:AAF

[920] Ruidi Yang, Yuxing Mao, and Jianyu Pan. AFMD: attention fusionbased multiscale decomposition network. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07565-4.

Albalawi:2025:DCO

[921] Nasser Albalawi. Dynamic computational offloading approaches for IoT devices in cloud computing. The Journal of Supercomputing, 81 (9):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07551-w.

Verma:2025:CSS

[922] Rahul Kumar Verma, Shubhra Jain, and Abhinesh Kaushik. A comparative study and survey of chain-based routing protocols in wireless sensor networks. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07412-6.

Kumar:2025:ORA

[923] Bablu Kumar, Anshul Verma, and Akram Bennour. Optimizing resource allocation in cloud-native applications through proactive autoscaling with the InformerAutoScale model. The Journal of Supercomputing, 81 (9):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07500-7.

Kazemi:2025:EAS

[924] Ferdos Kazemi, Behnam Barzegar, and Meisam Yadollahzadeh-Tabari. An energy-aware scheduling in DVFS-enabled heterogeneous edge computing environments. The Journal of Supercomputing, 81(9): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07432-2.

Zhang:2025:VLD

[925] Jianwei Zhang, Xinyu Yan, and Bineng Zhong. Vision-language discriminative fusion network for object tracking. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07472-8.

Chu:2025:WNM

[926] BoHua Chu, BaoJu Zhang, and CuiPing Zhang. WTDL-Net: medical image registration based on wavelet transform and multi-scale deep learning. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07567-2.

Zheng:2025:ITB

[927] Fang Zheng, Juanjuan Zhao, and Yifang Zheng. An interpretable twobranch network based on an prototype to promote group gene biomarkers discovery. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07554-7.

Zhou:2025:MEL

[928] Chengjie Zhou, Ke Jiang, and Qiang Zhang. MSTF: enhancing long-term forecasting with multi-scale temporal fusion in time series forecasting. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07572-5.

Rezapourshahir:2025:NLB

[929] Samira Rezapourshahir and Shahram Babaie. A new load-balancing approach for cloud computing based on an improved genetic algorithm. The Journal of Supercomputing, 81 (9):??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07368-7.

Alam:2025:ASF

[930] Khairul Habib Alam, Yumnam Rohen, and Anita Tomar. Approximating the solutions of fractional differential equations with a novel and more efficient iteration procedure. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07562-7.

Rasheed:2025:YOE

[931] Areeg Fahad Rasheed and Mahdi Zarkoosh. YOLOv11 optimization for efficient resource utilization. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07520-3.

Park:2025:QRS

[932] Jong Hyuk Park and Minji Kim. Quantum-resilient security for 6G networks: a comprehensive survey on challenges, solutions, and research opportunities. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07544-9. See cor-

rection [958].

Mujlid:2025:QDS

[933] Hana Mohammed Mujlid and Reem Alshahrani. Quantum-driven security evolution in IoT: AI-powered cryptography and anomaly detection. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07582-3.

Gu:2025:CFG

[934] Hangfan Gu, Jue Wang, and Jinzhang Li. Correction: FDAG-GAN: frequency-domain attention-guided GAN with feature restoration for underwater image enhancement. The Journal of Supercomputing, 81(9): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07557-4. See [840].

Song:2025:IAM

[935] Yingjie Song and Xidong Liu. An improved adaptive multi-tasking differential evolutionary algorithm with opposition-based learning. The Journal of Supercomputing, 81(9): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07535-w.

Sun:2025:UYN

[936] Guang-Ling Sun, Fen-Qi Zhang, and Fei Miao. UCS-YOLO: a novel approach for multi-scale small target detection in UAVs. The Journal of Supercomputing, 81(9): ??, June 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07536-9.

Jin:2025:GNS

[937] Tianchun Jin, Jindong Zhang, and Ma Mingzhou. GCE-net: single-image deraining with adaptive normalization and multi-scale feature integration. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07603-1.

Zhang:2025:ETK

[938] Yue Zhang, Guodong Ma, and Fuqiang Yang. Enhancing temporal knowledge graph completion with counterfactual insights. The Journal of Supercomputing, 81(9):??, June 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07579-y.

Akhtar:2025:DAB

[939] Muhammad Waseem Akhtar, Zia Bashir, and M. G. Abbas Malik. Dynamical analysis and bifurcation exploration in a biological model with neural network forecasting. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07599-8.

Kulasekara:2025:SMM

[940] Maithri Kulasekara, Juan F. Inglés-Romero, and José L. Abellán. Saffe: Multimodal model composition with semantic-alignment fusion of frozen encoders. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07473-7.

Zhang:2025:PMO

[941] Mandun Zhang, Chenyue Hou, and Xiangsheng Huang. Preaftrack: multiobject tracking based on adaptive feature matching from detection results. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07466-6.

Zaizi:2025:DAE

[942] Fatima Ezzahra Zaizi, Sara Qassimi, and Said Rakrak. A deep autoencoderenhanced multi-objective evolutionary algorithm for recommender systems. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07589-w.

Xu:2025:DAS

[943] Fei Xu and LinXuan Ye. A domain-adaptive sentiment analysis method integrating attentional mechanisms and adversarial training. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07608-w.

Salehi:2025:EPC

[944] Mojtaba Salehi, Saeed Parsa, and Hoshang Kolivand. Efficient path coverage-based test data generation using an enhanced pelican algorithm. The Journal of Supercomputing, 81 (10):??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07611-1.

Ornek:2025:AFD

[945] Bülent Nafi Örnek, Canan Oral, and Süleyman Dirik. Activation function design based on Schwarz lemma for neural networks. The Journal of Supercomputing, 81(10): ??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07606-y.

Gong:2025:ERT

[946] Boao Gong, Haoran Zhang, and Zhiyong Tao. Enhancing real-time lowlight object detection via multi-scale edge and illumination-guided features in YOLOv8. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07594-z.

Lu:2025:HHG

[947] Yun-Jun Lu, Xiao-Yao Xie, and Song Liu. HGFT: hybrid geometric feature transformer for point cloud registration. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07592-1.

Wu:2025:MSA

[948] Jiawei Wu and Yong Liu. A multi-strategy adaptive cultural algorithm for scheduling high-dimensional reentrant hybrid flowshops with incompatible job families. The Journal of Supercomputing, 81(10): ??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07581-4.

Lopez-Fernandez:2025:BBU

[949] Aurelio López-Fernández, Francisco A. Gomez-Vela, and Jorge Gonzalez-Dominguez. Biclustering in bioinformatics using big data and high performance computing applications: challenges and perspectives, a review. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07563-6.

Atli:2025:BBA

[950] Zafer Atli and Bahaeddin Turkoglu. BAHA: balanced artificial hummingbird algorithm and its application to space industry. The Journal of Supercomputing, 81(10): ??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07596-x.

Yin:2025:SYS

[951] Chunfang Yin, Zihan Wang, and Yuhao Guo. SBEW-YOLOV8: a small object detection algorithm for autonomous driving based on multi-scale feature fusion. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07577-0.

Dong:2025:SMT

[952] Mingshuai Dong, Yuxuan Bai, and Xiuli Yu. A single multi-task deep neural network with a multi-scale feature aggregation mechanism for manipulation relationship reasoning in robotic grasping. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07553-8.

Dwivedi:2025:BPE

[953] Pulkit Dwivedi and Benazir Islam. Balancing performance and energy efficiency: the method for sustainable deep learning. The Journal of Supercomputing, 81(10): ??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07610-2.

Xie:2025:HBM

[954] Xintong Xie, Tinghuai Ma, and Huan Rong. Hypergraph-based multimodal adaptive fusion for emotion recognition in conversation. The Journal of Supercomputing, 81(10): ??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07584-1.

Jouybari:2025:MFF

[955] Seyedeh Maryam Hashemi Jouybari, Alireza Janbaz, and Ahmad Esfandiari. A multivariate filter feature selection and stacking-based ensemble learning algorithm for network intrusion detection. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07609-9.

Quan:2025:MHM

[956] Weiming Quan, Yin Tang, and Lei Zhang. MSA-HAR: multi-scale segmented attention networks for human activity recognition using sensor signals. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07576-1.

Iqbal:2025:HEH

[957] Asif Iqbal, Muhammad Arslan Rauf, and Zhen Qin. E-harnet: an efficient hybrid transformer network for human activity recognition. The Journal of Supercomputing, 81(10): ??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07618-8.

Kim:2025:CQR

[958] Minji Kim and Jong Hyuk Park. Correction: Quantum-resilient security for 6G networks: a comprehensive survey on challenges, solutions, and research opportunities. The Journal of Supercomputing, 81 (10):??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07651-7. See [932].

Shi:2025:BRF

[959] Yinglong Shi, Xiaoming Hu, and Hao Lin. A Byzantine-robust federated learning against adversarial-majority attacks. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07619-7.

Mohawesh:2025:NMF

[960] Rami Mohawesh, Ahmed Abdallah AlQarni, and Haythem Bany Salameh. A new multilingual framework for fake reviews detection based on a large language model. The Journal of Supercomputing, 81 (10):??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07636-6.

Huang:2025:EGR

[961] Xingbin Huang, Rui Chen, and Weihua He. Estimating graph robustness through the Kirchhoff index and sampling techniques. The Journal of Supercomputing, 81(10): ??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07578-z.

Zhao:2025:RRM

[962] Shihai Zhao, Xiaodong Fu, and Jiaman Ding. RHFL: a robust method to defend against poisoning attacks for heterogeneous hierarchical federated learning. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07616-w.

Shirani:2025:ILB

[963] Leila Shirani, Homa Movahednejad, and Marjan Mahmoudi. Improving load balancing in distributed softwaredefined networks with a bidirectional long short-term memory-based controller design. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07595-
                                        y.

Bawa:2025:ICR

[964] Jyoti Bawa, Kuljit Kaur Chahal, and Kamaljit Kaur. Improving cloud resource management: an ensemble learning approach for workload prediction. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07560-9.

Mozaffari:2025:ETA

[965] Javad Mozaffari, Mohammad Abdollahi Azgomi, and Mohammad Reza Ebrahimi Dishabi. Energy and temperature-aware routing approach for congestion control in wireless body area networks. The Journal of Supercomputing, 81(10): ??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07622-y.

Qiao:2025:TCH

[966] Jian Hua Qiao, Yu Ting Guan, and Jin Ning Zhi. Transformer-cnn hybrid network for underwater image enhancement. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07612-0.

Cui:2025:MSF

[967] JingWen Cui, ChunHong Yuan, and Dan Zhang. Multi-scale feature aggregation with hierarchical semantics and uncertainty assessment: enabling high-accuracy visual retrieval. The Journal of Supercomputing, 81 (10):??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07623-x.

Song:2025:AFB

[968] Qingjun Song, Shirong Sun, and Lina Lu. An attentional fusion-based method for coal-gangue recognition in noisy environment of generalised workface. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07597-w.

Zhang:2025:HAM

[969] Youli Zhang, Hu Zhang, and Changjian Song. A hybrid algorithm for multiobjective task scheduling in heterogeneous cloud computing. The Journal of Supercomputing, 81(10): ??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07621-z.

Hu:2025:STB

[970] Feng Hu, Wei Pei, and Shuifa Sun. Star-transformer based semantic enhanced union relation extraction. The Journal of Supercomputing, 81 (10):??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07591-2.

Hong:2025:PEP

[971] Yang Hong, Xinyue Yao, and Jianqing Mo. Pointat: enhancing point cloud contrastive representation learning via an adaptive truncated loss.

The Journal of Supercomputing, 81 (10):??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07615-x.

Alsheavi:2025:ELB

[972] Amar N. Alsheavi, Naji Alhusaini, and A. S. Ismail. Efficient load balancing in cloud computing using hybrid ant colony optimization and crow search strategies. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07550-x.

Lu:2025:KDB

[973] Shixiang Lu, Yibo Yan, and Zhi-Wei Gao. A knowledge distillation-based network compression framework for lifecycle management of lithium-ion batteries. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07638-4.

Zhou:2025:RTS

[974] Zhuguo Zhou, Yujun Lu, and Liye Lv. Real-time surface defect detection algorithm for printed circuit boards based on improved YOLOv11n. The Journal of Supercomputing, 81 (10):??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07625-9.

Cheng:2025:HNN

[975] Nuo Cheng, Anbo Tang, and Nenggang Xie. A hybrid neural network prediction model based on timefrequency domain features and its application to industrial heat exchangers. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07642-8.

Jiang:2025:EST

[976] Xian Jiang, Yiyao Zhu, and Jie Zhang. Energy-saving transmission time and power management for D2D connection with a relay node using the same band with a cellular connection. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07624-w.

Oh:2025:QIC

[977] Seungeun Oh, Jinhyuk Kim, and Seong-Lyun Kim. Quantum infidelity codistillation for fast and accurate distributed quantum machine learning. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07614-y.

Feng:2025:KRD

[978] Siling Feng, QiaoYun Wang, and Mengxing Huang. Kds-radfnet: A distributed thermal infrared and visible image fusion framework based on knowledge distillation and semantic segmentation. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07640-
w.

Leng:2025:DSP

[979] Qiangkui Leng, Chunyue Lyu, and Changzhong Wang. Dynamic selfpaced undersampling ensemble for imbalanced classification. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07633-
9.

Qu:2025:DAB

[980] Na Qu, Long Yu, and Chaoyue Wu. Dual attention-based graph convolutional neural network for multimodal sentiment analysis. The Journal of Supercomputing, 81(10): ??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07613-z.

Xue:2025:MSI

[981] Yige Xue. Multi-strategy improved snake optimizer for library robot path planning problems. The Journal of Supercomputing, 81(10): ??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07570-7.

Wang:2025:EHF

[982] Di Wang, Gangyang Nan, and Fang Xia. Enhanced human fall detection via lightweight MDS-OpenPose framework. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07628-6.

Li:2025:MSF

[983] Xiaohui Li, Taicheng Bian, and Hongtao Liang. Multi-stage federated learning with group-wise bidirectional guidance. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07449-7.

Zhang:2025:CTR

[984] Yuan Zhang, Tiantian Sheng, and Lunping Duan. Click-through rate prediction via feature selection and layer-wise residual gated cross-network. The Journal of Supercomputing, 81 (10):??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07652-6.

Luo:2025:CCV

[985] Chuan Luo. CVSR: complex-valued networks for video super-resolution. The Journal of Supercomputing, 81 (10):??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07659-z.

Yu:2025:RTS

[986] Qianqian Yu, Guang Yang, and Ang Liu. A review of time series forecasting and spatio-temporal series forecasting in deep learning. The Journal of Supercomputing, 81 (10):??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07632-w.

Yang:2025:NHF

[987] Xueguang Yang, Yuefeng Zheng, and Yang Lu. A novel hybrid feature selection algorithm based on Mann– Whitney U test and double optimization. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07629-5.

Zheng:2025:HAT

[988] Gang Zheng, Minshu Chen, and Qiong Zhou. High-accuracy tool wear monitoring for micro-milling method based on improved Pyramid Vision Transformer. The Journal of Supercomputing, 81(10):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07650-8.

Ren:2025:ISL

[989] Weiwu Ren, Bingsen Wang, and Yu Hong. Inter-satellite link allocation in low-earth-orbit mega-constellation networks. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07655-3.

Lu:2025:EBC

[990] Chundong Lu, Junna Shang, and Yanzhao Ding. The EfficientNetV2based correction localization combining CSI and RSSI. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07648-
                                        2.

Wang:2025:CTF

[991] Wentao Wang, Haiyan Lu, and Jianzhou Wang. A complementary time-frequency domain approach for short-term photovoltaic power forecasting. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07654-4.

Kansal:2025:OEI

[992] Puneet Kansal, Manoj Kumar, and O. P. Verma. Optimizing edge intelligence: a DRL-driven service migration approach with enhanced feedback in mobile edge computing. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07630-y.

Xiao:2025:FMT

[993] Bin Xiao, Zehao Ge, and Yanxue Wu. FCAFormer: multivariate time series forecasting combining channel attention and transformer in the frequency domain. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07634-8.

Shafinezhad:2025:WES

[994] Negin Shafinezhad, Hamid Abrishami, and Anahita Morvaridi. Workflow ensemble scheduling in IaaS cloud: a gap analysis perspective under deadline and budget constraints. The Journal of Supercomputing, 81 (11):??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07649-1.

Chen:2025:APT

[995] Yang Chen, Junjiang He, and Xiaolong Lan. Automatic penetration testing model based on reinforcement learning for complex network environments. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07656-2.

Xu:2025:MSI

[996] Yujie Xu, Yao Jiang, and Liyun Jia. Multi-strategy improved coati optimization algorithm. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07617-

Huang:2025:MOT

[997] Jinjie Huang and Jiayu Cui. Multiobject tracking based on spatial super-resolution and spatiotemporal convolutional self-attention. The Journal of Supercomputing, 81(11): ??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07664-2.

Cuevas:2025:FSS

[998] Erik Cuevas, Óscar A. González-Sánchez, and Daniel Zaldívar. Filling space swarm optimization (FSSO): a metaheuristic algorithm with divided agent strategies and diamond crossover. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07645-5.

Chen:2025:JRT

[999] Zhe Chen, Sihao Xing, and Yinyang Zhang. Joint relational triple extraction based on topic constraints and multicore attention fusion. The Journal of Supercomputing, 81(11): ??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07631-x.

Hagras:2025:RSA

[1000] Tarek Hagras and Gamal A. El-Sayed. A replacement for the scaling axiomatic approach to scheduling dependent tasks on DVFS computing platforms. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07637-5.

Saha:2025:IIS

[1001] Nilanjana Saha, Amrita Namtirtha, and Animesh Dutta. Identifying influential spreaders on weighted directed networks based on spreading properties. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07658-0.

Ma:2025:IAP

[1002] Gengchen Ma, Haichao Xu, and Anji Li. Intelligent audio processing methodology for operational risk reduction in complex industrial process control systems. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07666-0.

Cai:2025:SFS

[1003] Qiang Cai, Shizeng Zhang, and Qiang Cui. ST-FPN: a Swin Transformerbased lightweight model for accurate VLSI congestion prediction. The Journal of Supercomputing, 81 (11):??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07641-9.

Wang:2025:MIS

[1004] Yuxiang Wang and Xuemei Lei. A multilevel integrated supervision self-distillation method. The Journal of Supercomputing, 81(11): ??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07661-5.

Kritsadakul:2025:HCQ

[1005] Jedsadakorn Kritsadakul and Sanpawat Kantabutra. Hybrid classical quantum computation for cybersecurity strategies in a layered cybersecurity model. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07662-4.

Toscani:2025:ESM

[1006] Giulio Toscani. Exaptation as a strategic mechanism in public R&D: challenges and solutions at the Barcelona Supercomputing Center. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07588-x.

Li:2025:GKD

[1007] Xiaolong Li, Huifang Ma, and Zhixin Li. Graph knowledge distillationguided few-shot learning for hyperspectral image classification. The Journal of Supercomputing, 81(11): ??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07653-5.

Madiyev:2025:EEO

[1008] Askar Madiyev, Daulet Bulegenov, and Dinh Mao Bui. Energy-efficient offloading framework for mobile edge/ cloud computing based on convex optimization and Deep Q-Network. The Journal of Supercomputing, 81 (11):??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07647-3.

Alzati:2025:KEB

[1009] Alberto Alzati and Alfonso Tortora. Key exchange based on threedimensional Segrè products and their projections. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07657-1.

You:2025:DDD

[1010] Peng You, Xinxin Wang, and Huangyining Gao. DT-DGSL: dynamic transformer using denoising graph structure learning for IoT time series anomaly detection. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07665-1.

Liu:2025:FGS

[1011] Genggeng Liu, Xuehui Su, and Ye Tian. A fast generalized surrogate-assisted evolutionary algorithm for solving expensive constrained optimization problems. The Journal of Supercomputing, 81(11): ??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07660-6.

Alotaibi:2025:FFL

[1012] Jamal Alotaibi. FuzOptRoute: a fuzzy logic-integrated optimizationbased energy-efficient cluster routing framework with edge computing for mobile communication networks. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07673-1.

Alqahtani:2025:QCI

[1013] Abdullah Alqahtani and Munish Bhatia. Quantum computing-inspired resource distribution in healthcare. The Journal of Supercomputing, 81 (11):??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07644-6.

Lyu:2025:MGC

[1014] Yiyang Lyu, Jiazhe Li, and Weiyi Wang. Mink-GraSNet: computationally efficient 6-DoF grasp detection for large-scale AI-powered robotics. The Journal of Supercomputing, 81 (11):??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07702-z.

Khan:2025:IPK

[1015] Abdul Wahid Khan, Saad Ullah Khan, and Rahmatullah Khan. Identifying and prioritizing key success factors for scaling DevOps methods in distributed software development environment using analytical hierarchy process. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07663-3.

Hakiki:2025:EII

[1016] Racha Ikram Hakiki, Abdennour Azerine, and Lhassane Idoumghar. Enhancing IoT intrusion detection with genetic algorithm-optimized convolutional neural networks. The Journal of Supercomputing, 81(11): ??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07626-8.

Castiglione:2025:ETD

[1017] Aniello Castiglione, Jacopo Gennaro Esposito, and Matteo Polsinelli. Enhancing trust of deep learning models with post-quantum digital signatures. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07669-x.

Wu:2025:EHF

[1018] Chengmao Wu and Xiaomin Wang. An enhanced harmonic fuzzy partition clustering algorithm incorporating a novel membership modification strategy. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07678-w.

Wang:2025:AFD

[1019] Yan Wang, Jiayi Li, and Shurong Li. Attention and frequency domain enhancement for image dehazing based on transformer. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07676-y.

Alzahrani:2025:EAQ

[1020] Abdullah I. A. Alzahrani. Exploring AI and quantum computing synergies in holographic counterpart frameworks for IoT security and privacy. The Journal of Supercomputing, 81 (11):??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07682-0.

He:2025:LNM

[1021] Ao He, Guanhua An, and Jichang Guo. LMF-Net: a Mamba-based enhancement network for low-light object detection. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07675-z.

Wang:2025:HAB

[1022] Yahui Wang, Hongchang Chen, and Yuxiang Hu. Hierarchy aware-based multi-task learning for user location prediction. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07643-7.

Kumar:2025:PPP

[1023] Mandeep Kumar, Gagandeep Kaur, and Prashant Singh Rana. Performance, portability, productivity, and security in HPC cloud: a systematic literature review. The Journal of Supercomputing, 81(11): ??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07685-x.

Zhang:2025:SGL

[1024] Ao Zhang, Han Liu, and HaiQiang Yang. Spatiotemporal graph learning driven by multivariate KAN for urban road network traffic prediction. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07671-3.

Touameur:2025:GGA

[1025] Ouissem Touameur, Fouzi Harrag, and Ladjel Bellatreche. GUITARES: graph attention network for building knowledge graph-based trustaware recommender systems. The Journal of Supercomputing, 81(11): ??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07674-0.

Bai:2025:MCS

[1026] Yun Bai, Yuxuan Gong, and Feng Wei. MSF-CSCNet: a supercomputingready 3D semantic segmentation network for urban point clouds via multiscale fusion and context-aware channel modeling. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07672-2.

Ding:2025:NMW

[1027] Pengfei Ding, Weiwei Hu, and Jingge Zhu. A new multi-wing hyperchaotic system and its application in image encryption. The Journal of Supercomputing, 81(11):??, July 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07670-4.

Arsoski:2025:MCS

[1028] Vladimir V. Arsoski. Multi-controlled single-qubit unitary gates based on the quantum Fourier transform and deep decomposition. The Journal of Supercomputing, 81(11): ??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07684-y.

Umer:2025:FTM

[1029] Asif Umer, Mushtaq Ali, and Ali Imran Jehangiri. Fault tolerant and mobility-aware task offloading and scheduling model for IoT logistics. The Journal of Supercomputing, 81 (11):??, July 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07627-7.

Gohel:2025:CHP

[1030] Prashant Gohel and Manjunath Joshi.

Correction: Hybrid and parallel GAN architecture with non-IID noise input. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07703-y. See [739].

Li:2025:SMA

[1031] Zhenpeng Li, Jingwei Bian, and Lihua Zhang. Safety measures in automotive operating system: a comprehensive review of trends and defense frameworks. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07635-7.

Wang:2025:AMS

[1032] Yan Wang, Huan Wang, and Shurong Li. Adaptive multi-scale contrastaware dynamic feature fusion transformer for image dehazing method.

The Journal of Supercomputing, 81 (12):??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07686-w.

Wang:2025:MMM

[1033] Xin Wang, Xiang Zhang, and Zhenyu Zuo. MA-MSTNet: mixed attentionbased multi-scale temporal network for stock trend prediction. The Journal of Supercomputing, 81(12): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07700-1.

Zhou:2025:FYS

[1034] Huaping Zhou, Wei Yin, and Bin Deng. FO-YOLO for small object detection in drone aerial imagery. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07688-8.

Peng:2025:SYI

[1035] Ende Peng, Qing Ai, and Tao Han. SOG-YOLO: an infrared road scene small object detection model. The Journal of Supercomputing, 81(12): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07692-y.

Xue:2025:OOG

[1036] Songdong Xue, Yan Chen, and Xinyu Zhao. OFSPNet: ODE-guided fusion network with the same paradigm for infrared small target detection. The Journal of Supercomputing, 81 (12):??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07695-9.

Liu:2025:FRL

[1037] Yanpei Liu, Haoyang Zhao, and Hongchan Li. Federated reinforcement learning-based multi-UAV cooperative dynamic caching replacement strategy. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07701-0.

Zhang:2025:SAT

[1038] Mingyue Zhang and Zhiheng Zhou.

Speed-accuracy trade-off in lightweightbased hand pose estimation. The Journal of Supercomputing, 81(12): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07699-5.

Muhyi:2025:IML

[1039] Diyar Fadhil Muhyi and Oguz Ata. Integrative machine learning approaches for enhanced cardiovascular disease prediction: a comparative analysis of XGBoost and ANFIS algorithms. The Journal of Supercomputing, 81 (12):??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07687-9.

Chen:2025:IFD

[1040] Huaiguang Chen, Wenyu Wei, and Jing Gao. Image fusion and denoising of optical coherence tomography based on ResNet and truncated Huber filter. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07690-0.

Liang:2025:RRC

[1041] Hong Liang, Xu Zhang, and Mingwen Shao. ReDepthNet: a radar and camera depth estimation model based on semantic segmentation mask region alignment. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07697-
                                        7.

Sun:2025:DEL

[1042] Yi Sun, Ruoxuan Huang, and Zhe Ming. A dual-enhanced long shortterm memory earthquake prediction method based on improved and hybrid rice-inspired gray wolf optimizers. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07677-x.

Zhou:2025:IAD

[1043] Yupeng Zhou, Hanhui Liu, and Minghao Yin. An incremental algorithm for dynamic graph coloring based on graph reduction and adaptive recoloring strategies. The Journal of Supercomputing, 81(12): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07696-8.

Wei:2025:MTC

[1044] Lei Wei, Pan Wang, and Anhua Wang. Multi-task classification offloading strategy based on reinforcement learning for mobile edge computing. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07693-x.

Zhao:2025:BDL

[1045] Hongzhi Zhao, Xun Liu, and Jinxiang Xie. Batched data layout optimization for Im2col-based convolutions on CPUs. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07706-9.

Shang:2025:DDL

[1046] Fanshu Shang, Yong Wang, and Jiaqi Liu. DLTPPM: a dual-layer model for trajectory privacy protection in autonomous driving. The Journal of Supercomputing, 81(12): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07689-7.

Li:2025:DDJ

[1047] Hao Li, Daowen Qiu, and Le Luo. Distributed Deutsch–Jozsa algorithm. The Journal of Supercomputing, 81 (12):??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07683-z.

Huang:2025:MAO

[1048] Xiaoxiao Huang, Kaibo He, and Dingchang Zheng. A model-agnostic ordinal regression pipeline for length of stay prediction. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07681-
                                        1.

Jiang:2025:SOM

[1049] Haiyin Jiang, Ruilin Wang, and Yanan Liu. Slerp-Opt: merging large language models via adaptive strategies. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07727-4.

Chen:2025:WTS

[1050] Jie Chen, Xinran Zhang, and Jiahao Guo. WARNs: a time series model integrating wavelet convolution and channel attention mechanism. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07704-x.

Nong:2025:CIE

[1051] Yumei Nong, Liwei Tian, and Zhaowen Li. Conditional information entropybased feature selection for partially labeled heterogeneous data via matrix operation and prediction label using k-nearest neighbor. The Journal of Supercomputing, 81(12): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07680-2.

Shen:2025:LYL

[1052] Qian Shen, Lei Zhang, and Shihao Liu.

LEDDR-YOLO: a lightweight and efficient distracted driving recognition algorithm with a particular pruning method. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07709-6.

Noori:2025:HHI

[1053] Sarah-Ahmad Noori, Mohammadreza Ramezanpour, and Reihaneh Khorsand. HIDE: high-integrity data embedding using dynamic carrier control in H.266/VVC streams. The Journal of Supercomputing, 81(12): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07736-3.

Liao:2025:PTZ

[1054] Bolin Liao, Jinsha Xu, and Shuai Li. Predefined-time ZNN model with noise reduction for solving quadratic programming and its application to binary assignment problem in logistics. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07694-w.

Zhang:2025:SRP

[1055] Xiuguo Zhang, Yuhang Cao, and Zhiwei Chen. Service reliability prediction methodology based on multivariate time series and improved AdaRNN model. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07712-x.

Wang:2025:IVM

[1056] Shuai Wang, Yifan Zhu, and Ming Yang. Infrared and visible military image fusion strategies and applications based on composite decomposition and multi-fuzzy theory. The Journal of Supercomputing, 81 (12):??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07731-8.

Mi:2025:DMA

[1057] Jian-Xun Mi, Lu Pan, and Xiao Cheng. Directing model attention to the discriminative foreground features. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07698-6.

Yin:2025:FEL

[1058] Jun Yin, Fei Wu, and Peng Huang.

FlowLoader: an efficient large language model loader via parallel transmissions and local multi-tier caching. The Journal of Supercomputing, 81 (12):??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07646-4.

Zhang:2025:FSM

[1059] Hongbo Zhang, Haohuan Nan, and Xueliang Gao. A feature selection method based on salp swarm algorithm with a multi-round voting mechanism. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07721-w.

Tang:2025:EAF

[1060] Yaqian Tang, Bai Yin, and Jianxi Fan.

An efficient algorithm to find a shorter fault-tolerant path in cycle composition networks. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07707-8.

Liu:2025:GFM

[1061] Yu Liu, Fu Li, and Chenhao Kong.

A generic framework for minimizing cold start times in serverless applications via resource serialization. The Journal of Supercomputing, 81 (12):??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07710-z.

Lin:2025:RRC

[1062] Yufeng Lin, Bo Liu, and Simon James Fong. RCCS: resource clusteringbased container scheduling for serverless edge intelligence. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07719-
4.

Deng:2025:SYE

[1063] Jiangzhou Deng, Lei Zheng, and Yong Wang. A simple yet effective enhanced collaborative filtering framework for mitigating noise and data sparsity: evidence from pervasive digital platform datasets. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07720-x.

Lei:2025:AGM

[1064] Tong Lei, Zongjing Chen, and Jian Zhang. Automatic GPU memory access optimization for AoSoA-based application in OP2 framework. The Journal of Supercomputing, 81(12): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07734-5.

Beikmohammadi:2025:HIF

[1065] Ali Beikmohammadi and Sindri Magnússon. Human-inspired framework to accelerate reinforcement learning. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07737-2.

Zheng:2025:SLO

[1066] Yingying Zheng, Ningjiang Chen, and Xinghui Gan. Sac-lstm: optimal resource allocation based on user intent in computility networks. The Journal of Supercomputing, 81(12): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07718-5.

Sha:2025:IHD

[1067] Mohemmed Sha and Adel Binbusayyis.

Implementing a hybrid deep learning technique for detecting malicious DNS over HTTPS (DoH) traffic. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07715-8.

Amer:2025:IEW

[1068] Dina A. Amer, Nader Y. Ibrahim, and Sarah A. Soliman. Intelligent eyes on water: YOLOv11-based realtime drowning detection system. The Journal of Supercomputing, 81(12): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07732-7.

Parizi:2025:QLQ

[1069] Mostafa Nazarian Parizi, Seyyed Hamid Ghafouri, and Mohammad Sadegh Hajmohammadi. QAT-LOADng: QoS aware trusted lightweight on-demand adhoc distance-vector routing-next generation. The Journal of Supercomputing, 81(12):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07722-
9.

Peng:2025:DEM

[1070] Yexin Peng, Mao Luo, and Hui Kong. A dual-evaluation-mode localsearch for the minimum dominating set problem on ultra-large sparse graphs. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07724-7.

Chen:2025:ILI

[1071] Fangshu Chen, Yilin Huang, and Xiankai Meng. Ifqa-llm: intelligent intention-driven financial questionanswering with large language models. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07726-5.

Nasrat:2025:SAR

[1072] Abdul Fatah Nasrat, Tuba Çaglikantar, and Necaattin Barisçi. SERDNet: adaptive residual dense learning with dual-path fusion and squeeze-and-excitation for robust image denoising. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07742-5.

Liu:2025:ICS

[1073] Dexin Liu, Xin Zhao, and Xuequan Zhang. Integrated control strategy for autonomous vehicle decision-making based on deep reinforcement learning. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07725-6.

Mei:2025:MMM

[1074] Wencheng Mei, Kangjian He, and Yiqiao Zhou. Multi-modality medical image fusion by edge supervising and multi-scale attention features extraction. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07601-3.

Singh:2025:PQS

[1075] Jagdeep Singh, Parminder Singh, and Mustapha Hedabou. Post-quantum secure fog-edge computing using federated learning with blockchain. The Journal of Supercomputing, 81(13):

??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07738-1.

Akhtar:2025:HAB

[1076] Md Akhtar, Goutam Mandal, and Asoke Kumar Bhunia. A hybrid algorithm based on differential evolution for reliability-redundancy allocation problem of a series system under multiple brands. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07708-7.

Liang:2025:MCS

[1077] Hong Liang, Mingchen Xu, and Mingwen Shao. Multi-cue SORT: integrating weak cues with appearance and motion for multi-object tracking. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07730-9.

Zhang:2025:BEO

[1078] Jiazheng Zhang, Shouwei Li, and Hongmei Pei. Blockchain-enabled one-stop efficient data retrieval privacy protection mechanism Industry 4.0. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07746-1.

Nandi:2025:IVS

[1079] Suman Nandi, Giridhar Maji, and Animesh Dutta. Identifying vital spreaders in multiplex networks: measurement of layer dominance and a closeness-based layer gravity method. The Journal of Supercomputing, 81 (13):??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07713-w.

Garcia-Vega:2025:PPP

[1080] Diego García-Vega, Fernando Plou-Llorente, and José Ranilla. pLazyQML: a parallel package for efficient execution of QML models on classical computers. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07714-9.

Bourian:2025:ESI

[1081] Imad Bourian, Chaimae Elfilali, and Khalid Choughdali. Enhancing security in intelligent transport system network by integrating blockchain-based smart contracts. The Journal of Supercomputing, 81 (13):??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07735-4.

Chen:2025:LSP

[1082] Siying Chen and Guoqiang Tang.

A layered stock prediction model based on novel feature selection and model parameter optimization. The Journal of Supercomputing, 81(13): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07745-2.

Chauhan:2025:AMB

[1083] Vikas Chauhan, Aruna Tiwari, and Arun Kumar. An attention mechanismbased hybrid TimeAttentionBiLSTM architecture for long-term traffic forecasting. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07747-0.

Hu:2025:PPT

[1084] Renyuan Hu, Zheyu Chen, and Zhiqiang Yao. Privacy preservation through makeup transfer for facial feature obfuscation. The Journal of Supercomputing, 81(13): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07756-z.

Ma:2025:EEP

[1085] Jin-Peng Ma and Yi-Wen Zhang. Energy-efficient partitioned semiclairvoyant scheduling in mixedcriticality system with graceful degradation. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07711-y.

Abdolhosseini:2025:SON

[1086] Hesam Abdolhosseini, Hamid R.

Zarandi, and Arman Gheysari. Simultaneous optimization of networkon-chip to improve reliability and reduce average packet latency considering buffer size constraints. The Journal of Supercomputing, 81(13): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07733-6.

Xi:2025:PHG

[1087] Qinghui Xi, Xi Li, and Lei Xu. Parallel heterogeneous graph learning based internet of things multivariate time series anomaly detection and explanation via cross-channel feature fusion. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07717-6.

Biswas:2025:PNS

[1088] Hillol Biswas. Power networks SCADA communication cybersecurity, a Qiskit implementation. The Journal of Supercomputing, 81(13): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07758-x.

Jiao:2025:GAD

[1089] Jiajia Jiao, Ling Jiang, and Ran Wen. GAN-assisted data augmentation to enhance detection accuracy of Evasive Spectre attacks. The Journal of Supercomputing, 81(13): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07691-z.

Liao:2025:FSR

[1090] Yanping Liao, Shengwen Lin, and Yihan He. Few-shot radar emitter signal recognition based on prototype network with filter system. The Journal of Supercomputing, 81 (13):??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07752-3.

Liu:2025:LMD

[1091] Dingshuo Liu, Yiran Liu, and Qingling Duan. LPTMono: monocular depth estimation for underwater images using local perception transformer and global-local context fusion. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07740-7.

Nayak:2025:FMA

[1092] Subhashish Nayak and P. M. Khilar. Fuzzy marine adaptive clustering with weighted similarity for accurate missing data imputation in electronic health records. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07729-2.

Yan:2025:SFC

[1093] Lirong Yan, Yongbing Zhang, and Xiaofen Tang. Spatial-frequency consistency and bias-corrected for few-shot object detection. The Journal of Supercomputing, 81(13): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07716-7.

Louja:2025:WSB

[1094] Ayoub Louja, Ines Drira, and Layth Sliman. Wearable sensorbased eye-rubbing monitoring: a hybrid CNN-attentionrub architecture for keratoconus prevention. The Journal of Supercomputing, 81(13): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07744-3.

Wan:2025:ORE

[1095] Hang Wan, Zhizhuo Qiu, and William Derigent. Optimizing renewable energy forecasting: a hybrid approach integrating MSADBO, BiGRU, and TCN for PV/wind power generation prediction. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07751-4.

Guan:2025:MCV

[1096] Chen Guan, Haihong Ai, and Ravi P.

Singh. Multi-CNNs with variational information bottleneck for chest X-ray classification. The Journal of Supercomputing, 81(13): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07767-w.

Liu:2025:PDS

[1097] Yuhang Liu, Yuqiao Zheng, and Yanqiang Zhang. Precise detection of surface defects on wind turbine blades for multi-scale target perception. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07761-2.

Chen:2025:CSD

[1098] Junfu Chen, Sirui Li, and Dechang Pi. Cross-subject domain adaptation for classifying working memory load with multi-frame EEG images. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07748-z.

Hamdouchi:2025:NDS

[1099] Abderahmane Hamdouchi and Ali Idri. New design strategies for IoT intrusion detection using boosting and feature selection. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07755-0.

Medjahed:2025:NOB

[1100] Seyyid Ahmed Medjahed and Fatima Boukhatem. A new optimizationbased framework for enhanced feature selection with the narwal optimizer. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07741-6.

Chari:2025:EFP

[1101] Ahmadreza Yosefzadeh Chari, Morteza Gholipour, and Mohammadreza Hassanzadeh. Efficient fully parallel convolutional neural network architecture using 1-memristor and 1-transistor (1M1T). The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07796-5.

Garcia:2025:TRT

[1102] Pedro Henrique Sachete Garcia, Marcelo Caggiani Luizelli, and Fábio Diniz Rossi. Toward real-time IoT multi-sensor data orchestration on wireless sensor networks. The Journal of Supercomputing, 81(13): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07749-y.

Ancha:2025:TYT

[1103] Vinod Kumar Ancha, Venkateswarlu Gonuguntla, and Ramesh Vaddi. TRSBi-YOLO: Transformer based lightweight and high-performance model for PCB defects detection. The Journal of Supercomputing, 81 (13):??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07771-0.

Elghandour:2025:PNB

[1104] Mohamed Elghandour. PCA-NCA based VIT features for classification of white blood cell images. The Journal of Supercomputing, 81(13): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07753-2.

Liang:2025:LML

[1105] Hong Liang, Quanyi Guo, and Qian Zhang. LSA-MEP: layer-wise sparsity allocation multi-metric evaluation pruning. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07786-7.

Miao:2025:NSO

[1106] JinFeng Miao, Yanyi Huang, and Xuechen Zhao. NoTNER: selfoptimizing text reconstruction for open named entity recognition on social media. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07764-z.

Cao:2025:UTM

[1107] Zhong Cao, Jiang Lu, and Yuheng Luo. Unsupervised tractive momentum: a novel unsupervised fewshot learning framework. The Journal of Supercomputing, 81(13): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07757-y.

Yuan:2025:NPP

[1108] Xiao Yuan, Shengwei Tian, and Zhezhe Zhu. Non parametric 3D point cloud understanding based on curvature guidance. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07766-x.

He:2025:EMN

[1109] Xinyu He, Shixin Li, and Binhe Li. Enhancing multimodal named entity recognition with multi-granularity knowledge distillation. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07772-
z.

Ghazaan:2025:ASB

[1110] Majid Ilchi Ghazaan and Kasra Laal Dehghani. An adaptive surrogatebased approach for solving constrained optimization problems. The Journal of Supercomputing, 81(13): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07759-w.

Wang:2025:CDC

[1111] Jiale Wang, Yitao Liang, and Juan Xia. CELAT-DiffNet: channelenhanced local attention transformer for underwater image enhancement based on diffusion models. The Journal of Supercomputing, 81(13): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07769-8.

Kayadibi:2025:MEY

[1112] Ismail Kayadibi, Utku Köse, and Bilgün Çetin. E-MTMYOLO: an explainable YOLOv5-based architecture for accurate detection of mandibular third molar using a novel expert-annotated dataset. The Journal of Supercomputing, 81(13): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07775-w.

Du:2025:RUM

[1113] Haishun Du, Linbing Cao, and Wenzhe Zhang. Refining unified model by test-time dynamic adaptation for domain shift problem. The Journal of Supercomputing, 81(13): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07770-1.

Keyhanipour:2025:GAW

[1114] Amir Hosein Keyhanipour. Graphaugmented web spam detection using evidential reasoning. The Journal of Supercomputing, 81(13): ??, August 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07778-7.

Singh:2025:EFF

[1115] Narendra Singh and Somanath Tripathy. FedShield: federated learning based robust online payment fraud detection. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07774-x.

Yao:2025:DBD

[1116] Daojin Yao, Hanxin Chen, and Yongxiang Yu. A dangerous behavior detection algorithm with the fusion of RGB data and skeleton information. The Journal of Supercomputing, 81(13):??, August 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07765-y.

Qiang:2025:UYI

[1117] Qiuxi Qiang, Mingshu Zhang, and Alin Hou. UDS-YOLO: an improved instance segmentation network for traffic scenarios. The Journal of Supercomputing, 81(14): ??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07779-6.

Ren:2025:MOE

[1118] Wei Ren, Fangzhen Ge, and Huaiyu Liu. A many-objective evolutionary algorithm based on decision variable classification mutation and indicator. The Journal of Supercomputing, 81 (14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07705-w.

Wang:2025:SYH

[1119] Zixuan Wang, Dezhi Han, and Xiaohu Huang. SVN-YOLO: a highprecision ship detection algorithm based on improved YOLOv10n. The Journal of Supercomputing, 81(14): ??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07743-4.

Zhou:2025:EHP

[1120] Yi-Hua Zhou, Wei-Yi Pu, and Wei-Min Shi. An efficient hybrid protocol for quantum secure multi-party summation, multiplication and comparison based on d-dimensional single-particle states. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07763-0.

Li:2025:DGN

[1121] Muyang Li, Mingjian Liu, and Dianchen Liu. Dynamic graph neural network-based vehicle trajectory prediction method. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07739-0.

Yan:2025:UGA

[1122] Lirong Yan, Yongbing Zhang, and Xiaofen Tang. Uncertainty-guided alignment optimization and pseudolabel refinement for cross-domain few-shot object detection. The Journal of Supercomputing, 81(14): ??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07750-5.

Danismaz:2025:TPD

[1123] Sibel Danismaz, Ramazan Ozgur Dogan, and Hulya Dogan. Two-phase deep learning method for image fusionbased extended depth of focus. The Journal of Supercomputing, 81(14): ??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07791-w.

Karimzadeh:2025:RDS

[1124] Masoud Karimzadeh, Parham Moradi, and Abdulbaghi Ghaderzadeh. Robust dual space factorization for unsupervised feature selection (RDSF-UFS). The Journal of Supercomputing, 81(14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07780-z.

Boudi:2025:ECP

[1125] Raid Boudi, Mohammed Lalou, and Nardjes Bouchemal. Enhancing controller placement in SD-IoT with dynamic heterogeneous switch traffic via weighted betweenness and GWO. The Journal of Supercomputing, 81 (14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07783-w.

Xu:2025:EUO

[1126] Xuebin Xu, Ziyang Xing, and Kuihe Yang. Enhancing UAV object detection through multi-scale deformable convolutions and adaptive fusion attention. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07788-5.

Alem:2025:RRR

[1127] Ismahene Alem, Samira Yessad, and Louiza Bouallouche-Medjkoune. RIMS-RPL: RL-based intelligent mobilitysupport of RPL for IIoT-based networks. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07723-8.

Ghaseminya:2025:ACV

[1128] Mohammad Mahdi Ghaseminya, Elahe Eslami, and Seyed Mehdi Karbassi.

Advancing cloud virtualization: a comprehensive survey on integrating IoT, Edge, and Fog computing with FaaS for heterogeneous smart environments. The Journal of Supercomputing, 81 (14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07799-2.

Tian:2025:MCA

[1129] Wentao Tian, Anqi Pan, and Yaqi Liao. A multi-classifier-assisted constrained optimization algorithm for obstacle avoidance trajectory planning of robotic arm. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07785-8.

Wang:2025:TTS

[1130] Chang Wang, Peiyu Li, and Yusong Lin. TSFN: a time-series fusion network for keystroke authentication. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07797-4.

Jiang:2025:ERS

[1131] Yuzhu Jiang, Kuaini Wang, and Jinge Li. Enhancing robustness and sparsity of extreme learning machine in regression problems using correntropy function. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07620-0.

Xiong:2025:DPC

[1132] Wei Xiong, Haina Song, and Yan Li. A dense pyramid convolutional neural network for MRI brain tumor segmentation. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07805-7.

Ali:2025:MAB

[1133] Zeeshan Ali, Muhammad Waqas, and Kostaq Hila. Multi-attributive border approximation area comparison model based on Yager weighted aggregation operators for circular Pythagorean fuzzy information and their application in shortest path problems. The Journal of Supercomputing, 81 (14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07327-2.

Zhang:2025:MTS

[1134] Qifan Zhang, Houqun Yang, and Jianqiang Huang. A multivariate time series data classification method based on dual-branch structure GSTT. The Journal of Supercomputing, 81 (14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07781-y.

Agrawal:2025:QIA

[1135] Umang Kumar Agrawal and Nibedan Panda. Quantum-inspired adaptive mutation operator enabled PSO (QAMO-PSO) for parallel optimization and tailoring parameters of Kolmogorov-Arnold network. The Journal of Supercomputing, 81(14): ??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07810-w.

Gumuscekicci:2025:SDN

[1136] Gizem Gumuscekicci and Rahim Dehkharghani. Sarcasm detection on news headlines using transformers. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07773-y.

Li:2025:LIB

[1137] Weiwei Li, Chen Chen, and Deyun Chen. Language identification based on multi-scale feature recursive fusion and adaptive loss. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07808-4.

Chen:2025:EMO

[1138] Lei Chen, Wenhao Liu, and Guanli Wang. An evolutionary multitask optimization algorithm based on block-level knowledge transfer and beluga whale optimization. The Journal of Supercomputing, 81(14): ??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07787-6.

Hong:2025:SST

[1139] Yuling Hong and Jiaqi Zhang. STMACN: spatial-temporal multiscale auto-correlation network for traffic forecasting. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07809-3.

Yang:2025:IDB

[1140] Gaoming Yang, Peng Zhu, and Xiangyu Yang. Interactive dualbranch network based on adversarial knowledge distillation for compressed deepfake detection. The Journal of Supercomputing, 81(14): ??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07795-6.

Wu:2025:ULS

[1141] Yuxiang Wu, Qianjin Zhao, and Kuan-Ching Li. Ultra-lightweight SAR ship object detection based on multiscale fusion and pruning distillation. The Journal of Supercomputing, 81 (14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07782-x.

Shuai:2025:POG

[1142] Yong Shuai, Dingcai Shen, and Manman Shen. Potential-optima guided adaptive neighborhood differential evolution algorithm for multimodal optimization problems. The Journal of Supercomputing, 81(14): ??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07762-1.

Chen:2025:FAL

[1143] Jinhua Chen, Mengmeng Li, and Xin He. FedCK:addressing label distribution skew in federated learning via clustering-efficient and knowledge distillation. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07728-3.

Asl:2025:REE

[1144] Arman Sanaei Asl and Massoud Reza Hashemi. Resilient and efficient edge server placement in mobile edge computing: a dual-strategy approach for optimized performance and fault tolerance. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07807-5.

Pandi:2025:TTL

[1145] Marziye Pandi, Mostafa Moghaddas, and Hakem Beitollahi. TestLock: a testability logic locking method against machine learning-based oracle-less attacks. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07803-9.

Meng:2025:RDF

[1146] Xiankai Meng, Xiang Ji, and Jiang Wu. RTL design flaws revisited: a data-driven study of systematic bug patterns in Verilog code. The Journal of Supercomputing, 81(14): ??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07811-9.

Jimenez:2025:SRL

[1147] Marvin Jiménez, Jose Aguilar, and Edwin Montoya. A systematic review of the literature on recent trends in dynamic combinatorial optimization problems linked to meta-learning in metaheuristics and quantum computing. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07792-9.

Yuan:2025:FBE

[1148] Ming Yuan, Qiang Liu, and Lin Gan. An FPGA-based efficient accelerator for fault interaction of rupture dynamics. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07801-x.

Rafiei:2025:CEF

[1149] Hanieh Rafiei, Mojtaba Mahdavi, and Ahmad Reza NaghshNilchi. A comprehensive evaluation framework for EMD-based steganography: balancing capacity and security with upper bound analysis. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07804-8.

Saeidi:2025:PQA

[1150] Mohammad Reza Saeidi and Hamid Mala. Post-quantum anonymous password authenticated key exchange with indirect public key validation. The Journal of Supercomputing, 81(14): ??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07793-8.

Trinh:2025:MTT

[1151] Thanh Binh Trinh, Van Hieu Vu, and Thi Van Nguyen. A multitask test case optimization framework with integrated explainable AI for customer churn prediction. The Journal of Supercomputing, 81(14): ??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07816-4.

Ken:2025:ECH

[1152] Pao-Yi Ken and Chao-Chin Wu. Efficient configuration of heterogeneous resources and task scheduling strategies in deep learning auto-tuning systems. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07798-3.

Wang:2025:FFG

[1153] Ying Wang, Jiuyun Xu, and Hina Batool. Fgcfl: a fine-grained clustering framework for federated learning with heterogeneity data. The Journal of Supercomputing, 81(14): ??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07813-7.

Liu:2025:ERO

[1154] Zifeng Liu, Lujiao Li, and Shigang Hu. Enhanced road object detection with DFPD-YOLO: focusing on small and occluded targets. The Journal of Supercomputing, 81(14): ??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07815-5.

Majidpour:2025:MOF

[1155] Jafar Majidpour and Hakem Beitollahi. Multi-objective feature selection of radiomics and deep learning features for breast cancer subtype detection. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07768-9.

Chen:2025:AGD

[1156] Lei Chen, Yikai Zhao, and Wenhao Liu.

Advancing green development in 3D medical image segmentation: a novel lightweight, semi-supervised framework. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07824-4.

Zhang:2025:RYS

[1157] Jinhang Zhang, Min Gao, and Chenglin Rong. REA-YOLO for small object detection in UAV aerial images. The Journal of Supercomputing, 81 (14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07836-0.

Xue:2025:WWH

[1158] Shuxin Xue, Zhaohui Zhang, and Jing Meng. WHANet: wavelet and hybrid attention network for vessel segmentation in OCTA fundus images. The Journal of Supercomputing, 81 (14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07794-7.

Swathi:2025:SOD

[1159] M. Swathi and Rajeshkannan Regunathan. Speech organ disease classification using bidirectional block selfadaptive attention mechanism enabled quantum convolutional neural network. The Journal of Supercomputing, 81 (14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07800-y.

Chitla:2025:PED

[1160] Vinay Santhosh Chitla, Hemantha Kumar Kalluri, and Satya Krishna Nunna. Performance evaluation of diverse graph-based models on homogeneous datasets. The Journal of Supercomputing, 81(14): ??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07826-2.

Al-Sarray:2025:RSI

[1161] Naseer Hameed Saadoon Al-Sarray, Ayşe Demirhan, and Javad Rahebi. A robust and scalable intrusion detection framework for SDN with GAN-CL-STO. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07821-7.

Zhang:2025:SYA

[1162] Runze Zhang, Ce Wang, and Xudong Xu. Splp-yolo: an all-weather real-time detector for airport runway foreign object debris. The Journal of Supercomputing, 81(14): ??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07818-2.

Frinkle:2025:CSC

[1163] Karl Frinkle, Saxon Rowland, James Pack, Madison Gordon, John Eskue, Sardonyx Egbe, and Mike Morris. Computing streaks of consecutive numbers with the same Collatz height. The Journal of Supercomputing, 81 (14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07820-8.

Omidian:2025:AAA

[1164] Fatemeh Omidian, Athena Abdi, and Alireza Hamed-Rouhbakhs. AdApTS: adaptive approximate computingbased traffic sign recognition unit for self-driving cars. The Journal of Supercomputing, 81(14):??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07825-3.

Huang:2025:PRM

[1165] Xing Huang, Kexin Wang, and Junnan Li. Prototype reduction method based on accelerated binary barebone particle swarm optimization for instance-based classifiers. The Journal of Supercomputing, 81(14): ??, September 2025. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07518-x.

Matinfar:2025:EEN

[1166] Ehsan Matinfar, Marjan Mahmoudi, and Behrang Barekatain. ELDM-EDSDN: a novel resource allocation method in edge-DSDN-based networks using improved manifold approach. The Journal of Supercomputing, 81 (14):??, September 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-025-07817-3.